### v1

# Joint Processing and Transmission Energy Optimization for ISAC in Cell-Free Massive MIMO with URLLC

Abstract
In this paper, we explore the concept of integrated sensing and communication (ISAC) within a downlink cell-free massive MIMO (multiple-input multiple-output) system featuring multi-static sensing and users requiring ultra-reliable low-latency communications (URLLC). Our focus involves the formulation of two non-convex algorithms that jointly solve power and blocklength allocation for end-to-end (E2E) minimization. The objectives are to jointly minimize sensing/communication processing and transmission energy consumption, while simultaneously meeting the requirements for sensing and URLLC. To address the inherent non-convexity of these optimization problems, we utilize techniques such as the Feasible Point Pursuit - Successive Convex Approximation (FPP-SCA), Concave-Convex Programming (CCP), and fractional programming. We conduct a comparative analysis of the performance of these algorithms in ISAC scenarios and against a URLLC-only scenario where sensing is not integrated. Our numerical results highlight the superior performance of the E2E energy minimization algorithm, especially in scenarios without sensing capability. Additionally, our study underscores the increasing prominence of energy consumption associated with sensing processing tasks as the number of sensing receive access points rises. Furthermore, the results emphasize that a higher sensing signal-to-interference-plus-noise ratio threshold is associated with an escalation in E2E energy consumption, thereby narrowing the performance gap between the two proposed algorithms.
Index Terms:
Integrated sensing and communication, cell-free massive MIMO, URLLC, power allocation, blocklength
I Introduction
Integrated sensing and communication (ISAC) aims to enhance spectral and hardware efficiency, offering potential applications in 6G communication networks, including autonomous vehicles, smart factories, and location/environment-aware scenarios [1, 2]. Particularly, in ultra-reliable low-latency communication (URLLC) use cases like industrial Internet-of-Things (IIoT) and vehicle-to-everything (V2X) communications, ISAC can play a crucial role. For instance, in V2X networks, ISAC systems can detect targets such as pedestrians and vehicles, delivering sensing information reliably and with minimal delay for URLLC users like autonomous vehicles [3]. Thus, a comprehensive study of these technologies together is imperative to enable mission-critical applications in future wireless networks.
For many URLLC applications, such as V2X and IIoT, short codewords are needed to satisfy latency constraints where codes with short blocklengths, e.g., 100 symbols are employed. Therefore, the finite blocklength regime needs to be considered to model decoding error probability (DEP) [4, 5, 6]. In such applications, short data packets should be delivered with a minimum reliability of 99.999% and an end-to-end (E2E) delay of less than 10-150 ms [7, 8]. Therefore, optimizing the blocklength can play an important role to meet the reliability and delay requirements.
Conversely, while ISAC introduces perceptive mobile networks through network-level sensing [9, 10, 11, 12], it is anticipated that ISAC networks may consume more energy compared to communication networks due to sensing tasks. This motivates a comprehensive study of the end-to-end (E2E) energy consumption of ISAC networks taking into account not only energy consumption for transmission but also energy consumption for processing tasks. Centralized radio access network (C-RAN) architecture has been already introduced as an energy-efficient solution for the mobile networks. There is also an inherent connection between cell-free massive MIMO (multiple-input multiple-output) and the C-RAN architecture due to the joint transmission and processing the UEs’ signals[13]. However, E2E energy-aware resource orchestration is a critical aspect that must be considered when optimizing cell-free massive MIMO systems on C-RAN architectures to fully benefit from energy savings potential [14].
I-A Related Work
To the best of our knowledge, there are few works that jointly consider URLLC and ISAC. In [3], a joint precoding scheme is proposed to minimize transmit power, satisfying sensing and delay requirements. Moreover, joint ISAC beamforming and scheduling design is addressed in [15] for periodic and aperiodic traffic. The existing works consider a single ISAC base station, while cell-free massive MIMO with multiple distributed access points (APs) has been shown to facilitate URLLC [6, 16, 17] and ISAC[18, 19] implementations.
In the conference version [20] of this work, we studied for the first time ISAC-enabled cell-free massive MIMO with URLLC users. In this paper, we extend [20] from an E2E energy minimization perspective. The consideration of E2E energy-awareness has been explored in various contexts, as reflected in prior works such as [21, 22, 23, 14, 24]. In particular, [14] studied fully virtualized E2E power minimization problem for cell-free massive MIMO on open radio access network (O-RAN) architecture by taking the radio, fronthaul, and processing resources into account. However, this work neither considered ISAC nor URLLC communications.
I-B Contributions
In this paper, we study E2E energy consumption in a cell-free massive MIMO system with URLLC user equipments (UEs) and multi-static sensing in a cluttered environment. Both sensing and communication signals contribute to the sensing task. However, the sensing signals can cause interference for the UEs. Thus, we design the precoding vectors to null the interference only for the UEs. We consider a maximum DEP threshold, representing the reliability requirement, together with a maximum transmission delay threshold as the URLLC requirements and a minimum signal-to-interference-plus-noise ratio (SINR) as the sensing requirement. The latter is motivated by the fact that given a fixed false alarm probability, detection probability increases with a higher sensing SINR [25], [26, Chap. 3 and 15]. In the conference version of this work [20], we propose a successive convex approximation-based power allocation algorithm that maximizes energy efficiency while satisfying the sensing and URLLC requirements for a given blocklength. However, the impact of blocklength optimization is not considered in [20].
Different from the conference paper, we conduct a comprehensive analysis of the processing requirements, specifically in terms of giga operations per second (GOPS), and examine its impact on processing energy consumption. The interplay between processing requirements and blocklength yields intriguing results, as demonstrated by the simulation findings presented in this paper.
The main contributions of this paper are outlined as follows:
  • We derive an upper bound for the DEP and transmission delay in finite blocklength cell-free massive MIMO communications.
  • We propose two joint blocklength and power allocation algorithms, which minimize the transmission energy consumption and E2E energy consumption, respectively while meeting URLLC and sensing requirements. The former algorithm is represented by SeURLLC+ and the latter is represented by E2E URLLC+. To solve the non-convex optimization problems, we benefit from the Feasible Point Pursuit - Successive Convex Approximation (FPP-SCA), Concave-Convex Programming (CCP), and fractional programming.
  • We investigate the impact of E2E energy minimization, encompassing both radio and processing energy consumption, in the above-mentioned ISAC scenario, as well as a scenario where only URLLC requirements are taken into account. The algorithms for URLLC-only scenarios are represented by URLLC-only and E2E URLLC-only. The former corresponds to transmission energy minimization and the latter corresponds to E2E energy minimization.
  • We conduct a sensitivity analysis by investigating the impact of DEP threhsold, maximum delay threshold, and sensing SINR threshold on E2E energy consumption.
The rest of the paper is organized as follows. In Section II, we present the system model. In Section III, we provide URLLC analysis accounting for decoding error probability and delay in finite blocklength regime. Section IV is devoted to the sensing analysis. In Section V, we describe the E2E power model and derive the GOPS analysis for both communication and sensing. Next, we present our optimization problems in Section VI. Finally, the numerical results and conclusion are provided in Section VII and Section VIII, respectively.
II System Model
We consider an ISAC-enabled cell-free massive MIMO system in URLLC scenarios. The system adopts a centralized radio access network (C-RAN) architecture [27] for uplink channel estimation and downlink communication, as well as multi-static sensing as shown in Fig. 1. All the APs are interconnected via fronthaul links to the edge cloud and operate in full synchronization. We consider the original form of cell-free massive MIMO [6], wherein all the $N_{\rm tx}$ ISAC APs jointly serve the $N_{\rm ue}$ URLLC UEs by transmitting precoded signals containing both communication and sensing symbols.
Figure 1: Illustration of the ISAC system in cell-free massive MIMO with URLLC UEs.
Concurrently, the $N_{\rm rx}$ sensing receive APs engage in simultaneous sensing of the candidate location to detect the target. Each AP is equipped with an array of $M$ antennas configured in a horizontal uniform linear array (ULA) with half-wavelength spacing. The respective array response vector is $\textbf{a}(\varphi,\vartheta)=\begin{bmatrix}1\!&\!e^{j\pi\sin(\varphi)\cos(\vartheta)}&\ldots&e^{j(M-1)\pi\sin(\varphi)\cos(\vartheta)}\end{bmatrix}^{T},$ where $\varphi$ and $\vartheta$ are the azimuth and elevation angles from the AP to the target location, respectively [28].
We consider the finite blocklength regime for URLLC UEs, where a packet of $b_{i}$ bits is sent to UE $i$ within a transmission block with blocklength $L=L_{p}+L_{d}$ symbols using the coherence bandwidth $B$. $L_{p}$ and $L_{d}$ are the number of symbols for pilot and data, respectively. It is expected that duration of each URLLC transmission, denoted by $T$, is shorter than one coherence time $T_{c}$, i.e., $T<T_{c}$ [17]. Without loss of generality, we assume the coherence time is equal to URLLC transmission time so that in for each transmission we first estimate the channel.
II-A Channel Modeling and Downlink Communications
The communication channels are modeled as spatially correlated Rician fading which are assumed to remain constant during each coherence block, and the channel realizations are independent of each other. Let $\textbf{h}_{i,k}\in\mathbb{C}^{M}$ denote the channel between ISAC AP $k$ and UE $i$, modeled as
$\textbf{h}_{i,k}=e^{j\varphi_{i,k}}\bar{\textbf{h}}_{i,k}\,+\,\tilde{\textbf{h}}_{i,k},$ (1)
which consists of a semi-deterministic line-of-sight (LOS) path, represented by $e^{j\varphi_{i,k}}\bar{\textbf{h}}_{i,k}$ with unknown phase-shift $\varphi_{i,k}\sim\mathcal{U}[0,2\pi)$, i.e., uniformly distributed on $[0,2\pi)$, and a stochastic non-LOS (NLOS) component $\tilde{\textbf{h}}_{i,k}\sim\mathcal{CN}(\textbf{0},\textbf{R}_{i,k})$ with the spatial correlation matrix $\textbf{R}_{i,k}\in\mathbb{C}^{M\times M}$. Both $\bar{\textbf{h}}_{i,k}$ and $\textbf{R}_{i,k}$ include the combined effect of geometric path loss and shadowing. We concatenate the channel vector $\textbf{h}_{i,k}$ in the collective channel vector
$\textbf{h}_{i}=\begin{bmatrix}\textbf{h}_{i,1}^{T}&\ldots&\textbf{h}_{i,N_{\rm tx}}^{T}\end{bmatrix}^{T}\in\mathbb{C}^{N_{\rm tx}M}$ (3)
for UE $i$.
Let $s_{i}[m]$ and $s_{0}[m]$ represent the downlink communication symbol for UE $i$ and sensing symbol, respectively at time instance $m$. The symbols are independent and have zero mean and unit power. Moreover, let $\rho_{i}\geq 0$ and $\rho_{0}\geq 0$ be, respectively, the power control coefficients for UE $i$ and the target, which are fixed throughout the transmission. Then, the transmitted signal from transmit AP $k\in\{1,\cdots,N_{\rm tx}\}$ at time instance $m$ is
$\vspace{-1mm}\textbf{x}_{k}[m]=\sum_{i=0}^{N_{\rm ue}}\sqrt{\rho_{i}}\textbf{w}_{i,k}s_{i}[m]=\textbf{W}_{k}\textbf{D}_{\rm s}[m]\boldsymbol{\rho},\vspace{-1mm}$ (4)
where the vectors $\textbf{w}_{i,k}\in\mathbb{C}^{M}$ and $\textbf{w}_{0,k}\in\mathbb{C}^{M}$ are the transmit precoding vectors for transmit AP $k$ corresponding to UE $i$ and the sensing signal, respectively. In (4), $\textbf{W}_{k}=\begin{bmatrix}\textbf{w}_{0,k}&\textbf{w}_{1,k}&\cdots&\textbf{w}_{N_{\rm ue},k}\end{bmatrix}$, ${\textbf{D}}_{\rm s}[m]=\textrm{diag}\left(s_{0}[m],s_{1}[m],\ldots,s_{N_{\rm ue}}[m]\right)$ is the diagonal matrix containing the sensing and communication symbols, and $\boldsymbol{\rho}=[\sqrt{\rho_{0}}\ \ldots\ \sqrt{\rho_{N_{\rm ue}}}]^{T}$.
Moreover, let $n_{i}[m]\!\sim\!\mathcal{CN}(0,\sigma_{n}^{2})$ be the independent receiver noise at UE $i$ at time instance $m$. The received signal at UE $i$ is given as
$y_{i}[m]=$ $\underbrace{\sqrt{\rho_{i}}\textbf{h}_{i}^{H}\textbf{w}_{i}s_{i}[m]}_{\textrm{Desired signal}}+\underbrace{\sum_{j=1,j\neq i}^{N_{\rm ue}}\sqrt{\rho_{j}}\textbf{h}_{i}^{H}\textbf{w}_{j}s_{j}[m]}_{\textrm{Interference signal due to the other UEs}}$
$+\underbrace{\sqrt{\rho_{0}}\textbf{h}_{i}^{H}\textbf{w}_{0}s_{0}[m]}_{\textrm{Interference signal due to the sensing}}+\underbrace{n_{i}[m]}_{\textrm{Noise}},$ (5)
where the collective precoding vectors
${\textbf{w}_{i}=\left[\textbf{w}_{i,1}^{T}\ \textbf{w}_{i,2}^{T}\ \ldots\ \textbf{w}_{i,N_{\rm tx}}^{T}\right]^{T}\in\mathbb{C}^{N_{\rm tx}M}},$ (6)
for $i=1,\ldots,N_{\rm ue}$, and
$\textbf{w}_{0}=\left[\textbf{w}_{0,1}^{T}\ \textbf{w}_{0,2}^{T}\ \ldots\ \textbf{w}_{0,N_{\rm tx}}^{T}\right]^{T}\in\mathbb{C}^{N_{\rm tx}M},$ (7)
are the centralized precoding vectors. The vectors $\textbf{w}_{i,k}\in\mathbb{C}^{M}$ and $\textbf{w}_{0,k}\in\mathbb{C}^{M}$ are the transmit precoding vectors for ISAC AP $k$ corresponding to UE $i$ and the sensing signal, respectively.
Using the independence of the data and sensing signals, the average transmit power for transmit AP $k\in\{1,\ldots,N_{\rm tx}\}$ is
$P_{k}=\sum_{i=0}^{N_{\rm ue}}\rho_{i}\mathbb{E}\left\{\|\textbf{w}_{i,k}\|^{2}\right\},\quad k=1,\ldots,N_{\rm tx}$ (8)
which should not exceed the maximum power limit $P_{\rm tx}$.
In the subsequent section, we will present the specifics of the linear minimum mean-squared error (LMMSE)-based channel estimation.
II-B LMMSE-Based Channel Estimation
Within the URLLC transmission time comprising $L$ symbols, $L_{p}<L$ symbols are allocated for pilot transmission. A collection of $L_{p}$ mutually orthogonal pilot sequences is represented by $\boldsymbol{\phi}_{1},\ldots,\boldsymbol{\phi}_{L_{p}}\in\mathbb{C}^{L_{p}}$, where $\|\boldsymbol{\phi}_{t}\|^{2}=L_{p}$ for $t=1,\ldots,L_{p}$. These pilot sequences are assigned to the $N_{\rm ue}$ UEs. If the number of pilot sequences is less than the number of UEs (i.e., $L_{p}<N_{\rm ue}$), then each pilot sequence may be assigned to multiple UEs. Let $t_{i}\in\{1,\ldots,L_{p}\}$ denote the index of the pilot sequence that is assigned to UE $i$ and we define $\mathcal{P}_{i}=\{j\in\{1,\ldots,N_{\rm ue}\}|t_{j}=t_{i}\}$ as the set of UEs that share the same pilot with UE $i$.
The received signal at the ISAC AP $k$ during the entire pilot transmission is denoted by $\textbf{Y}_{k}^{p}\in\mathbb{C}^{M\times L_{p}}$ where
$\textbf{Y}_{k}^{p}=\sum_{i=1}^{N_{\rm ue}}\sqrt{\eta_{i}}\textbf{h}_{i,k}\boldsymbol{\phi}_{t_{i}}^{T}+\textbf{N}_{k},$ (9)
where $\eta_{i}>0$ is the transmit power by UE $i$ and $\textbf{N}_{k}\in\mathbb{C}^{M\times L_{p}}$ is the received noise at AP $k$ with independent and identically distributed (i.i.d.) $\mathcal{CN}(0,\sigma_{n}^{2})$ entries. We assume that $\bar{\textbf{h}}_{i,k}$ and $\textbf{R}_{i,k}$, representing the long-term statistics for the channel in (1), are known while the random phase-shift $\varphi_{i,k}$ is unknown. With respect to estimating $\textbf{h}_{i,k}$, let us form $\textbf{y}_{t_{i},k}^{p}=\frac{1}{\sqrt{L_{p}}}\textbf{Y}_{k}^{p}\boldsymbol{\phi}_{t_{i}}^{*}\in\mathbb{C}^{M}$, where
$\textbf{y}_{t_{i},k}^{p}=\underbrace{\sqrt{\eta_{i}L_{p}}\textbf{h}_{i,k}}_{\textrm{Desired part}}+\underbrace{\sum_{j\in\mathcal{P}_{i}\setminus\{i\}}\sqrt{\eta_{j}L_{p}}\textbf{h}_{j,k}}_{\textrm{Interference part}}+\underbrace{\textbf{n}_{t_{i},k}}_{\textrm{Noise}},$ (10)
where $\textbf{n}_{t_{i},k}=\frac{1}{\sqrt{L_{p}}}\textbf{N}_{k}\boldsymbol{\phi}_{t_{i}}^{*}\sim\mathcal{CN}(\textbf{0},\sigma_{n}^{2}\textbf{I}_{M})$. If the number of UEs is greater than the number of mutually orthogonal pilot sequences, then each pilot sequence may be assigned to multiple UEs using the pilot assignment algorithm in [27, Alg. 4.1]. The LMMSE estimate of $\textbf{h}_{i,k}$ is given by [29]
$\hat{\textbf{h}}_{i,k}=\sqrt{\eta_{i}L_{p}}\textbf{R}_{i,k}^{\prime}\left(\boldsymbol{\Psi}^{\prime}_{t_{i},k}\right)^{-1}\textbf{y}_{t_{i},k}^{p},$ (11)
where $\textbf{R}_{i,k}^{\prime}$ and $\boldsymbol{\Psi}_{t_{i},k}^{\prime}$ are the spatial correlation matrix corresponding to channel $\textbf{h}_{i,k}$ and the correlation matrix of the received signal, respectively, given by
$\textbf{R}_{i,k}^{\prime}\triangleq\textbf{R}_{i,k}+\bar{\textbf{h}}_{i,k}\bar{\textbf{h}}_{i,k}^{H}$ (12)
$\boldsymbol{\Psi}^{\prime}_{t_{i},k}\triangleq\sum_{j\in\mathcal{P}_{i}}\eta_{j}L_{p}\textbf{R}_{j,k}^{\prime}+\sigma_{n}^{2}\textbf{I}_{M}.$ (13)
It is worth noting that deriving MMSE channel estimation proves challenging due to the non-Gaussian distribution of channels resulting from random phase-shifts in the LOS paths. For the sake of tractability, we adhere to the use of LMMSE-based channel estimation in this paper.
II-C Multi-Static Sensing
We investigate multi-static sensing, involving multiple transmit and receive APs in the network. The network engages in target location sensing during the downlink phase. We assume the existence of a LOS connection between the target location and each transmit/receive AP.
In the presence of the target, each receive AP captures both the reflected signals from the target and undesired signals, referred to as clutter. The clutter, being independent of the target’s presence, is treated as interference for sensing purposes. Without loss of generality, we assume that the LOS paths between transmit and receive APs are known and can be effectively canceled out. Consequently, the interference signals correspond to the reflected paths through obstacles and are henceforth denoted as target-free channels.
Let $\textbf{H}_{r,k}\in\mathbb{C}^{M\times M}$ denote the target-free channel matrix between transmit AP $k$ and receive AP $r$. We use the correlated Rayleigh fading model for the NLOS channels $\textbf{H}_{r,k}$, which is modeled using the Kronecker model [30]. We define the random matrix $\textbf{W}_{{\rm ch},(r,k)}\in\mathbb{C}^{M\times M}$ with i.i.d. entries with $\mathcal{CN}(0,1)$ distribution. The matrix $\textbf{R}_{{\rm rx},(r,k)}\in\mathbb{C}^{M\times M}$ represents the spatial correlation matrix corresponding to receive AP $r$ and with respect to the direction of transmit AP $k$. Similarly, $\textbf{R}_{{\rm tx},(r,k)}\in\mathbb{C}^{M\times M}$ is the spatial correlation matrix corresponding to transmit AP $k$ and with respect to the direction of receive AP $r$. The channel $\textbf{H}_{r,k}$ is written as
$\textbf{H}_{r,k}=\textbf{R}^{\frac{1}{2}}_{{\rm rx},(r,k)}\textbf{W}_{{\rm ch},{(r,k)}}\left(\textbf{R}^{\frac{1}{2}}_{{\rm tx},(r,k)}\right)^{T},$ (14)
where the channel gain is determined by the geometric path loss and shadowing, and is included in the spatial correlation matrices. The received signal at AP $r$ in the presence of the target and for $m\in\{1,\cdots,L_{d}\}$, can be formulated as
$\textbf{y}_{r}[m]=\sum_{k=1}^{N_{\rm tx}}\underbrace{\alpha_{r,k}\sqrt{\beta_{r,k}}\textbf{a}(\phi_{r},\theta_{r})\textbf{a}^{T}(\varphi_{k},\vartheta_{k})\textbf{x}_{k}[m]}_{\textrm{desired reflections from the target}}$
$+\sum_{k=1}^{N_{\rm tx}}\underbrace{\textbf{H}_{r,k}\textbf{x}_{k}[m]}_{\textrm{undesired signals}}+\textbf{n}_{r}[m],$ (15)
[…]
Following the same notation as in [25], in (II-C), the known part of each reflected path is denoted by ${\textbf{g}_{r,k}[m]\in\mathbb{C}^{M}}$, defined as
$\textbf{g}_{r,k}[m]\triangleq\sqrt{\beta_{r,k}}\textbf{a}(\phi_{r},\theta_{r})\textbf{a}^{T}(\varphi_{k},\vartheta_{k})\textbf{x}_{k}[m],$ (16)
where the matrix $\alpha_{r,k}\sqrt{\beta_{r,k}}\textbf{a}(\phi_{r},\theta_{r})\textbf{a}^{T}(\varphi_{k},\vartheta_{k})$ represents the reflected path through the target. Here, $\phi_{r}$ and $\theta_{r}$ denote the azimuth and elevation angles from the target location to receiver AP $r$, respectively. Similarly, $\varphi_{k}$ and $\vartheta_{k}$ represent the azimuth and elevation angles from transmit AP $k$ to the target location.
Each receive AP sends their respective signals $\textbf{y}_{r}[m]$, for $r=1,\ldots,N_{\rm rx}$, to the edge cloud, to form the collective received signal $\textbf{y}[m]=\begin{bmatrix}\textbf{y}_{1}^{T}[m]&\cdots&\textbf{y}_{N_{\rm rx}}^{T}[m]\end{bmatrix}^{T}$, which can be expressed as
$\textbf{y}[m]$ $=\underbrace{\mathrm{blkdiag}\left(\textbf{G}_{1}[m],\ldots,\textbf{G}_{N_{\rm rx}}[m]\right)}_{\triangleq\textbf{G}[m]}\boldsymbol{\alpha}$
$+\underbrace{\left(\textbf{I}_{N_{\rm rx}}\otimes\left(\textbf{x}^{T}[m]\otimes\textbf{I}_{M}\right)\right)}_{\triangleq\textbf{X}[m]}\boldsymbol{\mathfrak{h}}+\textbf{n}[m],$ (17)
where $\textbf{G}_{r}[m]=\begin{bmatrix}\textbf{g}_{r,1}[m]&\cdots&\textbf{g}_{r,N_{\rm tx}}[m]\end{bmatrix}\in\mathbb{C}^{M\times N_{\rm tx}}$, $\textbf{x}[m]=\begin{bmatrix}\textbf{x}^{T}_{1}[m]&\cdots&\textbf{x}^{T}_{N_{\rm tx}}[m]\end{bmatrix}^{T}\in\mathbb{C}^{N_{\rm tx}M}$, and $\mathfrak{h}\sim\mathcal{CN}\left(\textbf{0},\textbf{R}\right)$ is the vectorized target-free channels[25, Sec. V].
II-D Transmit ISAC Precoding Vectors
The communication and sensing transmit precoding vectors are obtained based on regularized zero forcing (RZF) and zero forcing (ZF) approaches, respectively. The unit-norm RZF precoding vector for UE $i$ is given as $\textbf{w}_{i}=\frac{\bar{\textbf{w}}_{i}}{\left\|\bar{\textbf{w}}_{i}\right\|}$, with
$\bar{\textbf{w}}_{i}\!=\!\left(\sum\limits_{j=1}^{N_{\rm ue}}\hat{\textbf{h}}_{j}\hat{\textbf{h}}_{j}^{H}+\delta\textbf{I}_{N_{\rm tx}M}\right)^{\!\!-1}\!\!\hat{\textbf{h}}_{i},\quad i=1,\ldots,N_{\rm ue},$ (18)
where $\delta$ is the regularization parameter, and $\hat{\textbf{h}}_{j}=\begin{bmatrix}\hat{\textbf{h}}_{j,1}^{T}&\ldots&\hat{\textbf{h}}_{j,N_{\rm tx}}^{T}\end{bmatrix}^{T}\in\mathbb{C}^{N_{\rm tx}M}$ is the LMMSE channel estimate of the communication channel $\textbf{h}_{j}$.
We aim to null the destructive interference from the sensing signal to the UEs by using the unit-norm ZF sensing precoding vector $\textbf{w}_{0}=\frac{\bar{\textbf{w}}_{0}}{\left\|\bar{\textbf{w}}_{0}\right\|}$, where
$\bar{\textbf{w}}_{0}=\left(\textbf{I}_{N_{\rm tx}M}-\textbf{U}\textbf{U}^{H}\right)\textbf{h}_{0},$ (19)
and U is the unitary matrix with the orthogonal columns that span the column space of the matrix $\begin{bmatrix}\hat{\textbf{h}}_{1}&\ldots&\hat{\textbf{h}}_{N_{\rm ue}}\end{bmatrix}$. $\textbf{h}_{0}=\begin{bmatrix}\sqrt{\beta_{1}}\textbf{a}^{T}(\varphi_{1},\vartheta_{1})&\ldots&\sqrt{\beta_{N_{\rm tx}}}\textbf{a}^{T}(\varphi_{N_{\rm tx}},\vartheta_{N_{\rm tx}})\end{bmatrix}^{T}\in\mathbb{C}^{N_{\rm tx}M}$ is the concatenated sensing channel between all the ISAC APs and the target, including corresponding path loss and the array response vectors $\textbf{a}(\varphi_{k},\vartheta_{k})$ for $k=1,\ldots,N_{\rm tx}$.
III Reliability and Delay Analysis for URLLC
In this section, we derive an upper bound on the DEP and the transmission delay, both crucial aspects considered as URLLC requirements. In the finite blocklength regime, the communication data cannot be transmitted without error. From [16], ergodic data rate of UE $i$ can be approximated as
$R_{i}\!\approx\!\mathbb{E}\Bigg{\{}\!\!\left(1\!-\!\beta\right)\!\log_{2}{\left(1+\mathsf{SINR}^{\rm(dl)}_{i}\right)}\!-\!\frac{Q^{-1}(\epsilon_{i})}{\ln(2)}\!\sqrt{\!\frac{(1\!-\!\beta)V_{i}}{L}}\Bigg{\}},\vspace{-0mm}$ (20)
where $\beta=\frac{L_{p}}{L}$, $\epsilon_{i}$ denotes the DEP when transmitting $b_{i}$ bits to UE $i$, $\mathsf{SINR}^{\rm(dl)}_{i}$ is the instantaneous downlink communication signal-to-interference-plus-noise ratio (SINR) for UE $i$, $V_{i}=1-\left(1+\mathsf{SINR}^{\rm(dl)}_{i}\right)^{-2}$ is the channel dispersion, and $Q(\cdot)$ refers to the Gaussian Q-function. Due to the fact that $V_{i}<1$, the ergodic data rate can be lower bounded by
$R_{i}$ $\!\geq\!\left(1\!-\!\beta\right)\!\mathbb{E}\Bigg{\{}\!\!\log_{2}{\left(1+\mathsf{SINR}^{(\rm dl)}_{i}\right)}\!\!\Bigg{\}}\!-\!\frac{Q^{-1}(\epsilon_{i})}{\ln(2)}\sqrt{\frac{(1-\beta)}{L}}.$ (21)
Moreover, given that only $\mathbb{E}\left\{\textbf{h}_{i}^{H}\textbf{w}_{i}\right\}$ is known at UE $i$, and according to [27, Thm. 6.1] and [25, Lem. 1],
$\mathbb{E}\left\{\log_{2}{\left(1+\mathsf{SINR}^{\rm(dl)}_{i}\right)}\right\}\geq\log_{2}{\left(1+\overline{\mathsf{SINR}}^{(\rm dl)}_{i}\right)}$ (22)
where
$\overline{\mathsf{SINR}}^{(\rm dl)}_{i}=\frac{\rho_{i}\mathsf{b}_{i}^{2}}{\sum_{j=0}^{{\rm N_{ue}}}\rho_{j}\mathsf{a}_{i,j}^{2}+\sigma_{n}^{2}},\quad i=1,\ldots,N_{\rm ue}$ (23)
with
$\mathsf{b}_{i}=\left|\mathbb{E}\left\{\textbf{h}_{i}^{H}\textbf{w}_{i}\right\}\right|,\quad\mathsf{a}_{i,i}=\sqrt{\mathbb{E}\left\{\left|\textbf{h}_{i}^{H}\textbf{w}_{i}\right|^{2}\right\}-\mathsf{b}_{i}^{2}}$ (24)
$\mathsf{a}_{i,j}=\sqrt{\mathbb{E}\left\{\left|\textbf{h}_{i}^{H}\textbf{w}_{j}\right|^{2}\right\}},\quad j=0,1,\ldots,N_{\rm ue},\quad j\neq i.$ (25)
The expectations are taken with respect to the random channel realizations. Now, using (22) and substituting $R_{i}=\frac{b_{i}}{L}$ into (21), we obtain an upper bound for the DEP as
$\epsilon_{i}\leq\epsilon_{i}^{\rm(ub)}\!\triangleq\!Q\!\left(\sqrt{L-L_{p}}\left[\ln\left(1+\overline{\mathsf{SINR}}^{(\rm dl)}_{i}\right)-\frac{b_{i}\ln{2}}{L-L_{p}}\right]\right).$ (26)
In this paper, we focus on the transmission delay and leave the analysis of E2E delay as future work. Let $D_{i}^{t}$ denote the transmission delay of UE $i$, expressed as
$D_{i}^{\rm t}=\frac{T}{1-\epsilon_{i}}=\frac{L}{B(1-\epsilon_{i})},\quad i=1,\ldots,N_{\rm ue}$ (27)
where $T=\frac{L}{B}$ is the time duration of one URLLC transmission with blocklength $L$. To satisfy the reliability requirement, $\epsilon_{i}$ should be less than the maximum tolerable DEP denoted by $\epsilon_{i}^{(\rm th)}$. Then, since $\epsilon_{i}\leq\epsilon_{i}^{\rm(ub)}\leq\epsilon_{i}^{\rm(th)}$, the transmission delay is upper-bounded as
$D_{i}^{\rm t}\leq\frac{L}{B\left(1-\epsilon_{i}^{\rm(th)}\right)}\triangleq D_{i}^{\rm(ub)}\leq D_{i}^{\rm(th)},$ (28)
where $D_{i}^{\rm(th)}$ is the maximum tolerable delay by UE $i$ and $D_{i}^{\rm(ub)}\leq D_{i}^{\rm(th)}$ should be satisfied to guarantee the delay requirement. This implies that the blocklength cannot exceed $D^{\rm(th)}_{i}B(1-\epsilon_{i}^{\rm(th)})$. Thus, we can define the maximum tolerable blocklength by $L_{\rm max}$, where
$L_{\rm max}=\min\left\{D^{\rm(th)}_{i}B\left(1-\epsilon_{i}^{\rm(th)}\right)|i=1,\cdots,N_{\rm ue}\right\}.$ (29)
IV Sensing Analysis
We have numerically shown in [19] and [25], that maximizing the sensing SINR improves the target detection probability under a constant false alarm probability. When studying other sensing tasks, it is naturally desired to keep the sensing SINR above a required threshold denoted by $\gamma_{\rm s}$, i.e., $\mathsf{SINR}_{s}\geq\gamma_{\rm s}$. This fact motivates the optimization problems presented in Section VI.
Using the previously defined vector $\boldsymbol{\rho}=[\sqrt{\rho_{0}}\ \ldots\ \sqrt{\rho_{N_{\rm ue}}}]^{T}$, from [25], the sensing SINR with independent RCS values is given by
$\mathsf{SINR_{\rm s}}=\frac{\boldsymbol{\rho}^{T}\textbf{A}\boldsymbol{\rho}}{L_{d}MN_{\rm rx}\sigma_{n}^{2}+\boldsymbol{\rho}^{T}\textbf{B}\boldsymbol{\rho}},$ (30)
where
[…]
×$\hskip 8.53581pt\times\textbf{W}_{k}\Bigg{)}\textbf{D}_{\rm s}[m],$ (31)
B $=\sum_{m=1}^{L_{d}}\textbf{D}_{\rm s}^{H}[m]\left(\sum_{r=1}^{N_{\rm rx}}\sum_{k=1}^{N_{\rm tx}}\mathrm{tr}\left(\textbf{R}_{\mathrm{rx},(r,k)}\right)\textbf{W}_{k}^{H}\textbf{R}^{T}_{{\rm tx},(r,k)}\textbf{W}_{k}\right)$
×$\hskip 8.53581pt\times\textbf{D}_{\rm s}[m].$ (32)
As we see, the sensing SINR is a function of symbols which are random for different blocklength values. When the blocklength is fixed, we can assume that the symbols are known in advance. This implies that although the symbols are known when processing the received sensing signals, they cannot be known during the resource allocation phase. To this end, we take the expectation with respect to the random symbols. Then, the average sensing SINR would be
$\overline{\mathsf{SINR}}_{s}$ $=\frac{\boldsymbol{\rho}^{T}\mathbb{E}\{\textbf{A}\}\boldsymbol{\rho}}{L_{d}MN_{\rm rx}\sigma_{n}^{2}+\boldsymbol{\rho}^{T}\mathbb{E}\{\textbf{B}\}\boldsymbol{\rho}},$ (33)
The symbols $s_{i}[m]$ are independent, having zero mean and unit variance. Therefore, $\mathbb{E}\{s_{i}^{*}[m]s_{i}[m]\}=1$ and $\mathbb{E}\{s_{i}^{*}[m]s_{j}[m]\}=0$ for $i\neq j$. Then, the expectations in (33) are obtained as follows
$\mathbb{E}\{\textbf{A}\}=L_{d}\,M\textbf{A}_{D},\quad$ $\mathbb{E}\{\textbf{B}\}=L_{d}\,\textbf{B}_{D},$ (34)
where
$\textbf{A}_{D}=\mathrm{diag}\Bigg{(}\mathrm{diag}\Bigg{(}$ $\sum_{r=1}^{N_{\rm rx}}\sum_{k=1}^{N_{\rm tx}}\beta_{r,k}\textbf{W}_{k}^{H}\textbf{a}^{*}(\varphi_{k},\vartheta_{k})$
×$\times\textbf{a}^{T}(\varphi_{k},\vartheta_{k})\textbf{W}_{k}\Bigg{)}\Bigg{)}$ (35)
$\textbf{B}_{D}=\mathrm{diag}\Bigg{(}\mathrm{diag}\Bigg{(}$ $\sum_{r=1}^{N_{\rm rx}}\sum_{k=1}^{N_{\rm tx}}\mathrm{tr}\left(\textbf{R}_{\mathrm{rx},(r,k)}\right)\textbf{W}_{k}^{H}\textbf{R}^{T}_{{\rm tx},(r,k)}\textbf{W}_{k}\Bigg{)}\Bigg{)}.$ (36)
The average sensing SINR with respect to symbols and RCS values is finally written as
$\overline{\mathsf{SINR}}_{s}$ $=\frac{M\boldsymbol{\rho}^{T}\textbf{A}_{D}\boldsymbol{\rho}}{MN_{\rm rx}\sigma_{n}^{2}+\boldsymbol{\rho}^{T}\textbf{B}_{D}\boldsymbol{\rho}}.$ (37)
V E2E Power Consumption Modeling
Compared to communication networks, ISAC networks are expected to consume more power due to sensing tasks. In general, E2E power consumption in a system with C-RAN architecture is consisting of two main components: i) the radio site power consumption, including the AP power consumption and ii) the power consumption at the edge cloud, denoted by $P_{\rm cloud}$ [23].
In the considered C-RAN architecture, all processing is done in the cloud. Let $P_{\textrm{AP},0}^{\rm tx}$ and $P_{\textrm{AP},0}^{\rm rx}$ be the static power consumption of the transmit ISAC AP and the receive sensing AP, respectively. The total power consumption, taking into account both communication and sensing, can be expressed as
$P_{\rm total}$ $=\underbrace{\sum_{k=1}^{N_{\rm tx}}\left(P_{\textrm{AP},0}^{\rm tx}+\Delta^{\rm tr}P_{k}\right)+\sum_{r=1}^{N_{\rm rx}}P_{\textrm{AP},0}^{\rm rx}}_{\textrm{radio site power consumption}}+P_{\rm cloud}$ (38)
where $\Delta^{\rm tr}$ is the slope of load-dependent transmission power consumption of each AP. The $P_{k}$ is the transmission power of the ISAC AP $k$ from (8). The power consumption at the cloud is modeled as
$P_{\rm cloud}$ $=P_{\rm fixed}+\frac{1}{\sigma_{\rm cool}}\Bigg{(}P_{\rm cloud,0}^{\rm proc}+\Delta_{\rm cloud}^{\rm proc}\frac{C_{\rm cloud}}{C_{\rm max}}\Bigg{)},$ (39)
where $P_{\rm fixed}$ is the fixed power consumption at the cloud which is independent of the load. $P_{\rm cloud,0}^{\rm proc}$ is the processing power consumption in the idle mode. $\sigma_{\rm cool}\in(0,1]$ and $\Delta_{\rm cloud}^{\rm proc}$ denote the cooling efficiency of the cloud and the slope of the load-dependent power consumption for processing at the digital unit (DU) in the cloud, respectively. Moreover, $C_{\rm max}$ and $C_{\rm cloud}\in[0,C_{\rm max}]$ are the maximum processing capacity of the processing resources and the total processing resource utilization in giga-operations per second (GOPS), respectively[14, 31]. The processing resource utilization can be expressed as
$C_{\rm cloud}=C_{\rm proc}^{\rm c}+C_{\rm proc}^{\rm s}$ (40)
where $C_{\rm proc}^{\rm c}$ and $C_{\rm proc}^{\rm s}$ are the processing resource utilization due to communication and sensing tasks, respectively [^1] In this paper, we focus on the GOPS analysis by taking into account only physical-layer communication and sensing processing and neglect high-layer operations..
In the following parts, the required GOPS for communication and sensing in our system model is computed, respectively. We assess the computational complexity where only the numbers of real multiplications and divisions are counted. Each complex multiplication is equal to four real multiplications. We also consider memory overhead in arithmetic operation calculations by multiplying each operation by two as done in [32, 14]. Hence, each complex multiplication is counted as $4\cdot 2=8$ operations in computing the total GOPS.
V-A GOPS Analysis of Digital Operations for Communication at the Cloud
In this section, we analyze the GOPS for digital signal processing corresponding to the communication tasks including the uplink channel estimation and downlink transmission. To compute the number of real multiplications, we mainly follow the GOPS analysis in [28, App. B], [23].
Let $C_{\textrm{ch-est}}$ denote the computational complexity of the LMMSE channel estimation approach for all the APs. To compute the channel estimates, we first obtain the vectors $\textbf{y}_{t_{i},k}^{p}$. From [28, App. B], the multiplication of one matrix of size $M\times L_{p}$ with a vector of size $L_{p}\times 1$ results in $ML_{p}$ complex multiplications. Hence, obtaining $\textbf{y}_{t_{i},k}^{p}$ for $N_{\rm ue}$ UEs at all $N_{\rm tx}$ APs costs $8ML_{p}N_{\rm ue}N_{\rm tx}$ real multiplications/division in total, if $L_{p}\geq N_{\rm ue}$. Otherwise, the number of real multiplications/divisions would be $8ML_{p}^{2}N_{\rm tx}$. Moreover, we need to compute the matrices $\textbf{R}_{i,k}^{\prime}$ and $\boldsymbol{\Psi}_{t_{i},k}^{\prime}$, given that $\bar{\textbf{h}}_{i,k}$ and $\textbf{R}_{i,k}$ are known. However, this pre-computation can be neglected since the channel statistics are usually constant for a while and there is no need to compute them every coherence block. Finally, the multiplications in (11) costs $8M^{2}N_{\rm ue}N_{\rm tx}$ real operations. Then, $C_{\textrm{ch-est}}$ is equal to
$C_{\textrm{ch-est}}=\left\{\begin{matrix}(8ML_{p}+8M^{2})N_{\rm ue}N_{\rm tx},&L_{p}\geq N_{\rm ue}\\ 8ML_{p}^{2}N_{\rm tx}+8M^{2}N_{\rm ue}N_{\rm tx},&L_{p}<N_{\rm ue}.&\end{matrix}\right.$ (43)
The number of real multiplications/divisions to compute centralized RZF precoding vector for all the UEs from [28, App. B] is
$C_{\textrm{prec-comp}}$ $=\underbrace{\frac{8\left((MN_{\rm tx})^{2}+MN_{\rm tx}\right)}{2}N_{\rm ue}}_{\textrm{inside the parentheses}}$
$\quad+\underbrace{\frac{8\left((MN_{\rm tx})^{3}-MN_{\rm tx}\right)}{3}}_{\rm inversion}+N_{\rm ue}\cdot\underbrace{8(MN_{\rm tx})^{2}}_{\textrm{for}(.)^{-1}\hat{\textbf{h}}_{i}}$
$\quad+N_{\rm ue}\cdot\left(\underbrace{8MN_{\rm tx}}_{\textrm{computing }||\textbf{w}_{i}||}+\underbrace{4MN_{\rm tx}}_{\textrm{normalization}}\right)$
$=\left(12(MN_{\rm tx})^{2}+16MN_{\rm tx}\right)N_{\rm ue}$
$\quad+\frac{8\left((MN_{\rm tx})^{3}-MN_{\rm tx}\right)}{3}.$ (44)
Reciprocity calibration and multiplication of the symbols by the precoding vectors, each costs $8L_{d}MN_{\rm ue}N_{\rm tx}$ real operations [33, 14]. Multiplying by the power coefficients also costs $4L_{d}MN_{\rm ue}N_{\rm tx}$. Finally, the GOPS corresponding to communication processing (i.e., channel estimation, precoding and reciprocity calibration) is computed as
$C_{\textrm{proc}}^{\rm c}=\frac{B}{L10^{9}}\Bigg{(}C_{\textrm{ch-est}}+C_{\textrm{prec-comp}}+20L_{d}MN_{\rm ue}N_{\rm tx}\Bigg{)},$ (45)
where we divided the total giga operations by the coherence time $L/B$.
V-B GOPS Analysis of Digital Operations for Sensing at the Cloud
In this subsection, we compute the GOPS for sensing operations. In downlink, first the sensing precoding vector $\textbf{w}_{0}$ should be obtained. Next operation is multiplying the sensing symbols by the sensing precoding vector and the sensing power coefficient which costs $12L_{d}MN_{\rm tx}$ real multiplications/divisions. Let $C_{\rm prec-comp}^{\rm s}$ be the number of real multiplications/divisions to compute the sensing precoder $\textbf{w}_{0}$. The unitary matrix U in (19) (subspace spanned by the UE channel estimation vectors) is already obtained when computing the RZF precoding vectors by matrix inversion and the corresponding $\mathbf{L}\mathbf{D}\mathbf{L}^{H}$ decomposition [28, Lem. B.2]. Therefore, the computational complexity of the ZF precoding vector $\textbf{w}_{0}$ is given by
$C_{\rm prec-comp}^{\rm s}$ $=8(MN_{\rm tx})^{2}+12MN_{\rm tx},$ (46)
where the first term stands for matrix-vector multiplication and the second term corresponds to the cost of computing $||\textbf{w}_{0}||$ and normalization, which are counted as $8MN_{\rm tx}+4MN_{\rm tx}$.
After transmitting the signal in downlink, the reflected signals along with the interference signals are received at the receiver APs and sent to the cloud. At the cloud, these signals are processed for a specific sensing application. In this work we consider target detection and assess the number of real multiplications/divisions required to compute the test statistics. For target detection problems, we usually compute the test statistics and compare them with a threshold. The target is declared detected if the value of the test statistics is greater than the threshold. We assume that the threshold is constant. Therefore, we can neglect the computational complexity of obtaining the threshold. However, test statistics should be obtained for each transmission. We use the maximum a posteriori ratio test (MAPRT) detector proposed in [25, Lem. 2]. The test statistics is given by
$T$ $=\frac{1}{\sigma_{n}^{2}}\left(\begin{bmatrix}\textbf{a}^{H}&\textbf{b}^{H}\end{bmatrix}\left(\begin{bmatrix}\textbf{C}&\textbf{E}\\ \textbf{E}^{H}&\textbf{D}\end{bmatrix}^{-1}-\begin{bmatrix}\textbf{0}&\textbf{0}\\ \textbf{0}&\textbf{D}^{-1}\end{bmatrix}\right)\begin{bmatrix}\textbf{a}\\ \textbf{b}\end{bmatrix}\right),$ (54)
where
$\textbf{a}=\sum_{m=1}^{L_{d}}\textbf{G}^{H}[m]\textbf{y}[m],\quad\textbf{b}=\sum_{m=1}^{L_{d}}\textbf{X}^{H}[m]\textbf{y}[m],$ (55)
$\textbf{C}=\sum_{m=1}^{L_{d}}\textbf{G}^{H}[m]\textbf{G}[m]+\sigma_{n}^{2}\textbf{I}_{N_{\rm rx}N_{\rm tx}},$ (56)
$\textbf{D}=\sum_{m=1}^{L_{d}}\textbf{X}^{H}[m]\textbf{X}[m]+\sigma_{n}^{2}\textbf{R}^{-1},$ (57)
$\textbf{E}=\sum_{m=1}^{L_{d}}\textbf{G}^{H}[m]\textbf{X}[m].$ (58)
Let $C_{\rm se-comp}$ denote the number of real multiplications/divisions to compute the test statistics $T$. The computational complexity can be expressed as
$C_{\rm se-comp}=C_{\rm se-prep}+C_{\rm se-detector}$ (59)
where $C_{\rm se-prep}$ and $C_{\rm se-detector}$ are the computational complexity of pre-processing and computing the test statistics for sensing, respectively. The former corresponds to computing (55)-(58), and the latter corresponds to computing the test statistics in (54). The computational complexity for pre-processing part and each step of computing the test statistics are listed in Table I. The total $C_{\rm se-prep}$ and $C_{\rm se-detector}$ are obtained as
$C_{\rm se-prep}=L_{d}C_{{\rm se-prep},m}$
$L_{d}\Bigg{(}20MN_{\rm tx}N_{\rm rx}+16M^{2}N_{\rm tx}N_{\rm rx}+4N_{\rm rx}\left(N_{\rm tx}^{2}+N_{\rm tx}\right)M$
$\quad+4\left((MN_{\rm tx})^{2}+MN_{\rm tx}\right)+8N_{\rm rx}\,N_{\rm tx}^{2}\,M^{2}\Bigg{)}$ (60)
$C_{\rm se-detector}=8\frac{\left((1+M^{2})N_{\rm tx}N_{\rm rx}\right)^{3}-(1+M^{2})N_{\rm tx}N_{\rm rx}}{3}$
$\quad+8\frac{(M^{2}N_{\rm tx}N_{\rm rx})^{3}-M^{2}N_{\rm tx}N_{\rm rx}}{3}$
$\quad+8\,\left(\left((1+M^{2})N_{\rm tx}N_{\rm rx}\right)^{2}+(1+M^{2})N_{\rm tx}N_{\rm rx}\right),$ (61)
where $C_{{\rm se-prep},m}$ is defined as $C_{\rm se-prep}/L_{d}$. Finally, the sensing GOPS is obtained as
$C_{\rm proc}^{\rm s}$ $=\frac{B}{L10^{9}}\Bigg{(}\underbrace{12L_{d}MN_{\rm tx}}_{\textrm{precoding + power}}+C_{\textrm{prec-comp}}^{\rm s}+\underbrace{C_{\rm se-comp}}_{\textrm{target detection}}\Bigg{)}.$ (62)
TABLE I: Number of real multiplications/divisions for computing test statistics.
Operation # of real multiplications/division
$\textbf{G}[m],\quad\forall m$ $20\,L_{d}\,MN_{\rm tx}N_{\rm rx}$
a $8\,L_{d}\,M^{2}\,N_{\rm rx}\,N_{\rm tx}\,$
b $8\,L_{d}\,M^{2}\,N_{\rm tx}\,N_{\rm rx}$
C $4\,L_{d}N_{\rm rx}\left(N_{\rm tx}^{2}+N_{\rm tx}\right)M$
D $4L_{d}\,\left((MN_{\rm tx})^{2}+MN_{\rm tx}\right)$
E $8\,L_{d}\,N_{\rm rx}\,N_{\rm tx}^{2}\,M^{2}$
$\begin{bmatrix}\textbf{C}&\textbf{E}\\ \textbf{E}^{H}&\textbf{D}\end{bmatrix}^{-1}$ $8\frac{\left((1+M^{2})N_{\rm tx}N_{\rm rx}\right)^{3}-(1+M^{2})N_{\rm tx}N_{\rm rx}}{3}$
$\textbf{D}^{-1}$ $8\frac{(M^{2}N_{\rm tx}N_{\rm rx})^{3}-M^{2}N_{\rm tx}N_{\rm rx}}{3}$
$\begin{bmatrix}\textbf{a}^{H}&\textbf{b}^{H}\end{bmatrix}\begin{bmatrix}(.)&(.)\\ (.)&(.)\end{bmatrix}\begin{bmatrix}\textbf{a}\\ \textbf{b}\end{bmatrix}$ $8\,\left(\left((1+M^{2})N_{\rm tx}N_{\rm rx}\right)^{2}+(1+M^{2})N_{\rm tx}N_{\rm rx}\right)$
VI Joint blocklength and Power Optimization
In this paper, we aim to optimize the blocklength and the power control coefficients to minimize the energy consumption while URLLC and sensing requirements are satisfied. In the following subsections, we cast two non-convex optimization problems: i) minimizing only the energy consumption for transmission; ii) minimizing E2E energy consumption by taking into account the energy consumption at the radio site and the cloud. [^2] In addition to considering radio and cloud processing energy consumption, it is important to take fronthaul energy consumption into account when assessing E2E energy consumption. However, in our analysis, where the primary focus is on the interplay between processing and radio resources, we treat fronthaul energy consumption as a fixed component and do not include it in our considerations.
VI-A Minimizing Energy Consumption for Transmission
The optimization problem can be cast as
[…]
subject to $\epsilon_{i}\leq\epsilon^{\rm(ub)}_{i}\leq\epsilon^{\rm(th)}_{i},\quad\forall i>0$ (63b)
$L\leq L_{\rm max},\quad$ (63c)
$\overline{\mathsf{SINR}}_{s}\geq\gamma_{\rm s},\quad$ (63d)
$P_{k}\leq P_{\rm tx},\quad\quad\textrm{for}\quad k=1,\cdots,N_{\rm tx}$ (63e)
$P_{\rm tr}=\sum_{k=1}^{N_{\rm tx}}P_{k}$ is the average power consumption for transmission, given as $P_{\rm tr}=\sum_{j=0}^{N_{\rm ue}}\rho_{j}=\|\boldsymbol{\rho}\|^{2}$, where $\boldsymbol{\rho}=[\sqrt{\rho_{0}}\ \ldots\ \sqrt{\rho_{N_{\rm ue}}}]^{T}$, due to the unit-power centralized precoding vectors. Constraints (63b) and (63c) correspond to the URLLC requirements. The $\gamma_{\rm s}$ is the required sensing SINR that is selected according to the target detection performance requirement and $P_{\rm tx}$ is the maximum transmit power per AP.
The aforementioned problem is challenging to solve due to its non-convex nature and the high coupling of variables. In the following theorem, we present an equivalent optimization problem by introducing newly defined auxiliary variables. This allows us to obtain a more tractable optimization problem.
Theorem 1.
Consider the problem in (VI-A) and define the collective vectors $\textbf{r}=[r_{1}\ \ldots\ r_{N_{\rm ue}}]^{T}$ and $\boldsymbol{\chi}=[\chi_{1}\ \ldots\ \chi_{N_{\rm ue}}]^{T}$. The optimal solution $\{\boldsymbol{\rho}^{\star},L^{\star}\}$ of the problem given below is also an optimal solution to (VI-A):
$\underset{\boldsymbol{\rho},\boldsymbol{\chi},\textbf{r}\geq\textbf{0},L>L_{p},\bar{L}>0}{\textrm{minimize}}\quad\frac{\|\boldsymbol{\rho}\|^{2}}{\overline{L}}\,$ (64a)
[…]
$\ln\left(1+\chi_{i}\right)-\chi_{i}+\frac{\rho_{i}\mathsf{b}_{i}^{2}}{r_{i}}\geq\frac{Q^{-1}\left(\epsilon_{i}^{\rm(th)}\right)}{\sqrt{L-L_{p}}}+\frac{b_{i}\ln{2}}{L-L_{p}}$ (64c)
$\left\|\begin{bmatrix}\sqrt{2}\mathsf{a}_{i,0}\sqrt{\rho_{0}}\\ \vdots\\ \sqrt{2}\mathsf{a}_{i,N_{\rm ue}}\sqrt{\rho_{N_{\rm ue}}}\\ \sqrt{2}\mathsf{b}_{i}\sqrt{\rho_{i}}\\ \sqrt{2}\sigma_{n}\\ 1+\chi_{i}\\ r_{i}\end{bmatrix}\!\right\|\leq 1+\chi_{i}+r_{i},\forall i,$ (64k)
$\boldsymbol{\rho}^{T}\left(\gamma_{\rm s}\textbf{B}_{D}-M\,\textbf{A}_{D}\,\right)\boldsymbol{\rho}\leq-\gamma_{\rm s}MN_{\rm rx}\sigma_{n}^{2}$ (64l)
$\left\|\textbf{F}_{k}\boldsymbol{\rho}\right\|\leq\sqrt{P_{\rm tx}},\quad k=1,\ldots,N_{\rm tx},$ (64m)
$L\leq L_{\rm max},$ (64n)
where ${\bf F}_{k}=\mathrm{diag}\left(\sqrt{\mathbb{E}\left\{\|{\bf w}_{0,k}\|^{2}\right\}},\ldots,\sqrt{\mathbb{E}\left\{\|{\bf w}_{N_{\rm ue},k}\|^{2}\right\}}\right)$.
Proof.
See Appendix A. ∎
The optimization problem in (1) is still not convex due to the non-convex constraints (64b), (64c) and (64l). The terms that destroy convexity are the convex terms $1/\overline{L}$ and $\frac{\rho_{i}\mathsf{b}_{i}^{2}}{r_{i}}$ (in terms of $\boldsymbol{\rho}$ and r) on the right-hand side of (64b) and the left-hand side of (64c), respectively. To this end, we apply the concave-convex procedure (CCP) approach to (64b) and (64c), and the FPP-SCA method [34] to (64l), wherein $-\boldsymbol{\rho}^{T}M\,\textbf{A}_{D}\boldsymbol{\rho}$ is a concave function. Moreover, to avoid any potential infeasibility issue regarding (64l) during the initial iterations of the algorithm, we add slack variable $\chi_{0}\geq 0$ and a slack penalty $\lambda$, to the convexified problem at the initial iterations. In subsequent iterations, we set $\chi_{0}$ to zero if it is less than a threshold, denoted as $\chi_{0}\leq\epsilon_{\chi}$. Finally, the convex problem that is solved at the $c^{\rm th}$ iteration becomes
$\underset{\boldsymbol{\rho},\boldsymbol{\chi},\textbf{r}\geq\textbf{0},L,\bar{L}>L_{p},\chi_{0}\geq 0}{\textrm{minimize}}\quad\frac{\|\boldsymbol{\rho}\|^{2}}{\overline{L}}+\lambda\chi_{0}\,,$ (65a)
subject $\textrm{subject to}\quad{L-L_{p}}\leq\frac{2}{\overline{L}^{(c-1)}}-\frac{\overline{L}}{\left(\overline{L}^{(c-1)}\right)^{2}},$ (65b)
$\ln\left(1+\chi_{i}\right)-\chi_{i}+2\frac{\sqrt{\rho_{i}}^{(c-1)}\mathsf{b}_{i}^{2}\,\sqrt{\rho_{i}}}{r_{i}^{(c-1)}}-r_{i}\left(\frac{\sqrt{\rho_{i}}^{(c-1)}\mathsf{b}_{i}}{r_{i}^{(c-1)}}\right)^{2}$
$\hskip 14.22636pt\geq\frac{Q^{-1}\left(\epsilon_{i}^{\rm(th)}\right)}{\sqrt{L-L_{p}}}+\frac{b_{i}\ln{2}}{L-L_{p}}$ (65c)
$\gamma_{\rm s}\boldsymbol{\rho}^{T}\textbf{B}_{D}\boldsymbol{\rho}-2M\,\Re\left(\left(\boldsymbol{\rho}^{(c-1)}\right)^{T}\textbf{A}_{D}\boldsymbol{\rho}\right)$
$\hskip 14.22636pt\leq-\gamma_{\rm s}MN_{\rm rx}\sigma_{n}^{2}-M\left(\boldsymbol{\rho}^{(c-1)}\right)^{T}\textbf{A}_{D}\boldsymbol{\rho}^{(c-1)}+\chi_{0},$ (65d)
[…]
The steps of FPP-SCA and CCP procedure are outlined in Algorithm 1. We empirically observed that setting $\overline{L}^{(0)}=\frac{1}{L_{\rm max}-L_{p}}$, $\rho_{0}^{(0)}=0$ and $\sqrt{\rho_{i}}^{(0)}=10^{-3}\sqrt{P_{\rm tx}/N_{\rm ue}}$ for $i>0$ yields satisfactory results.
Algorithm 1 FPP-SCA and CCP Procedure for Problem (1).
1:Initialization: Set arbitrary $\boldsymbol{\rho}^{(0)},\textbf{r}^{(0)}\geq\textbf{0}$, the solution accuracy parameters $\epsilon,\epsilon_{\chi}>0$, $\overline{L}^{(0)}=\frac{1}{L_{\rm max}-L_{p}}$, and $\lambda>0$. Set the iteration counter to $c=0$, and a maximum iteration number $c_{\rm max}$. Moreover, let $\mathsf{F}^{(0)}=\infty$ be the initial value of objective function and $\Delta\mathsf{F}^{(c)}=\mathsf{F}^{(c)}-\mathsf{F}^{(c-1)}$ with $\mathsf{F}^{(-1)}=0$.
2:$c\leftarrow c+1$.
3:While $\Delta\mathsf{F}^{(c-1)}\geq\epsilon$ and $c\leq c_{\rm max}$, do
4:Set $\boldsymbol{\rho}^{(c)}$, $\overline{L}^{(c)}$, and $\textbf{r}^{(c)}$ to the solution of the convex problem in (VI-A), where the previous iterates $\boldsymbol{\rho}^{(c-1)}$, $\overline{L}^{(c-1)}$, and $\textbf{r}^{(c-1)}$ are taken as constant.
5:If $\chi_{0}<\epsilon_{\chi}$, set $\chi_{0}=0$ for the next iteration.
6:End while
7:Output: The transmit power coefficients $\boldsymbol{\rho}^{(c)}$ and the blocklength $L^{(c)}$.
VI-B Minimizing E2E Energy Consumption
In this subsection, our objective is to minimize the E2E energy consumption. The optimization problem is formulated as follows
[…]
subject to $\epsilon_{i}\leq\epsilon^{\rm(ub)}_{i}\leq\epsilon^{\rm(th)}_{i},\quad\forall i>0$ (66b)
$L\leq L_{\rm max},\quad$ (66c)
$\overline{\mathsf{SINR}}_{s}\geq\gamma_{\rm s},\quad$ (66d)
$P_{k}\leq P_{\rm tx},\quad\quad\textrm{for}\quad k=1,\cdots,N_{\rm tx}.$ (66e)
The total power consumption can be rewritten as
$P_{\rm total}=P_{\rm FIXED}+\Delta^{\rm tr}\|\boldsymbol{\rho}\|^{2}+\frac{1}{L}f_{1}+\frac{L_{d}}{L}f_{2}$ (67)
where
$P_{\rm FIXED}=\sum_{k=1}^{N_{\rm tx}}P_{\textrm{AP},0}^{\rm tx}+\sum_{r=1}^{N_{\rm rx}}P_{\textrm{AP},0}^{\rm rx}+P_{\rm fixed}+\frac{P_{\rm cloud,0}^{\rm proc}}{\sigma_{\rm cool}}$ (68)
$f_{1}=\frac{\Delta_{\rm cloud}^{\rm proc}}{\sigma_{\rm cool}C_{\rm max}}\frac{B}{10^{9}}\Bigg{(}C_{\textrm{ch-est}}+C_{\textrm{prec-comp}}+C_{\textrm{prec-comp}}^{\rm s}$
$\quad+C_{\rm se-detector}\Bigg{)}$ (69)
$f_{2}=\frac{\Delta_{\rm cloud}^{\rm proc}}{\sigma_{\rm cool}C_{\rm max}}\frac{B}{10^{9}}\left(20MN_{\rm ue}N_{\rm tx}+12MN_{\rm tx}+C_{{\rm se-prep},m}\right).$ (70)
Similar to the approach of problem P1, we define a new variable $\bar{L}$ where $L-L_{p}\leq\frac{1}{\bar{L}}$. Then, the objective function (66a) is equivalent to
$\mathsf{F_{2}}\triangleq\frac{1}{B}\left(L\left(P_{\rm FIXED}\right)+\left(L-L_{p}\right)f_{2}+\Delta^{\rm tr}\frac{\|\boldsymbol{\rho}\|^{2}}{\bar{L}}\right).$ (71)
The non-convex optimization problem in (VI-B) can be solved using the same approach in P1 as follows:
$\underset{\boldsymbol{\rho},\boldsymbol{\chi},\textbf{r}\geq\textbf{0},L,\bar{L}>L_{p},\chi_{0}\geq 0}{\textrm{minimize}}\quad\mathsf{F_{2}}+\lambda\chi_{0},$ (72a)
[…]
VII Numerical Results
In this section, we present numerical results to evaluate the performance of the proposed joint blocklength and power allocation algorithms. We consider a total area of 500 $\rm m\times$ 500 $\rm m$. The sensing location is in the center of the area, i.e., $(250,250)$. The number of antenna elements per AP is set to $M=4$. The $N_{\rm tx}=16$ ISAC transmit APs are uniformly distributed in the area. We consider the number of sensing receive APs equal to $N_{\rm rx}=1$ or $N_{\rm rx}=2$ where the first AP is located at $(200,250)$ and the second one is located at $(300,250)$. The total number of URLLC UEs in the network is $N_{\rm ue}=8$, unless otherwise stated. We set $P_{\rm tx}=100$ mW and the uplink pilot transmission power for each UE is $50$ mW.
The large-scale fading coefficients, shadowing parameters, probability of LOS, and the Rician factors are simulated based on the 3GPP Urban Microcell model, defined in [35, Table B.1.2.1-1, Table B.1.2.1-2, Table B.1.2.2.1-4]. The path losses for the Rayleigh fading target-free channels are also modeled by the 3GPP Urban Microcell model with the difference that the channel gains are multiplied by an additional scaling parameter equal to $0.3$ to suppress the known parts of the target-free channels due to LOS and permanent obstacles [25]. The sensing channel gains are computed by the bi-static radar range equation [26]. The carrier frequency, the bandwidth, and the noise variance are set to $1.9$ GHz, $200$ KHz, and $-114$ dBm, respectively. The number of pilot symbols is $L_{p}=10$.
TABLE II: Simulation Parameters
$\Delta^{\rm tr},\sigma_{\rm cool}$ 4, 0.9 $P_{\rm fixed}$ 120 W
[…]
$\Delta_{\rm cloud}^{\rm proc}$ 740 W $C_{\rm max}$ 1800 GOPS
The spatial correlation matrices for the communication channels are generated by using the local scattering model in [27, Sec. 2.5.3]. The RCS of the target is modeled by the Swerling-I model with $\sigma_{\rm rcs}=0\,\rm dBsm$. For all the UEs, we set the same packet size, the maximum transmission delay, and the DEP threshold to $b_{i}=256$ bits, $D_{i}^{\rm(th)}=1$ ms, and $\epsilon_{\rm th}=10^{-5}$, respectively unless otherwise stated. In the proposed algorithms, $\epsilon=10^{-3}$, $\epsilon_{\chi}=10^{-6}$, $\lambda=10$, and $c_{\rm max}=30$. The sensing SINR threshold is set to $\gamma_{s}=0$ dB, unless otherwise stated. The remaining parameters are detailed in Table II, where the values are consistent with those in [14], except for a scaling of the total processing capacity and the corresponding power of a DU by a factor of ten to accommodate all the processing in a single DU.
We compare the performance of two proposed algorithms: i) SeURLLC+, which aims to minimize the transmission energy consumption for as ISAC cell-free massive MIMO system with URLLC UEs, and ii) E2E SeURLLC+, which aims to minimize the E2E energy consumption. We also consider two baseline algorithms for only URLLC systems without integrating sensing: i) URLLC-only, and ii) E2E URLLC-only, which represent minimizing transmission energy and E2E energy, respectively.
Figure 2: E2E energy consumption over one transmission with different algorithms and with $N_{\rm rx}=1,2$.
In Fig. 2, we illustrate E2E energy consumption with varying algorithms and two values for the number of sensing receive APs. The chart breaks down energy consumption by system components and different operations, including transmission, ISAC APs, sensing receive APs, communication and sensing processing, and “Others” representing load-independent and idle mode power consumption at the cloud. It is shown that in URLLC-only scenarios, the majority of the energy consumption is related to the ISAC APs and then the load-independent and idle mode power consumption at the cloud. Remarkably, E2E energy minimization demonstrates a substantial reduction in energy consumption by approximately $50\%$ in URLLC-only scenarios. This underscores the importance of E2E energy-aware optimization. Conversely, in ISAC scenarios, the impact of E2E energy minimization is less pronounced, resulting in only a marginal decrease. However, we observe that the energy consumption due to sensing processing becomes predominant when $N_{\rm rx}=2$. This observation suggests a potential optimization avenue, as turning off one sensing receive AP can halve the energy consumption, contributing significantly to overall energy efficiency.
(a) E2E URLLC-only
(b) E2E SeURLLC+ , $N_{\rm rx}=1$
(c) E2E SeURLLC+ , $N_{\rm rx}=2$
Figure 3: Energy consumption in one transmission with a) E2E URLLC-only, b) E2E SeURLLC+ with $N_{\rm rx}=1$, and E2E SeURLLC+ with $N_{\rm rx}=2$.
To complete the findings in Fig. 2, we provide a comparative analysis of the energy consumption distribution across various components within the system in Fig. 3. E2E energy minimization algorithms are considered for a) URLLC-only scenario, b) ISAC scenario with 1 sensing receive AP, and c) ISAC scenario with 2 sensing receive APs. As shown in Fig. 2(a) and Fig. 2(b), the ISAC APs are the most energy consuming components. Notably, the increment in the number of sensing receive APs is accompanied by a significant rise in energy consumption corresponding to the processing of the received signals for target detection. This is evident that with the addition of another receive AP, the share of sensing processing increases up to $52\%$, approximately four times larger than the case with $N_{\rm rx}=1$.
Figure 4: Blocklength vs. decoding error probability threshold.
Figure 5: E2E energy consumption vs. decoding error probability threshold.
VII-A Impact of Reliability and Delay Requirements
In this subsection, we investigate the impact of the reliability and delay requirements on the performance of the above-mentioned algorithms. Fig. 4 shows the blocklength as a function of decoding error probability (DEP) threshold. Generally, the blocklength decreases by relaxing the reliability threshold in all the algorithms except URLLC-only. It is shown that the URLLC-only algorithm always ends up with the maximum available blocklength, i.e., $L_{\rm max}$, since increasing the blocklength allows lower transmit power to satisfy the reliability constraint. We also observe that minimizing only the transmission energy consumption will lead to higher blocklength values compared to the E2E energy minimization algorithms, i.e., E2E URLLC-only and E2E SeURLLC+. Compared to URLLC-only, the E2E URLLC-only algorithm results in a significant decrease in the blocklength to minimize the E2E energy consumption over one transmission. For ISAC scenarios, we see the same behaviours as in E2E URLLC-only. However, the blocklength is slightly higher due to the interplay between sensing power consumption and the blocklength as transmit power consumption will decrease with larger blocklength.
Fig. 5 illustrates the E2E energy consumption in terms of the DEP threshold. The results show that relaxing the reliability constraint would generally reduce the E2E energy consumption. However, ISAC scenarios demand higher energy consumption as it is expected. Moreover, we observe that algorithms that minimize E2E energy consumption outperform the corresponding algorithms which only minimize the transmission energy consumption. Particularly, when we do not integrate sensing in the system, there is a noticeable gap between performance of URLLC-only and E2E URLLC-only algorithms.
Figure 6: E2E energy consumption vs. maximum delay threshold.
The results for E2E energy consumption versus maximum delay threshold, $D_{i}^{(\rm tx)}$, are depicted in Fig. 6. As it is shown, E2E energy consumption does not change significantly with different maximum delay threshold, except for URLLC-only algorithm. The reason is that the URLLC-only algorithm would always result in the highest possible blocklength to minimize the transmission energy consumption while the optimized value for the blocklength with other algorithms would be much lower than the one with URLLC-only to minimize the energy consumption. That is why the SeURLLC+ has lower E2E energy consumption than URLLC-only. However, the results of E2E URLLC-only algorithm show that minimum E2E energy consumption could be archived by optimizing the blocklength with respect to E2E energy consumption. This behavior highlights the importance of the E2E energy minimization.
VII-B Impact of Sensing Requirement
Figure 7: Blocklength vs. sensing SINR threshold.
Figure 8: E2E energy consumption vs. sensing SINR threshold.
The impact of the sensing SINR threshold, $\gamma_{s}$, on the blocklength and E2E energy consumption is investigated in Fig. 7 and Fig. 8, respectively. E2E energy minimization leads to lower energy consumption as it is shown that E2E SeURLLC+ outperforms SeURLLC+. The reason can be found in Fig. 7. It is shown that E2E SeURLLC+ leads to lower blocklength which means lower E2E energy consumption. Moreover, we observe that the gap between the performance of these two algorithms gradually decreases by demanding higher sensing SINR threshold.
VIII Conclusion
In this paper, we proposed a joint blocklength and power control algorithm for a downlink cell-free massive MIMO system with multi-static sensing and URLLC users. We formulated two non-convex optimization problems, focusing on transmission energy minimization and E2E energy minimization, respectively. Numerical results demonstrated the superiority of the E2E energy minimization algorithm, particularly in scenarios with no sensing capability. Furthermore, our results revealed a growing significance of energy consumption related to sensing processing tasks with the increase in the number of sensing receive access points. This observation highlights a potential optimization avenue, as deactivating one sensing receive AP can effectively mitigate this dominance and contribute to overall energy efficiency. Furthermore, our findings indicated that a higher sensing SINR threshold leads to increased E2E energy consumption. This ultimately narrows the performance gap between the two proposed algorithms: the transmission-only approach and the E2E energy minimization approach.
Appendix A Proof of Theorem 1
Let us define a new optimization variable, denoted by $\overline{L}$, where $L-L_{p}\leq 1/\overline{L}$. Then, minimizing the objective function is equivalent to minimizing the following convex function (quadratic-over-linear function)
$\frac{\|\boldsymbol{\rho}\|^{2}}{\overline{L}}\,.$ (73)
Since the above function is minimized, at the optimal solution, it holds that $\overline{L}=1/L_{d}=1/(L-L_{p})$. According to (26), the reliability constraints in (63b) can be written in the form of
$\ln\left(1+\frac{\rho_{i}\mathsf{b}_{i}^{2}}{\sum_{j=0}^{{\rm N_{ue}}}\rho_{j}\mathsf{a}_{i,j}^{2}+\sigma_{n}^{2}}\right)\!\geq\frac{Q^{-1}\left(\epsilon_{i}^{(\rm th)}\right)}{\sqrt{L-L_{p}}}+\frac{b_{i}\ln{2}}{L-L_{p}}$ (74)
where $\overline{\mathsf{SINR}}^{\rm(dl)}_{i}$ is substituted by (23). To handle the non-convexity of the left-hand side in (74), we define a new variable $\chi_{i}$ and use fractional programming [36] to write the left-hand side as
$\ln(1+\chi_{i})-\chi_{i}+(1+\chi_{i})\frac{\rho_{i}\mathsf{b}_{i}^{2}}{\sum_{j=0}^{{\rm N_{ue}}}\rho_{j}\mathsf{a}_{i,j}^{2}+\rho_{i}\mathsf{b}_{i}^{2}+\sigma_{n}^{2}}.$ (75)
Moreover, to represent the upper bound to $\left(\sum_{j=0}^{{\rm N_{ue}}}\rho_{j}\mathsf{a}_{i,j}^{2}+\rho_{i}\mathsf{b}_{i}^{2}+\sigma_{n}^{2}\right)/(1+\chi_{i})$, we introduce the optimization variable $r_{i}$, similarly as in [14], for $i=1,\ldots,N_{\rm ue}$, where
$\frac{\sum_{j=0}^{{\rm N_{ue}}}\rho_{j}\mathsf{a}_{i,j}^{2}+\rho_{i}\mathsf{b}_{i}^{2}+\sigma_{n}^{2}}{1+\chi_{i}}\leq r_{i},$ (76)
which can be written as a second-order cone (SOC) constraint in (64k). We then re-cast the constraint in (75) as
$\ln\left(1+\chi_{i}\right)-\chi_{i}+\frac{\rho_{i}\mathsf{b}_{i}^{2}}{r_{i}}\geq\frac{Q^{-1}\left(\epsilon_{i}^{\rm(th)}\right)}{\sqrt{L-L_{p}}}+\frac{b_{i}\ln{2}}{L-L_{p}}$ (77)
which will not destroy optimality since we want to minimize $r_{i}$ to increase the left-hand side of the SINR constraint.
From (37), the sensing constraint in (63d) are expressed as
$\boldsymbol{\rho}^{T}\left(\gamma_{\rm s}\textbf{B}_{D}-M\,\textbf{A}_{D}\,\right)\boldsymbol{\rho}\!\leq\!-\gamma_{\rm s}MN_{\rm rx}\sigma_{n}^{2}.$ (78)
Finally, we define $\textbf{F}_{k}=\textrm{diag}\left(\sqrt{\mathbb{E}\left\{\|\textbf{w}_{0,k}\|^{2}\right\}},\ldots,\sqrt{\mathbb{E}\left\{\|\textbf{w}_{N_{\rm ue},k}\|^{2}\right\}}\right)$ and rewrite the per-AP power constraints in (63e) in SOC form in terms of $\boldsymbol{\rho}$ as $\left\|\textbf{F}_{k}\boldsymbol{\rho}\right\|\leq\sqrt{P_{\rm tx}},$ , for $k=1,\ldots,N_{\rm tx}$.

### v2

@@
 In this paper, we explore the concept of integrated sensing and communication (ISAC) within a downlink cell-free massive MIMO (multiple-input multiple-output) system featuring multi-static sensing and users requiring ultra-reliable low-latency communications (URLLC).
-Our focus involves the formulation of two non-convex algorithms that jointly solve power and blocklength allocation for end-to-end (E2E) minimization.
-The objectives are to jointly minimize sensing/communication processing and transmission energy consumption, while simultaneously meeting the requirements for sensing and URLLC.
-To address the inherent non-convexity of these optimization problems, we utilize techniques such as the Feasible Point Pursuit - Successive Convex Approximation (FPP-SCA), Concave-Convex Programming (CCP), and fractional programming.
-We conduct a comparative analysis of the performance of these algorithms in ISAC scenarios and against a URLLC-only scenario where sensing is not integrated.
-Our numerical results highlight the superior performance of the E2E energy minimization algorithm, especially in scenarios without sensing capability.
-Additionally, our study underscores the increasing prominence of energy consumption associated with sensing processing tasks as the number of sensing receive access points rises.
-Furthermore, the results emphasize that a higher sensing signal-to-interference-plus-noise ratio threshold is associated with an escalation in E2E energy consumption, thereby narrowing the performance gap between the two proposed algorithms.
+Our work focuses on devising two algorithms that concurrently optimize power and blocklength allocation.
+The objectives of these algorithms are to minimize the energy consumption solely for transmission and the overall end-to-end (E2E) energy consumption, which encompasses sensing/communication processing and transmission.
+This is achieved while meeting the necessary requirements for sensing and URLLC.
+To address the inherent non-convexity of these optimization problems, we apply feasible point pursuit - successive convex approximation (FPP-SCA), concave-convex programming (CCP), and fractional programming.
+Our numerical analysis compares the performance of these algorithms in ISAC scenarios and against a baseline URLLC-only scenario without sensing integration.
+Results show the E2E energy minimization algorithm’s effectiveness, particularly in scenarios without sensing.
+Additionally, our study underscores the increasing prominence of energy consumption associated with sensing processing tasks as the number of sensing receive access points rises, indicating potential for energy savings by selectively deactivating APs.
+Furthermore, the results emphasize that although increasing the number of transmit APs can improve URLLC satisfaction and sensing update rate by reducing blocklength, careful management is essential to avoid excessive energy consumption and reduced detection probability.
 Index Terms:
@@
 I Introduction
-Integrated sensing and communication (ISAC) aims to enhance spectral and hardware efficiency, offering potential applications in 6G communication networks, including autonomous vehicles, smart factories, and location/environment-aware scenarios [1, 2].
-Particularly, in ultra-reliable low-latency communication (URLLC) use cases like industrial Internet-of-Things (IIoT) and vehicle-to-everything (V2X) communications, ISAC can play a crucial role.
-For instance, in V2X networks, ISAC systems can detect targets such as pedestrians and vehicles, delivering sensing information reliably and with minimal delay for URLLC users like autonomous vehicles [3].
-Thus, a comprehensive study of these technologies together is imperative to enable mission-critical applications in future wireless networks.
-For many URLLC applications, such as V2X and IIoT, short codewords are needed to satisfy latency constraints where codes with short blocklengths, e.g., 100 symbols are employed.
-Therefore, the finite blocklength regime needs to be considered to model decoding error probability (DEP) [4, 5, 6].
-In such applications, short data packets should be delivered with a minimum reliability of 99.999% and an end-to-end (E2E) delay of less than 10-150 ms [7, 8].
-Therefore, optimizing the blocklength can play an important role to meet the reliability and delay requirements.
-Conversely, while ISAC introduces perceptive mobile networks through network-level sensing [9, 10, 11, 12], it is anticipated that ISAC networks may consume more energy compared to communication networks due to sensing tasks.
-This motivates a comprehensive study of the end-to-end (E2E) energy consumption of ISAC networks taking into account not only energy consumption for transmission but also energy consumption for processing tasks.
-Centralized radio access network (C-RAN) architecture has been already introduced as an energy-efficient solution for the mobile networks.
-There is also an inherent connection between cell-free massive MIMO (multiple-input multiple-output) and the C-RAN architecture due to the joint transmission and processing the UEs’ signals[13].
-However, E2E energy-aware resource orchestration is a critical aspect that must be considered when optimizing cell-free massive MIMO systems on C-RAN architectures to fully benefit from energy savings potential [14].
-I-A Related Work
-To the best of our knowledge, there are few works that jointly consider URLLC and ISAC.
+6G mobile networks are expected to offer various sensing-based applications such as autonomous vehicles, smart homes/cities/factories, remote healthcare, and location/environment-aware scenarios [1, 2].
+Many of these applications, such as autonomous vehicles and industrial Internet-of-things (IIoT), demand ultra-reliable delivering of the sensing information to the user equipments (UEs) with a minimum reliability of 99.999% and within a low end-to-end (E2E) delay of less than 10-150 ms [3, 4, 5].
+This demand necessitates two key aspects defined in 6G networks: integrated sensing and communication (ISAC) together with ultra-reliable low-latency communication (URLLC).
+For most URLLC applications, short codewords are usually needed to satisfy latency constraints where codes with short blocklengths, e.g., 50-400 symbols are employed.
+Short blocklengths result in decoding errors; therefore, to ensure reliability, it is essential to consider the finite blocklength regime for modeling the decoding error probability (DEP) [6, 7, 8].
+In mission-critical applications such as traffic control and autonomous vehicles, real-time transmission of sensing information to URLLC UEs is essential, as the speed of information updates directly impacts the effectiveness of these applications.
+In this paper, we introduce the concept of the “sensing update rate (SUR),” which quantifies the number of sensing measurements and updates per time unit [9].
+The SUR is influenced by the blocklength, with longer blocklengths resulting in lower update rates due to the extended duration of each sensing operation.
+This dependency creates an opportunity to optimize blocklength to balance reliable communication with high SUR.
+Although URLLC and ISAC have typically been studied as separate domains, a joint design and optimization approach is required to meet the distinct requirements of both and maximize integration gains.
+Cell-free massive multiple-input multiple-output (CF-mMIMO) has been shown to be a good candidate to facilitate stringent requirements of URLLC applications [8, 10, 11] as it provides high reliability by handling large path-loss variations and improves cell-edge user performance.
+CF-mMIMO can also facilitate ISAC implementations leveraging distributed access points (APs) to deploy bi-static/ multi-static sensing eliminating the need for full-duplex capability at the APs [12, 13, 14, 15].
+The central processing of network functions within a centralized radio access network (C-RAN) architecture facilitates coordination among distributed APs, enabling joint transmission and processing in CF-mMIMO[16].
+C-RAN has evolved into virtualized C-RAN and, with the open interfaces, further into Open RAN (O-RAN), where virtualization enables the use of general-purpose servers for jointly processing of baseband and sensing signals.
+However, deploying these technologies in 6G is anticipated to result in substantial energy consumption.
+In CF-mMIMO networks, baseband processing combined with sensing tasks is expected to lead to higher energy usage in ISAC networks compared to conventional communication systems.
+The impact of sensing on energy consumption in such networks has not been well-studied.
+Existing studies on green CF-mMIMO primarily focus on optimizing transmit power consumption in radio access, with limited work addressing processing energy, aside from [17].
+Virtualized C-RAN architecture facilitates centralized processing and enables the sharing of processing resources between APs and tasks, paving the way for end-to-end (E2E) optimization of both processing and radio resources in ISAC networks.
+This motivates a comprehensive study of the E2E energy consumption in ISAC networks, considering not only energy consumption for transmission but also for processing tasks.
+The primary research question addressed in this paper is how to incorporate sensing in CF-mMIMO systems with URLLC while minimizing E2E energy consumption.
+I-A Literature Review
+The finite blocklength regime and URLLC have primarily been studied in the context of cellular networks (see [18, 19, 20] and references therein), with limited research addressing URLLC in CF-mMIMO networks.
+The potentials of massive MIMO in meeting URLLC requirements for industrial automation are investigated in [21, 22, 23].
+Network availability is investigated in [21] and [22], and [23] proposes a power allocation algorithm to maximize the minimum signal-to-interference-plus-noise ratio (SINR).
+The authors in [24] propose two power optimization algorithms to provide URLLC for traditional ground UEs and unmanned aerial vehicles (UAVs).
+The objectives in [24] are maximizing either sum or minimum URLLC rate and they study neither energy minimization nor blocklength optimization.
+Max-min rate optimization is also studied in [11], where the authors adapt a special class of the conjugate beamforming for a CF-mMIMO with single-antenna APs.
+Moreover, the authors aim to address an energy efficiency optimization problem.
+Joint power and blocklength optimization is studied in [25] and [26] to optimize the reliability in a factory automation scenario and heterogeneous cellular networks, respectively.
+Cell-free networks as well as energy minimization are not studied in these works.
+Moreover, sensing requirements are not considered in [23, 21, 24, 11, 25, 26].
+There are few works that jointly consider URLLC and ISAC.
 In [5], a joint precoding scheme is proposed to minimize transmit power, satisfying sensing and delay requirements.
 Moreover, joint ISAC beamforming and scheduling design is addressed in [27] for periodic and aperiodic traffic.
-The existing works consider a single ISAC base station, while cell-free massive MIMO with multiple distributed access points (APs) has been shown to facilitate URLLC [6, 16, 17] and ISAC[18, 19] implementations.
-In the conference version [20] of this work, we studied for the first time ISAC-enabled cell-free massive MIMO with URLLC users.
-In this paper, we extend [20] from an E2E energy minimization perspective.
 The consideration of E2E energy-awareness has been explored in various contexts, as reflected in prior works such as [28, 29, 17, 30, 31].
-In particular, [14] studied fully virtualized E2E power minimization problem for cell-free massive MIMO on open radio access network (O-RAN) architecture by taking the radio, fronthaul, and processing resources into account.
-However, this work neither considered ISAC nor URLLC communications.
+In particular, [30] studied fully virtualized E2E power minimization problem for CF-mMIMO on O-RAN architecture by taking the radio, fronthaul, and processing resources into account.
+Joint user scheduling and power allocation scheme for cell-free ISAC systems is studied in [32] that aims to maximize the sum-rate of the communication UEs and the sensing targets.
+However, the authors do not consider either URLLC requirements or energy minimization.
+To the best of our knowledge, for the first time, ISAC in CF-mMIMO with URLLC UEs and with the focus on energy efficiency was studied in our previous work[33].
+In [33], we propose a successive convex approximation-based power allocation algorithm that maximizes energy efficiency while satisfying the sensing and URLLC requirements for a given blocklength.
+However, the impact of blocklength optimization is not considered in [33].
+In this paper, we extend [33] from an E2E energy minimization perspective and aim to address a joint blocklength and power optimization problem.
+Here, E2E refers to considering energy consumption at both radio site and the central cloud.
 I-B Contributions
-In this paper, we study E2E energy consumption in a cell-free massive MIMO system with URLLC user equipments (UEs) and multi-static sensing in a cluttered environment.
-Both sensing and communication signals contribute to the sensing task.
-However, the sensing signals can cause interference for the UEs.
-Thus, we design the precoding vectors to null the interference only for the UEs.
-We consider a maximum DEP threshold, representing the reliability requirement, together with a maximum transmission delay threshold as the URLLC requirements and a minimum signal-to-interference-plus-noise ratio (SINR) as the sensing requirement.
-The latter is motivated by the fact that given a fixed false alarm probability, detection probability increases with a higher sensing SINR [25], [26, Chap.
+In this paper, we study E2E energy consumption in a CF-mMIMO system with URLLC UEs and multi-static sensing in a cluttered environment.
+We sense one candidate location of the target utilizing communication signals and an additional sensing signal.
+The sensing signals can cause interference for the UEs.
+Thus, we design the sensing precoding vectors to null the interference for the UEs.
+We consider a maximum DEP threshold, representing the reliability requirement, together with a maximum transmission delay threshold as the URLLC requirements and a minimum SINR as the sensing requirement.
+The latter is motivated by the fact that given a fixed false alarm probability, detection probability increases with a higher sensing SINR [34], [9, Chap.
 3 and 15].
-In the conference version of this work [20], we propose a successive convex approximation-based power allocation algorithm that maximizes energy efficiency while satisfying the sensing and URLLC requirements for a given blocklength.
-However, the impact of blocklength optimization is not considered in [20].
-Different from the conference paper, we conduct a comprehensive analysis of the processing requirements, specifically in terms of giga operations per second (GOPS), and examine its impact on processing energy consumption.
-The interplay between processing requirements and blocklength yields intriguing results, as demonstrated by the simulation findings presented in this paper.
+Different from [33], we conduct a comprehensive analysis of the processing requirements, specifically in terms of giga-operations per second (GOPS), and examine its impact on processing energy consumption.
 The main contributions of this paper are outlined as follows:
-  • We derive an upper bound for the DEP and transmission delay in finite blocklength cell-free massive MIMO communications.
-  • We propose two joint blocklength and power allocation algorithms, which minimize the transmission energy consumption and E2E energy consumption, respectively while meeting URLLC and sensing requirements.
-The former algorithm is represented by SeURLLC+ and the latter is represented by E2E URLLC+.
-To solve the non-convex optimization problems, we benefit from the Feasible Point Pursuit - Successive Convex Approximation (FPP-SCA), Concave-Convex Programming (CCP), and fractional programming.
-  • We investigate the impact of E2E energy minimization, encompassing both radio and processing energy consumption, in the above-mentioned ISAC scenario, as well as a scenario where only URLLC requirements are taken into account.
-The algorithms for URLLC-only scenarios are represented by URLLC-only and E2E URLLC-only.
-The former corresponds to transmission energy minimization and the latter corresponds to E2E energy minimization.
-  • We conduct a sensitivity analysis by investigating the impact of DEP threhsold, maximum delay threshold, and sensing SINR threshold on E2E energy consumption.
-The rest of the paper is organized as follows.
-In Section II, we present the system model.
-In Section III, we provide URLLC analysis accounting for decoding error probability and delay in finite blocklength regime.
-Section IV is devoted to the sensing analysis.
-In Section V, we describe the E2E power model and derive the GOPS analysis for both communication and sensing.
-Next, we present our optimization problems in Section VI.
-Finally, the numerical results and conclusion are provided in Section VII and Section VIII, respectively.
+  • We propose two joint blocklength and power allocation algorithms aimed at minimizing transmission energy consumption and E2E energy consumption, respectively, while satisfying URLLC and sensing requirements.
+To address the non-convex optimization challenges, we employ feasible point pursuit - successive convex approximation (FPP-SCA), concave-convex programming (CCP), and fractional programming.
+  • We analyze the impact of E2E energy minimization, covering both radio and processing energy consumption, within the proposed ISAC scenario and a comparative scenario that considers only URLLC requirements.
+  • We examine two distinct target detection schemes, comparing their performance in terms of E2E energy consumption, detection probability, and sensing update rate.
+  • We conduct sensitivity analyses to evaluate the effects of DEP threshold, sensing SINR threshold, and the number of ISAC transmit APs and sensing receive APs on E2E energy consumption, detection probability, and sensing update rate.
+The rest of the paper is organized as follows: Section II introduces the system model.
+Section III provides URLLC analysis, considering DEP and delay in finite blocklength regime.
+Section IV covers the sensing analysis.
+Section V describes the E2E power model and derives GOPS analysis for both communication and sensing.
+Optimization problems are presented in Section VI, followed by numerical results and conclusions in Sections VII and VIII, respectively.
 II System Model
-We consider an ISAC-enabled cell-free massive MIMO system in URLLC scenarios.
-The system adopts a centralized radio access network (C-RAN) architecture [27] for uplink channel estimation and downlink communication, as well as multi-static sensing as shown in Fig.
+We study ISAC in a CF-mMIMO system in URLLC scenarios.
+The system adopts a virtualized C-RAN architecture [35] for uplink channel estimation and downlink communication, as well as multi-static sensing as shown in Fig.
 1.
 All the APs are interconnected via fronthaul links to the edge cloud and operate in full synchronization.
-We consider the original form of cell-free massive MIMO [6], wherein all the $N_{\rm tx}$ ISAC APs jointly serve the $N_{\rm ue}$ URLLC UEs by transmitting precoded signals containing both communication and sensing symbols.
-Figure 1: Illustration of the ISAC system in cell-free massive MIMO with URLLC UEs.
+We consider the original form of CF-mMIMO [8], wherein all the $N_{\rm tx}$ ISAC APs jointly serve the $N_{\rm ue}$ URLLC UEs by transmitting precoded signals containing both communication and sensing symbols.
+Figure 1: Illustration of an ISAC system in CF-mMIMO with URLLC UEs.
 Concurrently, the $N_{\rm rx}$ sensing receive APs engage in simultaneous sensing of the candidate location to detect the target.
@@
 It is expected that duration of each URLLC transmission, denoted by $T$, is shorter than one coherence time $T_{c}$, i.e., $T<T_{c}$ [11].
-Without loss of generality, we assume the coherence time is equal to URLLC transmission time so that in for each transmission we first estimate the channel.
-II-A Channel Modeling and Downlink Communications
+Without loss of generality, we assume that we have a channel estimation phase in each transmission.
+II-A Downlink Communication
+Let $s_{i}[m]$ and $s_{0}[m]$ represent the downlink communication symbol for UE $i$ and sensing symbol, respectively at time instance $m$.
+The symbols are independent and have zero mean and unit power.
+Moreover, let $\rho_{i}\geq 0$ and $\rho_{0}\geq 0$ be, respectively, the power control coefficients for UE $i$ and the target, which are fixed throughout the transmission.
+Then, the transmitted signal from transmit AP $k\in\{1,\cdots,N_{\rm tx}\}$ at time instance $m$ is
+$\vspace{-1mm}\textbf{x}_{k}[m]=\sum_{i=0}^{N_{\rm ue}}\sqrt{\rho_{i}}\textbf{w}_{i,k}s_{i}[m]=\textbf{W}_{k}\textbf{D}_{\rm s}[m]\bm{\rho},\vspace{-1mm}$ (1)
+where the vectors $\textbf{w}_{i,k}\in\mathbb{C}^{M}$ and $\textbf{w}_{0,k}\in\mathbb{C}^{M}$ are the transmit precoding vectors for transmit AP $k$ corresponding to UE $i$ and the sensing signal, respectively.
+In (1), $\textbf{W}_{k}=\begin{bmatrix}\textbf{w}_{0,k}&\textbf{w}_{1,k}&\cdots&\textbf{w}_{N_{\rm ue},k}\end{bmatrix}$, ${\textbf{D}}_{\rm s}[m]=\textrm{diag}\left(s_{0}[m],s_{1}[m],\ldots,s_{N_{\rm ue}}[m]\right)$ is the diagonal matrix containing the sensing and communication symbols, and $\bm{\rho}=[\sqrt{\rho_{0}}\ \ldots\ \sqrt{\rho_{N_{\rm ue}}}]^{T}$.
 The communication channels are modeled as spatially correlated Rician fading which are assumed to remain constant during each coherence block, and the channel realizations are independent of each other.
@@
 We concatenate the channel vector $\textbf{h}_{i,k}$ in the collective channel vector
-$\textbf{h}_{i}=\begin{bmatrix}\textbf{h}_{i,1}^{T}&\ldots&\textbf{h}_{i,N_{\rm tx}}^{T}\end{bmatrix}^{T}\in\mathbb{C}^{N_{\rm tx}M}$ (3)
-for UE $i$.
-Let $s_{i}[m]$ and $s_{0}[m]$ represent the downlink communication symbol for UE $i$ and sensing symbol, respectively at time instance $m$.
-The symbols are independent and have zero mean and unit power.
-Moreover, let $\rho_{i}\geq 0$ and $\rho_{0}\geq 0$ be, respectively, the power control coefficients for UE $i$ and the target, which are fixed throughout the transmission.
-Then, the transmitted signal from transmit AP $k\in\{1,\cdots,N_{\rm tx}\}$ at time instance $m$ is
-$\vspace{-1mm}\textbf{x}_{k}[m]=\sum_{i=0}^{N_{\rm ue}}\sqrt{\rho_{i}}\textbf{w}_{i,k}s_{i}[m]=\textbf{W}_{k}\textbf{D}_{\rm s}[m]\boldsymbol{\rho},\vspace{-1mm}$ (4)
-where the vectors $\textbf{w}_{i,k}\in\mathbb{C}^{M}$ and $\textbf{w}_{0,k}\in\mathbb{C}^{M}$ are the transmit precoding vectors for transmit AP $k$ corresponding to UE $i$ and the sensing signal, respectively.
-In (4), $\textbf{W}_{k}=\begin{bmatrix}\textbf{w}_{0,k}&\textbf{w}_{1,k}&\cdots&\textbf{w}_{N_{\rm ue},k}\end{bmatrix}$, ${\textbf{D}}_{\rm s}[m]=\textrm{diag}\left(s_{0}[m],s_{1}[m],\ldots,s_{N_{\rm ue}}[m]\right)$ is the diagonal matrix containing the sensing and communication symbols, and $\boldsymbol{\rho}=[\sqrt{\rho_{0}}\ \ldots\ \sqrt{\rho_{N_{\rm ue}}}]^{T}$.
-Moreover, let $n_{i}[m]\!\sim\!\mathcal{CN}(0,\sigma_{n}^{2})$ be the independent receiver noise at UE $i$ at time instance $m$. The received signal at UE $i$ is given as
+$\textbf{h}_{i}=\begin{bmatrix}\textbf{h}_{i,1}^{T}&\ldots&\textbf{h}_{i,N_{\rm tx}}^{T}\end{bmatrix}^{T}\in\mathbb{C}^{N_{\rm tx}M},$ (3)
+for UE $i$. The received signal at UE $i$ is given as
 $y_{i}[m]=$ $\underbrace{\sqrt{\rho_{i}}\textbf{h}_{i}^{H}\textbf{w}_{i}s_{i}[m]}_{\textrm{Desired signal}}+\underbrace{\sum_{j=1,j\neq i}^{N_{\rm ue}}\sqrt{\rho_{j}}\textbf{h}_{i}^{H}\textbf{w}_{j}s_{j}[m]}_{\textrm{Interference signal due to the other UEs}}$
 $+\underbrace{\sqrt{\rho_{0}}\textbf{h}_{i}^{H}\textbf{w}_{0}s_{0}[m]}_{\textrm{Interference signal due to the sensing}}+\underbrace{n_{i}[m]}_{\textrm{Noise}},$ (4)
-where the collective precoding vectors
+where $n_{i}[m]\!\sim\!\mathcal{CN}(0,\sigma_{n}^{2})$ is the independent receiver noise at UE $i$ at time instance $m$ and the collective precoding vectors
 ${\textbf{w}_{i}=\left[\textbf{w}_{i,1}^{T}\ \textbf{w}_{i,2}^{T}\ \ldots\ \textbf{w}_{i,N_{\rm tx}}^{T}\right]^{T}\in\mathbb{C}^{N_{\rm tx}M}},$ (5)
@@
 are the centralized precoding vectors.
-The vectors $\textbf{w}_{i,k}\in\mathbb{C}^{M}$ and $\textbf{w}_{0,k}\in\mathbb{C}^{M}$ are the transmit precoding vectors for ISAC AP $k$ corresponding to UE $i$ and the sensing signal, respectively.
 Using the independence of the data and sensing signals, the average transmit power for transmit AP $k\in\{1,\ldots,N_{\rm tx}\}$ is
@@
 Within the URLLC transmission time comprising $L$ symbols, $L_{p}<L$ symbols are allocated for pilot transmission.
-A collection of $L_{p}$ mutually orthogonal pilot sequences is represented by $\boldsymbol{\phi}_{1},\ldots,\boldsymbol{\phi}_{L_{p}}\in\mathbb{C}^{L_{p}}$, where $\|\boldsymbol{\phi}_{t}\|^{2}=L_{p}$ for $t=1,\ldots,L_{p}$.
+A collection of $L_{p}$ mutually orthogonal pilot sequences is represented by $\bm{\phi}_{1},\ldots,\bm{\phi}_{L_{p}}\in\mathbb{C}^{L_{p}}$, where $\|\bm{\phi}_{t}\|^{2}=L_{p}$ for $t=1,\ldots,L_{p}$.
 These pilot sequences are assigned to the $N_{\rm ue}$ UEs.
@@
 The received signal at the ISAC AP $k$ during the entire pilot transmission is denoted by $\textbf{Y}_{k}^{p}\in\mathbb{C}^{M\times L_{p}}$ where
-$\textbf{Y}_{k}^{p}=\sum_{i=1}^{N_{\rm ue}}\sqrt{\eta_{i}}\textbf{h}_{i,k}\boldsymbol{\phi}_{t_{i}}^{T}+\textbf{N}_{k},$ (9)
+$\textbf{Y}_{k}^{p}=\sum_{i=1}^{N_{\rm ue}}\sqrt{\eta_{i}}\textbf{h}_{i,k}\bm{\phi}_{t_{i}}^{T}+\textbf{N}_{k},$ (8)
 where $\eta_{i}>0$ is the transmit power by UE $i$ and $\textbf{N}_{k}\in\mathbb{C}^{M\times L_{p}}$ is the received noise at AP $k$ with independent and identically distributed (i.i.d.) $\mathcal{CN}(0,\sigma_{n}^{2})$ entries.
 We assume that $\bar{\textbf{h}}_{i,k}$ and $\textbf{R}_{i,k}$, representing the long-term statistics for the channel in (2), are known while the random phase-shift $\varphi_{i,k}$ is unknown.
-With respect to estimating $\textbf{h}_{i,k}$, let us form $\textbf{y}_{t_{i},k}^{p}=\frac{1}{\sqrt{L_{p}}}\textbf{Y}_{k}^{p}\boldsymbol{\phi}_{t_{i}}^{*}\in\mathbb{C}^{M}$, where
+With respect to estimating $\textbf{h}_{i,k}$, let us form $\textbf{y}_{t_{i},k}^{p}=\textbf{Y}_{k}^{p}\bm{\phi}_{t_{i}}^{*}\in\mathbb{C}^{M}$, where
 $\textbf{y}_{t_{i},k}^{p}=\underbrace{\sqrt{\eta_{i}}\,L_{p}\textbf{h}_{i,k}}_{\textrm{Desired part}}+\underbrace{\sum_{j\in\mathcal{P}_{i}\setminus\{i\}}\sqrt{\eta_{j}}\,L_{p}\textbf{h}_{j,k}}_{\textrm{Interference part}}+\underbrace{\textbf{n}_{t_{i},k}}_{\textrm{Noise}},$ (9)
-where $\textbf{n}_{t_{i},k}=\frac{1}{\sqrt{L_{p}}}\textbf{N}_{k}\boldsymbol{\phi}_{t_{i}}^{*}\sim\mathcal{CN}(\textbf{0},\sigma_{n}^{2}\textbf{I}_{M})$.
-If the number of UEs is greater than the number of mutually orthogonal pilot sequences, then each pilot sequence may be assigned to multiple UEs using the pilot assignment algorithm in [27, Alg.
+where $\textbf{n}_{t_{i},k}=\textbf{N}_{k}\bm{\phi}_{t_{i}}^{*}\sim\mathcal{CN}(\textbf{0},\sigma_{n}^{2}L_{p}\textbf{I}_{M})$.
+If the number of UEs is greater than the number of mutually orthogonal pilot sequences, then each pilot sequence may be assigned to multiple UEs using the pilot assignment algorithm in [35, Alg.
 4.1].
 The LMMSE estimate of $\textbf{h}_{i,k}$ is given by [37]
-$\hat{\textbf{h}}_{i,k}=\sqrt{\eta_{i}L_{p}}\textbf{R}_{i,k}^{\prime}\left(\boldsymbol{\Psi}^{\prime}_{t_{i},k}\right)^{-1}\textbf{y}_{t_{i},k}^{p},$ (11)
-where $\textbf{R}_{i,k}^{\prime}$ and $\boldsymbol{\Psi}_{t_{i},k}^{\prime}$ are the spatial correlation matrix corresponding to channel $\textbf{h}_{i,k}$ and the correlation matrix of the received signal, respectively, given by
+$\hat{\textbf{h}}_{i,k}=\sqrt{\eta_{i}}\textbf{R}_{i,k}^{\prime}\left(\bm{\Psi}^{\prime}_{t_{i},k}\right)^{-1}\textbf{y}_{t_{i},k}^{p},$ (10)
+where $\textbf{R}_{i,k}^{\prime}$ and $\bm{\Psi}_{t_{i},k}^{\prime}$ are the spatial correlation matrix corresponding to channel $\textbf{h}_{i,k}$ and the correlation matrix of the received signal, respectively, given by
 $\textbf{R}_{i,k}^{\prime}\triangleq\textbf{R}_{i,k}+\bar{\textbf{h}}_{i,k}\bar{\textbf{h}}_{i,k}^{H}$ (11)
-$\boldsymbol{\Psi}^{\prime}_{t_{i},k}\triangleq\sum_{j\in\mathcal{P}_{i}}\eta_{j}L_{p}\textbf{R}_{j,k}^{\prime}+\sigma_{n}^{2}\textbf{I}_{M}.$ (13)
+$\bm{\Psi}^{\prime}_{t_{i},k}\triangleq\sum_{j\in\mathcal{P}_{i}}\eta_{j}L_{p}\textbf{R}_{j,k}^{\prime}+\sigma_{n}^{2}\textbf{I}_{M}.$ (12)
 It is worth noting that deriving MMSE channel estimation proves challenging due to the non-Gaussian distribution of channels resulting from random phase-shifts in the LOS paths.
@@
 Each receive AP sends their respective signals $\textbf{y}_{r}[m]$, for $r=1,\ldots,N_{\rm rx}$, to the edge cloud, to form the collective received signal $\textbf{y}[m]=\begin{bmatrix}\textbf{y}_{1}^{T}[m]&\cdots&\textbf{y}_{N_{\rm rx}}^{T}[m]\end{bmatrix}^{T}$, which can be expressed as
-$\textbf{y}[m]$ $=\underbrace{\mathrm{blkdiag}\left(\textbf{G}_{1}[m],\ldots,\textbf{G}_{N_{\rm rx}}[m]\right)}_{\triangleq\textbf{G}[m]}\boldsymbol{\alpha}$
-$+\underbrace{\left(\textbf{I}_{N_{\rm rx}}\otimes\left(\textbf{x}^{T}[m]\otimes\textbf{I}_{M}\right)\right)}_{\triangleq\textbf{X}[m]}\boldsymbol{\mathfrak{h}}+\textbf{n}[m],$ (17)
-where $\textbf{G}_{r}[m]=\begin{bmatrix}\textbf{g}_{r,1}[m]&\cdots&\textbf{g}_{r,N_{\rm tx}}[m]\end{bmatrix}\in\mathbb{C}^{M\times N_{\rm tx}}$, $\textbf{x}[m]=\begin{bmatrix}\textbf{x}^{T}_{1}[m]&\cdots&\textbf{x}^{T}_{N_{\rm tx}}[m]\end{bmatrix}^{T}\in\mathbb{C}^{N_{\rm tx}M}$, and $\mathfrak{h}\sim\mathcal{CN}\left(\textbf{0},\textbf{R}\right)$ is the vectorized target-free channels[25, Sec.
+$\textbf{y}[m]$ $=\underbrace{\mathrm{blkdiag}\left(\textbf{G}_{1}[m],\ldots,\textbf{G}_{N_{\rm rx}}[m]\right)}_{\triangleq\textbf{G}[m]}\bm{\alpha}$
+$+\underbrace{\left(\textbf{I}_{N_{\rm rx}}\otimes\left(\textbf{x}^{T}[m]\otimes\textbf{I}_{M}\right)\right)}_{\triangleq\textbf{X}[m]}\bm{\mathfrak{h}}+\textbf{n}[m],$ (16)
+where $\textbf{G}_{r}[m]=\begin{bmatrix}\textbf{g}_{r,1}[m]&\cdots&\textbf{g}_{r,N_{\rm tx}}[m]\end{bmatrix}\in\mathbb{C}^{M\times N_{\rm tx}}$, $\textbf{x}[m]=\begin{bmatrix}\textbf{x}^{T}_{1}[m]&\cdots&\textbf{x}^{T}_{N_{\rm tx}}[m]\end{bmatrix}^{T}\in\mathbb{C}^{N_{\rm tx}M}$, and $\mathfrak{h}\sim\mathcal{CN}\left(\textbf{0},\textbf{R}\right)$ is the vectorized target-free channels[34, Sec.
 V].
-II-D Transmit ISAC Precoding Vectors
+II-D Target Detection
+We employ two maximum a posteriori ratio test (MAPRT) detectors with two levels of complexity: the clutter-unaware detector [13] and the clutter-aware detector [34, Lem.
+2].
+The clutter-unaware detector ignores the presence of clutter for the sake of complexity reduction, whereas the clutter-aware detector accounts for unknown clutter.
+The corresponding test statistics are given by
+$T_{\text{c-unaware}}=\textbf{a}^{H}\textbf{C}^{-1}\textbf{a},$ (17)
+$T_{\text{c-aware}}=\begin{bmatrix}\textbf{a}\\ \textbf{b}\end{bmatrix}^{H}\left(\begin{bmatrix}\textbf{C}&\textbf{E}\\ \textbf{E}^{H}&\textbf{D}\end{bmatrix}^{-1}-\begin{bmatrix}\textbf{0}&\textbf{0}\\ \textbf{0}&\textbf{D}^{-1}\end{bmatrix}\right)\begin{bmatrix}\textbf{a}\\ \textbf{b}\end{bmatrix},$ (18)
+respectively, where
+$\textbf{a}=\sum_{m=1}^{L_{d}}\textbf{G}^{H}[m]\textbf{y}[m],\quad\textbf{b}=\sum_{m=1}^{L_{d}}\textbf{X}^{H}[m]\textbf{y}[m],$ (19)
+$\textbf{C}=\sum_{m=1}^{L_{d}}\textbf{G}^{H}[m]\textbf{G}[m]+\sigma_{n}^{2}\textbf{I}_{N_{\rm rx}N_{\rm tx}},$ (20)
+$\textbf{D}=\sum_{m=1}^{L_{d}}\textbf{X}^{H}[m]\textbf{X}[m]+\sigma_{n}^{2}\textbf{R}^{-1},$ (21)
+$\textbf{E}=\sum_{m=1}^{L_{d}}\textbf{G}^{H}[m]\textbf{X}[m].$ (22)
+II-E Transmit ISAC Precoding Vectors
 The communication and sensing transmit precoding vectors are obtained based on regularized zero forcing (RZF) and zero forcing (ZF) approaches, respectively.
@@
 $R_{i}$ $\!\geq\!\left(1\!-\!\beta\right)\!\mathbb{E}\Bigg{\{}\!\!\log_{2}{\left(1+\mathsf{SINR}^{(\rm dl)}_{i}\right)}\!\!\Bigg{\}}\!-\!\frac{Q^{-1}(\epsilon_{i})}{\ln(2)}\sqrt{\frac{(1-\beta)}{L}}.$ (26)
-Moreover, given that only $\mathbb{E}\left\{\textbf{h}_{i}^{H}\textbf{w}_{i}\right\}$ is known at UE $i$, and according to [27, Thm. 6.1] and [25, Lem. 1],
+Moreover, given that only $\mathbb{E}\left\{\textbf{h}_{i}^{H}\textbf{w}_{i}\right\}$ is known at UE $i$, and according to [35, Thm. 6.1] and [34, Lem. 1],
 $\mathbb{E}\left\{\log_{2}{\left(1+\mathsf{SINR}^{\rm(dl)}_{i}\right)}\right\}\geq\log_{2}{\left(1+\overline{\mathsf{SINR}}^{(\rm dl)}_{i}\right)}$ (27)
@@
 IV Sensing Analysis
-We have numerically shown in [19] and [25], that maximizing the sensing SINR improves the target detection probability under a constant false alarm probability.
+We assess the sensing performance in terms of detection probability under a certain false alarm probability and sensing update rate in terms of the number of sensing measurement updates per time unit.
+It is numerically shown in [13] and [34], that maximizing the sensing SINR improves the target detection probability under a constant false alarm probability.
 When studying other sensing tasks, it is naturally desired to keep the sensing SINR above a required threshold denoted by $\gamma_{\rm s}$, i.e., $\mathsf{SINR}_{s}\geq\gamma_{\rm s}$.
 This fact motivates the optimization problems presented in Section VI.
-Using the previously defined vector $\boldsymbol{\rho}=[\sqrt{\rho_{0}}\ \ldots\ \sqrt{\rho_{N_{\rm ue}}}]^{T}$, from [25], the sensing SINR with independent RCS values is given by
-$\mathsf{SINR_{\rm s}}=\frac{\boldsymbol{\rho}^{T}\textbf{A}\boldsymbol{\rho}}{L_{d}MN_{\rm rx}\sigma_{n}^{2}+\boldsymbol{\rho}^{T}\textbf{B}\boldsymbol{\rho}},$ (30)
-where
+In the following subsections, we present sensing SINR and sensing update rate.
+IV-A Sensing SINR
+From [34] , the sensing SINR is given by
+$\mathsf{SINR_{\rm s}}=\frac{\bm{\rho}^{T}\textbf{A}\bm{\rho}}{L_{d}MN_{\rm rx}\sigma_{n}^{2}+\bm{\rho}^{T}\textbf{B}\bm{\rho}},$ (35)
+where $\bm{\rho}=[\sqrt{\rho_{0}}\ \ldots\ \sqrt{\rho_{N_{\rm ue}}}]^{T}$ and
 […]
@@
 ×$\hskip 8.53581pt\times\textbf{D}_{\rm s}[m].$ (37)
-As we see, the sensing SINR is a function of symbols which are random for different blocklength values.
-When the blocklength is fixed, we can assume that the symbols are known in advance.
+The sensing SINR is a function of symbols, which vary randomly for different blocklength values.
 This implies that although the symbols are known when processing the received sensing signals, they cannot be known during the resource allocation phase.
 To this end, we take the expectation with respect to the random symbols.
-Then, the average sensing SINR would be
-$\overline{\mathsf{SINR}}_{s}$ $=\frac{\boldsymbol{\rho}^{T}\mathbb{E}\{\textbf{A}\}\boldsymbol{\rho}}{L_{d}MN_{\rm rx}\sigma_{n}^{2}+\boldsymbol{\rho}^{T}\mathbb{E}\{\textbf{B}\}\boldsymbol{\rho}},$ (33)
-The symbols $s_{i}[m]$ are independent, having zero mean and unit variance.
-Therefore, $\mathbb{E}\{s_{i}^{*}[m]s_{i}[m]\}=1$ and $\mathbb{E}\{s_{i}^{*}[m]s_{j}[m]\}=0$ for $i\neq j$.
-Then, the expectations in (33) are obtained as follows
-$\mathbb{E}\{\textbf{A}\}=L_{d}\,M\textbf{A}_{D},\quad$ $\mathbb{E}\{\textbf{B}\}=L_{d}\,\textbf{B}_{D},$ (34)
-where
-$\textbf{A}_{D}=\mathrm{diag}\Bigg{(}\mathrm{diag}\Bigg{(}$ $\sum_{r=1}^{N_{\rm rx}}\sum_{k=1}^{N_{\rm tx}}\beta_{r,k}\textbf{W}_{k}^{H}\textbf{a}^{*}(\varphi_{k},\vartheta_{k})$
-×$\times\textbf{a}^{T}(\varphi_{k},\vartheta_{k})\textbf{W}_{k}\Bigg{)}\Bigg{)}$ (35)
-$\textbf{B}_{D}=\mathrm{diag}\Bigg{(}\mathrm{diag}\Bigg{(}$ $\sum_{r=1}^{N_{\rm rx}}\sum_{k=1}^{N_{\rm tx}}\mathrm{tr}\left(\textbf{R}_{\mathrm{rx},(r,k)}\right)\textbf{W}_{k}^{H}\textbf{R}^{T}_{{\rm tx},(r,k)}\textbf{W}_{k}\Bigg{)}\Bigg{)}.$ (36)
-The average sensing SINR with respect to symbols and RCS values is finally written as
-$\overline{\mathsf{SINR}}_{s}$ $=\frac{M\boldsymbol{\rho}^{T}\textbf{A}_{D}\boldsymbol{\rho}}{MN_{\rm rx}\sigma_{n}^{2}+\boldsymbol{\rho}^{T}\textbf{B}_{D}\boldsymbol{\rho}}.$ (37)
+Thus, the average sensing SINR would be
+$\overline{\mathsf{SINR}}_{s}$ $=\frac{M\bm{\rho}^{T}\textbf{A}_{D}\bm{\rho}}{MN_{\rm rx}\sigma_{n}^{2}+\bm{\rho}^{T}\textbf{B}_{D}\bm{\rho}},$ (38)
+where $\textbf{A}_{D}$ and $\textbf{B}_{D}$ are diagonal matrices with
+$\left[\textbf{A}_{D}\right]_{ii}$ $=\sum_{r=1}^{N_{\rm rx}}\sum_{k=1}^{N_{\rm tx}}\beta_{r,k}\left(\textbf{W}_{k}^{H}\textbf{a}^{*}(\varphi_{k},\vartheta_{k})\textbf{a}^{T}(\varphi_{k},\vartheta_{k})\textbf{W}_{k}\right)_{ii},$ (39)
+$\left[\textbf{B}_{D}\right]_{ii}$ $=\sum_{r=1}^{N_{\rm rx}}\sum_{k=1}^{N_{\rm tx}}\mathrm{tr}\left(\textbf{R}_{\mathrm{rx},(r,k)}\right)\left(\textbf{W}_{k}^{H}\textbf{R}^{T}_{{\rm tx},(r,k)}\textbf{W}_{k}\right)_{ii}.$ (40)
+IV-B Sensing Update Rate (SUR)
+In our system model, the sensing information is updated once throughout each transmission block. Thus, the sensing update rate, denoted as $R_{\rm s}$, is
+$R_{\rm s}=\frac{B}{L}\quad(\rm updates/second).$ (41)
 V E2E Power Consumption Modeling
 Compared to communication networks, ISAC networks are expected to consume more power due to sensing tasks.
-In general, E2E power consumption in a system with C-RAN architecture is consisting of two main components: i) the radio site power consumption, including the AP power consumption and ii) the power consumption at the edge cloud, denoted by $P_{\rm cloud}$ [23].
-In the considered C-RAN architecture, all processing is done in the cloud.
+In general, E2E power consumption in a system with virtualized C-RAN architecture is consisting of two main components: i) the radio site power consumption, including the AP power consumption and ii) the power consumption at the edge cloud, denoted by $P_{\rm cloud}$ [17].
+In the considered virtualized C-RAN architecture, all processing is done in the cloud.
 Let $P_{\textrm{AP},0}^{\rm tx}$ and $P_{\textrm{AP},0}^{\rm rx}$ be the static power consumption of the transmit ISAC AP and the receive sensing AP, respectively.
@@
 $\sigma_{\rm cool}\in(0,1]$ and $\Delta_{\rm cloud}^{\rm proc}$ denote the cooling efficiency of the cloud and the slope of the load-dependent power consumption for processing at the digital unit (DU) in the cloud, respectively.
-Moreover, $C_{\rm max}$ and $C_{\rm cloud}\in[0,C_{\rm max}]$ are the maximum processing capacity of the processing resources and the total processing resource utilization in giga-operations per second (GOPS), respectively[14, 31].
+Moreover, $C_{\rm max}$ and $C_{\rm cloud}\in[0,C_{\rm max}]$ are the maximum processing capacity of the processing resources and the total processing resource utilization in GOPS, respectively[30, 39].
 The processing resource utilization can be expressed as
@@
 In this section, we analyze the GOPS for digital signal processing corresponding to the communication tasks including the uplink channel estimation and downlink transmission.
-To compute the number of real multiplications, we mainly follow the GOPS analysis in [28, App.
+To compute the number of real multiplications, we mainly follow the GOPS analysis in [36, App.
 B], [17].
@@
 To compute the channel estimates, we first obtain the vectors $\textbf{y}_{t_{i},k}^{p}$.
-From [28, App.
+From [36, App.
 B], the multiplication of one matrix of size $M\times L_{p}$ with a vector of size $L_{p}\times 1$ results in $ML_{p}$ complex multiplications.
@@
 Otherwise, the number of real multiplications/divisions would be $8ML_{p}^{2}N_{\rm tx}$.
-Moreover, we need to compute the matrices $\textbf{R}_{i,k}^{\prime}$ and $\boldsymbol{\Psi}_{t_{i},k}^{\prime}$, given that $\bar{\textbf{h}}_{i,k}$ and $\textbf{R}_{i,k}$ are known.
+Moreover, we need to compute the matrices $\textbf{R}_{i,k}^{\prime}$ and $\bm{\Psi}_{t_{i},k}^{\prime}$, given that $\bar{\textbf{h}}_{i,k}$ and $\textbf{R}_{i,k}$ are known.
 However, this pre-computation can be neglected since the channel statistics are usually constant for a while and there is no need to compute them every coherence block.
@@
 $C_{\textrm{ch-est}}=\left\{\begin{matrix}(8ML_{p}+8M^{2})N_{\rm ue}N_{\rm tx},&L_{p}\geq N_{\rm ue}\\ 8ML_{p}^{2}N_{\rm tx}+8M^{2}N_{\rm ue}N_{\rm tx},&L_{p}<N_{\rm ue}.&\end{matrix}\right.$ (45)
-The number of real multiplications/divisions to compute centralized RZF precoding vector for all the UEs from [28, App. B] is
+The number of real multiplications/divisions to compute centralized RZF precoding vector for all the UEs from [36, App. B] is
 $C_{\textrm{prec-comp}}$ $=\underbrace{\frac{8\left((MN_{\rm tx})^{2}+MN_{\rm tx}\right)}{2}N_{\rm ue}}_{\textrm{inside the parentheses}}$
@@
 V-B GOPS Analysis of Digital Operations for Sensing at the Cloud
-In this subsection, we compute the GOPS for sensing operations.
-In downlink, first the sensing precoding vector $\textbf{w}_{0}$ should be obtained.
-Next operation is multiplying the sensing symbols by the sensing precoding vector and the sensing power coefficient which costs $12L_{d}MN_{\rm tx}$ real multiplications/divisions.
-Let $C_{\rm prec-comp}^{\rm s}$ be the number of real multiplications/divisions to compute the sensing precoder $\textbf{w}_{0}$.
-The unitary matrix U in (19) (subspace spanned by the UE channel estimation vectors) is already obtained when computing the RZF precoding vectors by matrix inversion and the corresponding $\mathbf{L}\mathbf{D}\mathbf{L}^{H}$ decomposition [28, Lem.
+In this subsection, we present the GOPS analysis for sensing operations, including signal transmission and processing the received signal.
+The GOPS corresponding to the sensing transmissions consists of computing the sensing precoding vector $\textbf{w}_{0}$, denoted by $C_{\rm prec-comp}^{\rm s}$, and obtaining the sensing signal by multiplying the sensing symbols by the sensing precoding vector and the sensing power coefficient.
+Multiplying the sensing symbols by the sensing precoding vector and the sensing power coefficient costs $12L_{d}MN_{\rm tx}$ real multiplications/divisions.
+The unitary matrix U in (24) (subspace spanned by the UE channel estimation vectors) is already obtained when computing the RZF precoding vectors by matrix inversion and the corresponding $\mathbf{L}\mathbf{D}\mathbf{L}^{H}$ decomposition [36, Lem.
 B.2].
@@
 However, test statistics should be obtained for each transmission.
-We use the maximum a posteriori ratio test (MAPRT) detector proposed in [25, Lem.
-2].
-The test statistics is given by
-$T$ $=\frac{1}{\sigma_{n}^{2}}\left(\begin{bmatrix}\textbf{a}^{H}&\textbf{b}^{H}\end{bmatrix}\left(\begin{bmatrix}\textbf{C}&\textbf{E}\\ \textbf{E}^{H}&\textbf{D}\end{bmatrix}^{-1}-\begin{bmatrix}\textbf{0}&\textbf{0}\\ \textbf{0}&\textbf{D}^{-1}\end{bmatrix}\right)\begin{bmatrix}\textbf{a}\\ \textbf{b}\end{bmatrix}\right),$ (54)
-where
-$\textbf{a}=\sum_{m=1}^{L_{d}}\textbf{G}^{H}[m]\textbf{y}[m],\quad\textbf{b}=\sum_{m=1}^{L_{d}}\textbf{X}^{H}[m]\textbf{y}[m],$ (55)
-$\textbf{C}=\sum_{m=1}^{L_{d}}\textbf{G}^{H}[m]\textbf{G}[m]+\sigma_{n}^{2}\textbf{I}_{N_{\rm rx}N_{\rm tx}},$ (56)
-$\textbf{D}=\sum_{m=1}^{L_{d}}\textbf{X}^{H}[m]\textbf{X}[m]+\sigma_{n}^{2}\textbf{R}^{-1},$ (57)
-$\textbf{E}=\sum_{m=1}^{L_{d}}\textbf{G}^{H}[m]\textbf{X}[m].$ (58)
-Let $C_{\rm se-comp}$ denote the number of real multiplications/divisions to compute the test statistics $T$. The computational complexity can be expressed as
-$C_{\rm se-comp}=C_{\rm se-prep}+C_{\rm se-detector}$ (59)
-where $C_{\rm se-prep}$ and $C_{\rm se-detector}$ are the computational complexity of pre-processing and computing the test statistics for sensing, respectively.
-The former corresponds to computing (55)-(58), and the latter corresponds to computing the test statistics in (54).
-The computational complexity for pre-processing part and each step of computing the test statistics are listed in Table I.
-The total $C_{\rm se-prep}$ and $C_{\rm se-detector}$ are obtained as
-$C_{\rm se-prep}=L_{d}C_{{\rm se-prep},m}$
-$L_{d}\Bigg{(}20MN_{\rm tx}N_{\rm rx}+16M^{2}N_{\rm tx}N_{\rm rx}+4N_{\rm rx}\left(N_{\rm tx}^{2}+N_{\rm tx}\right)M$
-$\quad+4\left((MN_{\rm tx})^{2}+MN_{\rm tx}\right)+8N_{\rm rx}\,N_{\rm tx}^{2}\,M^{2}\Bigg{)}$ (60)
-$C_{\rm se-detector}=8\frac{\left((1+M^{2})N_{\rm tx}N_{\rm rx}\right)^{3}-(1+M^{2})N_{\rm tx}N_{\rm rx}}{3}$
-$\quad+8\frac{(M^{2}N_{\rm tx}N_{\rm rx})^{3}-M^{2}N_{\rm tx}N_{\rm rx}}{3}$
-$\quad+8\,\left(\left((1+M^{2})N_{\rm tx}N_{\rm rx}\right)^{2}+(1+M^{2})N_{\rm tx}N_{\rm rx}\right),$ (61)
-where $C_{{\rm se-prep},m}$ is defined as $C_{\rm se-prep}/L_{d}$. Finally, the sensing GOPS is obtained as
-$C_{\rm proc}^{\rm s}$ $=\frac{B}{L10^{9}}\Bigg{(}\underbrace{12L_{d}MN_{\rm tx}}_{\textrm{precoding + power}}+C_{\textrm{prec-comp}}^{\rm s}+\underbrace{C_{\rm se-comp}}_{\textrm{target detection}}\Bigg{)}.$ (62)
-TABLE I: Number of real multiplications/divisions for computing test statistics.
-Operation # of real multiplications/division
-$\textbf{G}[m],\quad\forall m$ $20\,L_{d}\,MN_{\rm tx}N_{\rm rx}$
-a $8\,L_{d}\,M^{2}\,N_{\rm rx}\,N_{\rm tx}\,$
-b $8\,L_{d}\,M^{2}\,N_{\rm tx}\,N_{\rm rx}$
-C $4\,L_{d}N_{\rm rx}\left(N_{\rm tx}^{2}+N_{\rm tx}\right)M$
-D $4L_{d}\,\left((MN_{\rm tx})^{2}+MN_{\rm tx}\right)$
-E $8\,L_{d}\,N_{\rm rx}\,N_{\rm tx}^{2}\,M^{2}$
-$\begin{bmatrix}\textbf{C}&\textbf{E}\\ \textbf{E}^{H}&\textbf{D}\end{bmatrix}^{-1}$ $8\frac{\left((1+M^{2})N_{\rm tx}N_{\rm rx}\right)^{3}-(1+M^{2})N_{\rm tx}N_{\rm rx}}{3}$
-$\textbf{D}^{-1}$ $8\frac{(M^{2}N_{\rm tx}N_{\rm rx})^{3}-M^{2}N_{\rm tx}N_{\rm rx}}{3}$
-$\begin{bmatrix}\textbf{a}^{H}&\textbf{b}^{H}\end{bmatrix}\begin{bmatrix}(.)&(.)\\ (.)&(.)\end{bmatrix}\begin{bmatrix}\textbf{a}\\ \textbf{b}\end{bmatrix}$ $8\,\left(\left((1+M^{2})N_{\rm tx}N_{\rm rx}\right)^{2}+(1+M^{2})N_{\rm tx}N_{\rm rx}\right)$
+Hence, the sensing GOPS is obtained as
+$C_{\rm proc}^{\rm s}$ $=\frac{B}{L10^{9}}\Bigg{(}\underbrace{12L_{d}MN_{\rm tx}}_{\textrm{precoding + power}}+C_{\textrm{prec-comp}}^{\rm s}+\underbrace{C_{\rm se-comp}}_{\textrm{target detection}}\Bigg{)},$ (49)
+where $C_{\rm se-comp}$ denotes the computational complexity for computing the detector test statistics. Given the clutter-unaware detector in (17), the computational complexity is computed as
+$C_{\rm se-comp}^{\rm c-unaware}$ $=L_{d}\Bigg{(}\underbrace{20MN_{\rm tx}N_{\rm rx}}_{\rm{for}\,\textbf{G}}+\underbrace{8\,M^{2}\,N_{\rm rx}\,N_{\rm tx}}_{\rm{for}\,\textbf{a}}$
+$\quad+\underbrace{4\,N_{\rm rx}\left(N_{\rm tx}^{2}+N_{\rm tx}\right)M}_{\rm{for}\,\textbf{C}}\Bigg{)}+\underbrace{8\frac{(N_{\rm tx}^{3}-N_{\rm tx})N_{\rm rx}}{3}}_{\textrm{for inverse {C}}}$
+$\quad+\underbrace{8\left((N_{\rm tx}N_{\rm rx})^{2}+N_{\rm tx}N_{\rm rx}\right)}_{\textrm{for final step}},$ (50)
+where we used the block diagonal structure of C in getting the inverse of it.
+The computational complexity for each step of computing the test statistics of clutter-aware detector in (18) are listed in Table I.
+TABLE I: Computational complexity with clutter-aware MAPRT detector
+Operation Computational complexity
+$\textbf{G}[m],\quad\forall m$ $\mathcal{O}\left(20\,L_{d}\,MN_{\rm tx}N_{\rm rx}\right)$
+a $\mathcal{O}\left(8\,L_{d}\,M^{2}\,N_{\rm rx}\,N_{\rm tx}\right)\,$
+b $\mathcal{O}\left(8\,L_{d}\,M^{2}\,N_{\rm rx}\,N_{\rm tx}\,\right)$
+C $\mathcal{O}\left(4\,L_{d}MN_{\rm rx}\left(N_{\rm tx}^{2}+N_{\rm tx}\right)\right)$
+D $\mathcal{O}\left(4L_{d}\,\left((MN_{\rm tx})^{2}+MN_{\rm tx}\right)\right)$
+E $\mathcal{O}\left(8\,L_{d}\,M^{2}\,N_{\rm rx}\,N_{\rm tx}^{2}\right)$
+$\begin{bmatrix}\textbf{C}&\textbf{E}\\ \textbf{E}^{H}&\textbf{D}\end{bmatrix}^{-1}$ $\mathcal{O}\left(\frac{8}{3}\left(\left((1+M^{2})N_{\rm tx}N_{\rm rx}\right)^{3}-(1+M^{2})N_{\rm tx}N_{\rm rx}\right)\right)$
+$\textbf{D}^{-1}$ $\mathcal{O}\left(\frac{8}{3}\left((M^{2}N_{\rm tx}N_{\rm rx})^{3}-M^{2}N_{\rm tx}N_{\rm rx}\right)\right)$
+$\begin{bmatrix}\textbf{a}^{H}&\textbf{b}^{H}\end{bmatrix}\begin{bmatrix}(.)&(.)\\ (.)&(.)\end{bmatrix}\begin{bmatrix}\textbf{a}\\ \textbf{b}\end{bmatrix}$ $\mathcal{O}\left(8\,\left(\left((1+M^{2})N_{\rm tx}N_{\rm rx}\right)^{2}+(1+M^{2})N_{\rm tx}N_{\rm rx}\right)\right)$
 VI Joint blocklength and Power Optimization
@@
 However, in our analysis, where the primary focus is on the interplay between processing and radio resources, we treat fronthaul energy consumption as a fixed component and do not include it in our considerations.
+The corresponding algorithms to these optimization problems are represented by SeURLLC+ and E2E SeURLLC+, respectively.
 VI-A Minimizing Energy Consumption for Transmission
@@
 $\overline{\mathsf{SINR}}_{s}\geq\gamma_{\rm s},\quad$ (51d)
-$P_{k}\leq P_{\rm tx},\quad\quad\textrm{for}\quad k=1,\cdots,N_{\rm tx}$ (63e)
-$P_{\rm tr}=\sum_{k=1}^{N_{\rm tx}}P_{k}$ is the average power consumption for transmission, given as $P_{\rm tr}=\sum_{j=0}^{N_{\rm ue}}\rho_{j}=\|\boldsymbol{\rho}\|^{2}$, where $\boldsymbol{\rho}=[\sqrt{\rho_{0}}\ \ldots\ \sqrt{\rho_{N_{\rm ue}}}]^{T}$, due to the unit-power centralized precoding vectors.
+$P_{k}\leq P_{\rm tx},\quad\quad k=1,\cdots,N_{\rm tx}$ (51e)
+$P_{\rm tr}=\sum_{k=1}^{N_{\rm tx}}P_{k}$ is the average power consumption for transmission, given as $P_{\rm tr}=\sum_{j=0}^{N_{\rm ue}}\rho_{j}=\|\bm{\rho}\|^{2}$, where $\bm{\rho}=[\sqrt{\rho_{0}}\ \ldots\ \sqrt{\rho_{N_{\rm ue}}}]^{T}$, due to the unit-power centralized precoding vectors.
 Constraints (51b) and (51c) correspond to the URLLC requirements.
@@
 Theorem 1.
-Consider the problem in (VI-A) and define the collective vectors $\textbf{r}=[r_{1}\ \ldots\ r_{N_{\rm ue}}]^{T}$ and $\boldsymbol{\chi}=[\chi_{1}\ \ldots\ \chi_{N_{\rm ue}}]^{T}$.
-The optimal solution $\{\boldsymbol{\rho}^{\star},L^{\star}\}$ of the problem given below is also an optimal solution to (VI-A):
-$\underset{\boldsymbol{\rho},\boldsymbol{\chi},\textbf{r}\geq\textbf{0},L>L_{p},\bar{L}>0}{\textrm{minimize}}\quad\frac{\|\boldsymbol{\rho}\|^{2}}{\overline{L}}\,$ (64a)
+Consider the problem in (51) and define the collective vectors $\textbf{r}=[r_{1}\ \ldots\ r_{N_{\rm ue}}]^{T}$ and $\bm{\chi}=[\chi_{1}\ \ldots\ \chi_{N_{\rm ue}}]^{T}$.
+The optimal solution $\{\bm{\rho}^{\star},L^{\star}\}$ of the problem given below is also an optimal solution to (51):
+$\underset{\bm{\rho},\bm{\chi},\textbf{r}\geq\textbf{0},L>L_{p},\bar{L}>0}{\textrm{minimize}}\quad\frac{\|\bm{\rho}\|^{2}}{\overline{L}}\,$ (52a)
 […]
 $\ln\left(1+\chi_{i}\right)-\chi_{i}+\frac{\rho_{i}\mathsf{b}_{i}^{2}}{r_{i}}\geq\frac{Q^{-1}\left(\epsilon_{i}^{\rm(th)}\right)}{\sqrt{L-L_{p}}}+\frac{b_{i}\ln{2}}{L-L_{p}}$ (52c)
-$\left\|\begin{bmatrix}\sqrt{2}\mathsf{a}_{i,0}\sqrt{\rho_{0}}\\ \vdots\\ \sqrt{2}\mathsf{a}_{i,N_{\rm ue}}\sqrt{\rho_{N_{\rm ue}}}\\ \sqrt{2}\mathsf{b}_{i}\sqrt{\rho_{i}}\\ \sqrt{2}\sigma_{n}\\ 1+\chi_{i}\\ r_{i}\end{bmatrix}\!\right\|\leq 1+\chi_{i}+r_{i},\forall i,$ (64k)
-$\boldsymbol{\rho}^{T}\left(\gamma_{\rm s}\textbf{B}_{D}-M\,\textbf{A}_{D}\,\right)\boldsymbol{\rho}\leq-\gamma_{\rm s}MN_{\rm rx}\sigma_{n}^{2}$ (64l)
-$\left\|\textbf{F}_{k}\boldsymbol{\rho}\right\|\leq\sqrt{P_{\rm tx}},\quad k=1,\ldots,N_{\rm tx},$ (64m)
-$L\leq L_{\rm max},$ (64n)
+$\left\|\begin{bmatrix}\sqrt{2}\mathsf{a}_{i,0}\sqrt{\rho_{0}}\\ \vdots\\ \sqrt{2}\mathsf{a}_{i,N_{\rm ue}}\sqrt{\rho_{N_{\rm ue}}}\\ \sqrt{2}\mathsf{b}_{i}\sqrt{\rho_{i}}\\ \sqrt{2}\sigma_{n}\\ 1+\chi_{i}\\ r_{i}\end{bmatrix}\!\right\|\leq 1+\chi_{i}+r_{i},\forall i,$ (52d)
+$\bm{\rho}^{T}\left(\gamma_{\rm s}\textbf{B}_{D}-M\,\textbf{A}_{D}\,\right)\bm{\rho}\leq-\gamma_{\rm s}MN_{\rm rx}\sigma_{n}^{2}$ (52e)
+$\left\|\textbf{F}_{k}\bm{\rho}\right\|\leq\sqrt{P_{\rm tx}},\quad k=1,\ldots,N_{\rm tx},$ (52f)
+$L\leq L_{\rm max},$ (52g)
 where ${\bf F}_{k}=\mathrm{diag}\left(\sqrt{\mathbb{E}\left\{\|{\bf w}_{0,k}\|^{2}\right\}},\ldots,\sqrt{\mathbb{E}\left\{\|{\bf w}_{N_{\rm ue},k}\|^{2}\right\}}\right)$.
@@
 See Appendix A. ∎
-The optimization problem in (1) is still not convex due to the non-convex constraints (64b), (64c) and (64l).
-The terms that destroy convexity are the convex terms $1/\overline{L}$ and $\frac{\rho_{i}\mathsf{b}_{i}^{2}}{r_{i}}$ (in terms of $\boldsymbol{\rho}$ and r) on the right-hand side of (64b) and the left-hand side of (64c), respectively.
-To this end, we apply the concave-convex procedure (CCP) approach to (64b) and (64c), and the FPP-SCA method [34] to (64l), wherein $-\boldsymbol{\rho}^{T}M\,\textbf{A}_{D}\boldsymbol{\rho}$ is a concave function.
-Moreover, to avoid any potential infeasibility issue regarding (64l) during the initial iterations of the algorithm, we add slack variable $\chi_{0}\geq 0$ and a slack penalty $\lambda$, to the convexified problem at the initial iterations.
+The optimization problem in (52) is still not convex due to the non-convex constraints (52b), (52c) and (52e).
+The terms that destroy convexity are the convex terms $1/\overline{L}$ and $\frac{\rho_{i}\mathsf{b}_{i}^{2}}{r_{i}}$ (in terms of $\bm{\rho}$ and r) on the right-hand side of (52b) and the left-hand side of (52c), respectively.
+To this end, we apply the concave-convex procedure (CCP) approach to (52b) and (52c), and the FPP-SCA method [42] to (52e), wherein $-\bm{\rho}^{T}M\,\textbf{A}_{D}\bm{\rho}$ is a concave function.
+Moreover, to avoid any potential infeasibility issue regarding (52e) during the initial iterations of the algorithm, we add slack variable $\chi_{0}\geq 0$ and a slack penalty $\lambda$, to the convexified problem at the initial iterations.
 In subsequent iterations, we set $\chi_{0}$ to zero if it is less than a threshold, denoted as $\chi_{0}\leq\epsilon_{\chi}$.
 Finally, the convex problem that is solved at the $c^{\rm th}$ iteration becomes
-$\underset{\boldsymbol{\rho},\boldsymbol{\chi},\textbf{r}\geq\textbf{0},L,\bar{L}>L_{p},\chi_{0}\geq 0}{\textrm{minimize}}\quad\frac{\|\boldsymbol{\rho}\|^{2}}{\overline{L}}+\lambda\chi_{0}\,,$ (65a)
+$\underset{\bm{\rho},\bm{\chi},\textbf{r}\geq\textbf{0},L,\bar{L}>L_{p},\chi_{0}\geq 0}{\textrm{minimize}}\quad\frac{\|\bm{\rho}\|^{2}}{\overline{L}}+\lambda\chi_{0}\,,$ (53a)
 subject $\textrm{subject to}\quad{L-L_{p}}\leq\frac{2}{\overline{L}^{(c-1)}}-\frac{\overline{L}}{\left(\overline{L}^{(c-1)}\right)^{2}},$ (53b)
@@
 $\hskip 14.22636pt\geq\frac{Q^{-1}\left(\epsilon_{i}^{\rm(th)}\right)}{\sqrt{L-L_{p}}}+\frac{b_{i}\ln{2}}{L-L_{p}}$ (53c)
-$\gamma_{\rm s}\boldsymbol{\rho}^{T}\textbf{B}_{D}\boldsymbol{\rho}-2M\,\Re\left(\left(\boldsymbol{\rho}^{(c-1)}\right)^{T}\textbf{A}_{D}\boldsymbol{\rho}\right)$
-$\hskip 14.22636pt\leq-\gamma_{\rm s}MN_{\rm rx}\sigma_{n}^{2}-M\left(\boldsymbol{\rho}^{(c-1)}\right)^{T}\textbf{A}_{D}\boldsymbol{\rho}^{(c-1)}+\chi_{0},$ (65d)
+$\gamma_{\rm s}\bm{\rho}^{T}\textbf{B}_{D}\bm{\rho}-2M\,\Re\left(\left(\bm{\rho}^{(c-1)}\right)^{T}\textbf{A}_{D}\bm{\rho}\right)$
+$\hskip 14.22636pt\leq-\gamma_{\rm s}MN_{\rm rx}\sigma_{n}^{2}-M\left(\bm{\rho}^{(c-1)}\right)^{T}\textbf{A}_{D}\bm{\rho}^{(c-1)}+\chi_{0},$ (53d)
 […]
@@
 Algorithm 1 FPP-SCA and CCP Procedure for Problem (52).
-1:Initialization: Set arbitrary $\boldsymbol{\rho}^{(0)},\textbf{r}^{(0)}\geq\textbf{0}$, the solution accuracy parameters $\epsilon,\epsilon_{\chi}>0$, $\overline{L}^{(0)}=\frac{1}{L_{\rm max}-L_{p}}$, and $\lambda>0$.
+1:Initialization: Set arbitrary $\bm{\rho}^{(0)},\textbf{r}^{(0)}\geq\textbf{0}$, the solution accuracy parameters $\epsilon,\epsilon_{\chi}>0$, $\overline{L}^{(0)}=\frac{1}{L_{\rm max}-L_{p}}$, and $\lambda>0$.
 Set the iteration counter to $c=0$, and a maximum iteration number $c_{\rm max}$.
@@
 3:While $\Delta\mathsf{F}^{(c-1)}\geq\epsilon$ and $c\leq c_{\rm max}$, do
-4:Set $\boldsymbol{\rho}^{(c)}$, $\overline{L}^{(c)}$, and $\textbf{r}^{(c)}$ to the solution of the convex problem in (VI-A), where the previous iterates $\boldsymbol{\rho}^{(c-1)}$, $\overline{L}^{(c-1)}$, and $\textbf{r}^{(c-1)}$ are taken as constant.
+4:Set $\bm{\rho}^{(c)}$, $\overline{L}^{(c)}$, and $\textbf{r}^{(c)}$ to the solution of the convex problem in (53), where the previous iterates $\bm{\rho}^{(c-1)}$, $\overline{L}^{(c-1)}$, and $\textbf{r}^{(c-1)}$ are taken as constant.
 5:If $\chi_{0}<\epsilon_{\chi}$, set $\chi_{0}=0$ for the next iteration.
 6:End while
-7:Output: The transmit power coefficients $\boldsymbol{\rho}^{(c)}$ and the blocklength $L^{(c)}$.
+7:Output: The transmit power coefficients $\bm{\rho}^{(c)}$ and the blocklength $L^{(c)}$.
 VI-B Minimizing E2E Energy Consumption
@@
 […]
-subject to $\epsilon_{i}\leq\epsilon^{\rm(ub)}_{i}\leq\epsilon^{\rm(th)}_{i},\quad\forall i>0$ (66b)
-$L\leq L_{\rm max},\quad$ (66c)
-$\overline{\mathsf{SINR}}_{s}\geq\gamma_{\rm s},\quad$ (66d)
-$P_{k}\leq P_{\rm tx},\quad\quad\textrm{for}\quad k=1,\cdots,N_{\rm tx}.$ (66e)
+[…]
 The total power consumption can be rewritten as
-$P_{\rm total}=P_{\rm FIXED}+\Delta^{\rm tr}\|\boldsymbol{\rho}\|^{2}+\frac{1}{L}f_{1}+\frac{L_{d}}{L}f_{2}$ (67)
+$P_{\rm total}=P_{\rm FIXED}+\Delta^{\rm tr}\|\bm{\rho}\|^{2}+\frac{1}{L}f_{1}+\frac{L_{d}}{L}f_{2}$ (55)
 where
@@
 Similar to the approach of problem P1, we define a new variable $\bar{L}$ where $L-L_{p}\leq\frac{1}{\bar{L}}$. Then, the objective function (54a) is equivalent to
-$\mathsf{F_{2}}\triangleq\frac{1}{B}\left(L\left(P_{\rm FIXED}\right)+\left(L-L_{p}\right)f_{2}+\Delta^{\rm tr}\frac{\|\boldsymbol{\rho}\|^{2}}{\bar{L}}\right).$ (71)
-The non-convex optimization problem in (VI-B) can be solved using the same approach in P1 as follows:
-$\underset{\boldsymbol{\rho},\boldsymbol{\chi},\textbf{r}\geq\textbf{0},L,\bar{L}>L_{p},\chi_{0}\geq 0}{\textrm{minimize}}\quad\mathsf{F_{2}}+\lambda\chi_{0},$ (72a)
+$\mathsf{F_{2}}\triangleq\frac{1}{B}\left(L\left(P_{\rm FIXED}\right)+\left(L-L_{p}\right)f_{2}+\Delta^{\rm tr}\frac{\|\bm{\rho}\|^{2}}{\bar{L}}\right).$ (59)
+The non-convex optimization problem in (54) can be solved using the same approach in P1 as follows:
+$\underset{\bm{\rho},\bm{\chi},\textbf{r}\geq\textbf{0},L,\bar{L}>L_{p},\chi_{0}\geq 0}{\textrm{minimize}}\quad\mathsf{F_{2}}+\lambda\chi_{0},$ (60a)
 […]
@@
 In this section, we present numerical results to evaluate the performance of the proposed joint blocklength and power allocation algorithms.
-We consider a total area of 500 $\rm m\times$ 500 $\rm m$.
-The sensing location is in the center of the area, i.e., $(250,250)$.
-The number of antenna elements per AP is set to $M=4$.
+We consider a total area of 500 $\rm m\times$ 500 $\rm m$, with the sensing location positioned at the center.
 The $N_{\rm tx}=16$ ISAC transmit APs are uniformly distributed in the area.
 We consider the number of sensing receive APs equal to $N_{\rm rx}=1$ or $N_{\rm rx}=2$ where the first AP is located at $(200,250)$ and the second one is located at $(300,250)$.
-The total number of URLLC UEs in the network is $N_{\rm ue}=8$, unless otherwise stated.
+The number of antenna elements per AP is set to $M=4$.
+The total number of URLLC UEs in the network is $N_{\rm ue}=8$.
 We set $P_{\rm tx}=100$ mW and the uplink pilot transmission power for each UE is $50$ mW.
-The large-scale fading coefficients, shadowing parameters, probability of LOS, and the Rician factors are simulated based on the 3GPP Urban Microcell model, defined in [35, Table B.1.2.1-1, Table B.1.2.1-2, Table B.1.2.2.1-4].
+The large-scale fading coefficients, shadowing parameters, probability of LOS, and the Rician factors are simulated based on the 3GPP Urban Microcell model, defined in [43, Table B.1.2.1-1, Table B.1.2.1-2, Table B.1.2.2.1-4].
 The path losses for the Rayleigh fading target-free channels are also modeled by the 3GPP Urban Microcell model with the difference that the channel gains are multiplied by an additional scaling parameter equal to $0.3$ to suppress the known parts of the target-free channels due to LOS and permanent obstacles [34].
@@
 The number of pilot symbols is $L_{p}=10$.
+The spatial correlation matrices for the communication channels are generated by using the local scattering model in [35, Sec.
+2.5.3].
+The RCS of the target is modeled by the Swerling-I model with $\sigma_{\rm rcs}=0\,\rm dBsm$.
+For all the UEs, the packet size, maximum transmission delay, and DEP threshold are $b_{i}=256$ bits, $D_{i}^{\rm(th)}=1$ ms, and $\epsilon_{\rm th}=10^{-5}$, respectively.
+In the proposed algorithms, $\epsilon=10^{-3}$, $\epsilon_{\chi}=10^{-6}$, $\lambda=10$, and $c_{\rm max}=30$.
+The sensing SINR threshold is set to $\gamma_{s}=0$ dB, unless otherwise stated.
+The remaining parameters are detailed in Table II, where the values are consistent with those in [30], except for a scaling of the total processing capacity and the corresponding power of a DU by a factor of ten to accommodate all the processing in a single DU.
 TABLE II: Simulation Parameters
@@
 $\Delta_{\rm cloud}^{\rm proc}$ 740 W $C_{\rm max}$ 1800 GOPS
-The spatial correlation matrices for the communication channels are generated by using the local scattering model in [27, Sec.
-2.5.3].
-The RCS of the target is modeled by the Swerling-I model with $\sigma_{\rm rcs}=0\,\rm dBsm$.
-For all the UEs, we set the same packet size, the maximum transmission delay, and the DEP threshold to $b_{i}=256$ bits, $D_{i}^{\rm(th)}=1$ ms, and $\epsilon_{\rm th}=10^{-5}$, respectively unless otherwise stated.
-In the proposed algorithms, $\epsilon=10^{-3}$, $\epsilon_{\chi}=10^{-6}$, $\lambda=10$, and $c_{\rm max}=30$.
-The sensing SINR threshold is set to $\gamma_{s}=0$ dB, unless otherwise stated.
-The remaining parameters are detailed in Table II, where the values are consistent with those in [14], except for a scaling of the total processing capacity and the corresponding power of a DU by a factor of ten to accommodate all the processing in a single DU.
-We compare the performance of two proposed algorithms: i) SeURLLC+, which aims to minimize the transmission energy consumption for as ISAC cell-free massive MIMO system with URLLC UEs, and ii) E2E SeURLLC+, which aims to minimize the E2E energy consumption.
-We also consider two baseline algorithms for only URLLC systems without integrating sensing: i) URLLC-only, and ii) E2E URLLC-only, which represent minimizing transmission energy and E2E energy, respectively.
-Figure 2: E2E energy consumption over one transmission with different algorithms and with $N_{\rm rx}=1,2$.
-In Fig.
-2, we illustrate E2E energy consumption with varying algorithms and two values for the number of sensing receive APs.
-The chart breaks down energy consumption by system components and different operations, including transmission, ISAC APs, sensing receive APs, communication and sensing processing, and “Others” representing load-independent and idle mode power consumption at the cloud.
-It is shown that in URLLC-only scenarios, the majority of the energy consumption is related to the ISAC APs and then the load-independent and idle mode power consumption at the cloud.
-Remarkably, E2E energy minimization demonstrates a substantial reduction in energy consumption by approximately $50\%$ in URLLC-only scenarios.
-This underscores the importance of E2E energy-aware optimization.
-Conversely, in ISAC scenarios, the impact of E2E energy minimization is less pronounced, resulting in only a marginal decrease.
-However, we observe that the energy consumption due to sensing processing becomes predominant when $N_{\rm rx}=2$.
-This observation suggests a potential optimization avenue, as turning off one sensing receive AP can halve the energy consumption, contributing significantly to overall energy efficiency.
-(a) E2E URLLC-only
-(b) E2E SeURLLC+ , $N_{\rm rx}=1$
-(c) E2E SeURLLC+ , $N_{\rm rx}=2$
-Figure 3: Energy consumption in one transmission with a) E2E URLLC-only, b) E2E SeURLLC+ with $N_{\rm rx}=1$, and E2E SeURLLC+ with $N_{\rm rx}=2$.
-To complete the findings in Fig.
-2, we provide a comparative analysis of the energy consumption distribution across various components within the system in Fig.
-3.
-E2E energy minimization algorithms are considered for a) URLLC-only scenario, b) ISAC scenario with 1 sensing receive AP, and c) ISAC scenario with 2 sensing receive APs.
-As shown in Fig.
-2(a) and Fig.
-2(b), the ISAC APs are the most energy consuming components.
-Notably, the increment in the number of sensing receive APs is accompanied by a significant rise in energy consumption corresponding to the processing of the received signals for target detection.
-This is evident that with the addition of another receive AP, the share of sensing processing increases up to $52\%$, approximately four times larger than the case with $N_{\rm rx}=1$.
-Figure 4: Blocklength vs. decoding error probability threshold.
-Figure 5: E2E energy consumption vs. decoding error probability threshold.
-VII-A Impact of Reliability and Delay Requirements
-In this subsection, we investigate the impact of the reliability and delay requirements on the performance of the above-mentioned algorithms.
+Figure 2: (a) Transmission energy consumption, (b) E2E energy consumption, and (c) breakdown of energy consumption over one transmission for different optimization algorithms, for clutter-aware target detector with $N_{\rm rx}=2$ and $\gamma_{s}=0$ dB.
+We compare the performance of two proposed algorithms: i) SeURLLC+ aiming to minimize the transmission energy consumption, and ii) E2E SeURLLC+ aiming to minimize the E2E energy consumption.
+Additionally, we include two baseline algorithms for URLLC systems without sensing integration: (i) URLLC-only, which minimizes transmission energy consumption, and (ii) E2E URLLC-only, which minimizes E2E energy consumption.
+We further compare results using clutter-aware and clutter-unaware MAPRT detectors, as defined in (17) and (18), respectively.
 Fig.
-4 shows the blocklength as a function of decoding error probability (DEP) threshold.
-Generally, the blocklength decreases by relaxing the reliability threshold in all the algorithms except URLLC-only.
-It is shown that the URLLC-only algorithm always ends up with the maximum available blocklength, i.e., $L_{\rm max}$, since increasing the blocklength allows lower transmit power to satisfy the reliability constraint.
-We also observe that minimizing only the transmission energy consumption will lead to higher blocklength values compared to the E2E energy minimization algorithms, i.e., E2E URLLC-only and E2E SeURLLC+.
-Compared to URLLC-only, the E2E URLLC-only algorithm results in a significant decrease in the blocklength to minimize the E2E energy consumption over one transmission.
-For ISAC scenarios, we see the same behaviours as in E2E URLLC-only.
-However, the blocklength is slightly higher due to the interplay between sensing power consumption and the blocklength as transmit power consumption will decrease with larger blocklength.
+2(a) illustrates the transmission energy consumption for four algorithms with the clutter-aware detector.
+The results show that integrating sensing into the system necessitates higher transmission energy.
+Notably, the URLLC-only and SeURLLC+ achieve lower transmission energy compared to their E2E counterparts, as their primary objective is to minimize transmission energy consumption.
+The E2E energy consumption is depicted in Fig.
+2(b).
+We observe that E2E energy minimization results in a substantial reduction in total energy consumption by approximately $50\%$ in URLLC-only scenarios, while its impact is less pronounced in ISAC scenarios, resulting in only a marginal decrease.
+The E2E algorithms try to minimize the total energy consumption by reducing the blocklength.
+However, higher power coefficients are required to meet the reliability requirements for URLLC UEs.
+In SeURLLC+, higher power coefficients for sensing allow lower blocklength, narrowing the gap in optimal blocklengths between SeURLLC+ and E2E SeURLLC+.
 Fig.
-5 illustrates the E2E energy consumption in terms of the DEP threshold.
-The results show that relaxing the reliability constraint would generally reduce the E2E energy consumption.
-However, ISAC scenarios demand higher energy consumption as it is expected.
-Moreover, we observe that algorithms that minimize E2E energy consumption outperform the corresponding algorithms which only minimize the transmission energy consumption.
-Particularly, when we do not integrate sensing in the system, there is a noticeable gap between performance of URLLC-only and E2E URLLC-only algorithms.
-Figure 6: E2E energy consumption vs. maximum delay threshold.
-The results for E2E energy consumption versus maximum delay threshold, $D_{i}^{(\rm tx)}$, are depicted in Fig.
-6.
-As it is shown, E2E energy consumption does not change significantly with different maximum delay threshold, except for URLLC-only algorithm.
-The reason is that the URLLC-only algorithm would always result in the highest possible blocklength to minimize the transmission energy consumption while the optimized value for the blocklength with other algorithms would be much lower than the one with URLLC-only to minimize the energy consumption.
-That is why the SeURLLC+ has lower E2E energy consumption than URLLC-only.
-However, the results of E2E URLLC-only algorithm show that minimum E2E energy consumption could be archived by optimizing the blocklength with respect to E2E energy consumption.
-This behavior highlights the importance of the E2E energy minimization.
+2c provides a detailed breakdown of energy consumption by system components and different operations, including ISAC transmit APs, sensing receive APs, communication and sensing processing, and “Others,” which represent load-independent and idle mode power consumption in the cloud.
+For both algorithms, most of the energy consumption is attributed to sensing processing tasks, ISAC APs, and load-independent and idle mode power consumption in the cloud.
+However, the E2E SeURLLC+ algorithm slightly reduces the corresponding energy consumption.
+Figure 3: (a) Blocklength, (b) E2E energy consumption, (c) detection probability, and (d) sensing update rate vs.
+decoding error probability threshold with clutter-aware target detector, $N_{\rm rx}=2$ and $\gamma_{s}=0$ dB.
+VII-A Impact of URLLC Requirements
+Fig.
+3 illustrates the blocklength, E2E energy consumption, detection probability and sensing update rate as functions of DEP threshold.
+Generally, the blocklength in Fig.
+3(a) decreases as the reliability threshold is relaxed in all algorithms except for URLLC-only.
+The URLLC-only algorithm consistently uses the maximum available blocklength, i.e., $L_{\rm max}$.
+By integrating sensing, the optimal blocklength is significantly reduced compared to URLLC-only, as meeting sensing constraint requires higher power consumption resulting in lower blocklength.
+Relaxing the reliability constraint would generally reduce the E2E energy consumption.
+However, ISAC scenarios demand higher energy consumption as it is expected as shown in Fig.
+3(b).
+Fig.
+3(c) and Fig.
+3(d) demonstrate the impact of the DEP threshold on sensing performance.
+Relaxing the DEP constraint slightly reduces detection probability while increasing the sensing update rate, due to smaller blocklengths at higher DEP thresholds.
+Notably, the E2E SeURLLC+ algorithm outperforms the SeURLLC+ algorithm in both sensing and communication performance.
+It is worth mentioning that higher delay threshold allows the system to operate with higher blocklength.
+However, it does not effect the results since the algorithm aims to minimize the energy consumption.
+Therefore, it still chooses a smaller blocklength.
 VII-B Impact of Sensing Requirement
-Figure 7: Blocklength vs. sensing SINR threshold.
-Figure 8: E2E energy consumption vs. sensing SINR threshold.
-The impact of the sensing SINR threshold, $\gamma_{s}$, on the blocklength and E2E energy consumption is investigated in Fig.
-7 and Fig.
-8, respectively.
-E2E energy minimization leads to lower energy consumption as it is shown that E2E SeURLLC+ outperforms SeURLLC+.
-The reason can be found in Fig.
-7.
-It is shown that E2E SeURLLC+ leads to lower blocklength which means lower E2E energy consumption.
-Moreover, we observe that the gap between the performance of these two algorithms gradually decreases by demanding higher sensing SINR threshold.
+The impact of the sensing SINR threshold, $\gamma_{s}$, on the performance of the E2E SeURLLC+ algorithm is evaluated in Figs.
+4a-c in terms of E2E energy consumption, detection probability, and sensing update rate, respectively.
+Comparing Fig.
+4(a) with Fig.
+4(b), we observe that both detection probability and energy consumption generally increase with a higher SINR threshold.
+The clutter-aware detector with $N_{\rm rx}=2$ provides the highest detection probability, consistently above 0.9, whereas the clutter-unaware detector cannot achieve a detection probability over 0.6, despite offering the lowest energy consumption.
+Notably, using a clutter-aware detector while disabling one RX-AP maintains a detection probability above 0.8 with around 43% energy saving.
+This suggests a potential optimization strategy: turning off one sensing receive AP can halve the energy consumption, significantly enhancing overall energy efficiency.
+Higher sensing SINR threshold demands higher blocklength, resulting in lower sensing update rate as shown in Fig.
+4(c).
+Comparing Fig.
+4(b) and Fig.
+4(c), it is evident that increasing the sensing SINR threshold significantly decreases the sensing update rate without substantially improving detection probability.
+Figure 4: (a) E2E energy consumption, (b) detection probability, and (c) sensing update rate vs. sensing SINR threshold with E2E SeURLLC+ algorithm.
+VII-C Impact of Number of APs
+We explore the impact of the number of transmit and receive APs on the network performance in terms of E2E energy consumption, detection probability, and sensing update rate as shown in Figs.
+5a-c, respectively.
+The results in Fig.
+5(a) and Fig.
+5(b) indicate that increasing the number of receive APs generally raises both E2E energy consumption and detection probability.
+For $N_{\rm rx}=2$, the detection probability substantially improves with only a marginal raise in energy consumption.
+However, further increasing to $N_{\rm rx}=4$ offers little improvement in detection probability while significantly increasing energy consumption.
+Moreover, the results in Fig.
+5(c) show a consistent sensing update rate across all curves, especially for $N_{\rm tx}=8$ and $N_{\rm tx}=12$, indicating that $N_{\rm rx}$ has minimal impact on $R_{s}$.
+This suggests the optimization algorithm is more constrained by URLLC requirements than by sensing requirements.
+A comparison between $N_{\rm tx}=8$ and $N_{\rm tx}=12$ in Fig.
+5(c) reveals that a higher $N_{\rm tx}$ enhances URLLC satisfaction due to elevated SINR, resulting in a reduced blocklength and consequently a higher sensing update rate.
+It is observed that when $N_{\rm rx}=1$ or $N_{\rm rx}=2$, increasing the number of transmit APs does not significantly affect E2E energy consumption.
+This is because the rise in energy consumption from the additional APs is offset by the decrease in blocklength.
+Thus, the overall E2E energy consumption remains relatively unchanged.
+Conversely, when $N_{\rm rx}=4$, the influence of $N_{\rm tx}$ becomes more pronounced, as the energy consumption for sensing tasks increases non-linearly with the product of $N_{\rm tx}$ and $N_{\rm rx}$.
+When $N_{\rm tx}=16$, the system’s behavior concerning E2E energy consumption and detection probability follows a similar trend, showing an increase in energy consumption and a decrease in detection probability.
+However, $R_{s}$ declines compared to when $N_{\rm tx}=12$, indicating a higher blocklength for $N_{\rm tx}=16$.
+This occurs because common power coefficients are used to maintain a favorable beam pattern, and increasing the number of transmit APs may constrain power due to some APs being closer to the UEs.
+As a result, a higher blocklength is necessary to fulfill the reliability requirements.
+These findings underscore the importance of optimizing the number of transmit and receive APs.
+The results of detection probability is depicted in Fig.
+5(b).
+In general, a decrease in blocklength translates to a lower detection probability explaining the results for $N_{\rm tx}=12$.
+However, we observe that detection probability for $N_{\rm tx}=16$ still decreases although blocklength is higher compared to $N_{\rm tx}=12$.
+This is because the power coefficients required to satisfy the sensing SINR requirement might be lower with higher $N_{\rm tx}$, resulting in a more pronounced negative impact on detection probability.
+This negative impact outweighs the benefits of increased blocklength, leading to a lower detection probability.
+The findings emphasize the importance of optimizing the number of transmit and receive APs to balance energy consumption, detection probability and sensing update rate.
+Higher numbers of transmit APs can enhance URLLC satisfaction and reduce blocklength resulting in higher sensing update rate, but must be carefully managed to avoid an unnecessary increase in energy consumption and detection probability degradation.
+Figure 5: (a) E2E energy consumption, (b) detection probability, and (c) sensing update rate vs. the number of transmit ISAC APs for $N_{\rm rx}=1,2,4$ and E2E SeURLLC+ algorithm.
 VIII Conclusion
-In this paper, we proposed a joint blocklength and power control algorithm for a downlink cell-free massive MIMO system with multi-static sensing and URLLC users.
-We formulated two non-convex optimization problems, focusing on transmission energy minimization and E2E energy minimization, respectively.
-Numerical results demonstrated the superiority of the E2E energy minimization algorithm, particularly in scenarios with no sensing capability.
-Furthermore, our results revealed a growing significance of energy consumption related to sensing processing tasks with the increase in the number of sensing receive access points.
-This observation highlights a potential optimization avenue, as deactivating one sensing receive AP can effectively mitigate this dominance and contribute to overall energy efficiency.
-Furthermore, our findings indicated that a higher sensing SINR threshold leads to increased E2E energy consumption.
-This ultimately narrows the performance gap between the two proposed algorithms: the transmission-only approach and the E2E energy minimization approach.
+In this paper, we proposed a joint blocklength and power control algorithm for a downlink CF-mMIMO system with multi-static sensing and URLLC UEs.
+We formulated two non-convex optimization problems focused on minimizing transmission and E2E energy consumption, and conducted GOPS analysis for both communication and sensing tasks.
+Numerical results demonstrated the effectiveness of the E2E energy minimization algorithm, especially in scenarios without sensing capabilities.
+Integrating sensing reduces the optimal blocklength compared to URLLC-only scenarios due to the higher power demands of sensing, which leads to a lower optimal blocklength.
+Moreover, the E2E algorithms minimize total energy consumption by further reducing the blocklength.
+Additionally, we observed that higher sensing SINR thresholds significantly lower the sensing update rate without substantially improving detection probability.
+Our findings also indicated that increasing the number of sensing receive APs raises sensing-related processing energy, suggesting that deactivating some APs can help reduce E2E energy consumption in CF-mMIMO ISAC systems.
+While increasing the number of transmit APs can enhance URLLC satisfaction and sensing update rate by reducing blocklength, careful management is necessary to prevent excessive energy consumption and degradation in detection probability.
 Appendix A Proof of Theorem 1
 Let us define a new optimization variable, denoted by $\overline{L}$, where $L-L_{p}\leq 1/\overline{L}$.
-Then, minimizing the objective function is equivalent to minimizing the following convex function (quadratic-over-linear function)
-$\frac{\|\boldsymbol{\rho}\|^{2}}{\overline{L}}\,.$ (73)
-Since the above function is minimized, at the optimal solution, it holds that $\overline{L}=1/L_{d}=1/(L-L_{p})$. According to (26), the reliability constraints in (63b) can be written in the form of
+Then, minimizing the objective function is equivalent to minimizing the following convex function (quadratic-over-linear function) $\frac{\|\bm{\rho}\|^{2}}{\overline{L}}$ Minimizing this function, at the optimal solution, holds that $\overline{L}=1/L_{d}=1/(L-L_{p})$.
+The reliability constraints in (51b) can be written in the form of
 $\ln\left(1+\frac{\rho_{i}\mathsf{b}_{i}^{2}}{\sum_{j=0}^{{\rm N_{ue}}}\rho_{j}\mathsf{a}_{i,j}^{2}+\sigma_{n}^{2}}\right)\!\geq\frac{Q^{-1}\left(\epsilon_{i}^{(\rm th)}\right)}{\sqrt{L-L_{p}}}+\frac{b_{i}\ln{2}}{L-L_{p}}$ (61)
-where $\overline{\mathsf{SINR}}^{\rm(dl)}_{i}$ is substituted by (23).
+according to (31), where $\overline{\mathsf{SINR}}^{\rm(dl)}_{i}$ is substituted by (28).
 To handle the non-convexity of the left-hand side in (61), we define a new variable $\chi_{i}$ and use fractional programming [44] to write the left-hand side as
@@
 $\frac{\sum_{j=0}^{{\rm N_{ue}}}\rho_{j}\mathsf{a}_{i,j}^{2}+\rho_{i}\mathsf{b}_{i}^{2}+\sigma_{n}^{2}}{1+\chi_{i}}\leq r_{i},$ (63)
-which can be written as a second-order cone (SOC) constraint in (64k). We then re-cast the constraint in (75) as
+which can be written as a second-order cone (SOC) constraint in (52d). We then re-cast the constraint in (62) as
 $\ln\left(1+\chi_{i}\right)-\chi_{i}+\frac{\rho_{i}\mathsf{b}_{i}^{2}}{r_{i}}\geq\frac{Q^{-1}\left(\epsilon_{i}^{\rm(th)}\right)}{\sqrt{L-L_{p}}}+\frac{b_{i}\ln{2}}{L-L_{p}}$ (64)
@@
 From (38), the sensing constraint in (51d) are expressed as
-$\boldsymbol{\rho}^{T}\left(\gamma_{\rm s}\textbf{B}_{D}-M\,\textbf{A}_{D}\,\right)\boldsymbol{\rho}\!\leq\!-\gamma_{\rm s}MN_{\rm rx}\sigma_{n}^{2}.$ (78)
-Finally, we define $\textbf{F}_{k}=\textrm{diag}\left(\sqrt{\mathbb{E}\left\{\|\textbf{w}_{0,k}\|^{2}\right\}},\ldots,\sqrt{\mathbb{E}\left\{\|\textbf{w}_{N_{\rm ue},k}\|^{2}\right\}}\right)$ and rewrite the per-AP power constraints in (63e) in SOC form in terms of $\boldsymbol{\rho}$ as $\left\|\textbf{F}_{k}\boldsymbol{\rho}\right\|\leq\sqrt{P_{\rm tx}},$ , for $k=1,\ldots,N_{\rm tx}$.
+$\bm{\rho}^{T}\left(\gamma_{\rm s}\textbf{B}_{D}-M\,\textbf{A}_{D}\,\right)\bm{\rho}\!\leq\!-\gamma_{\rm s}MN_{\rm rx}\sigma_{n}^{2}.$ (65)
+Finally, we define $\textbf{F}_{k}=\textrm{diag}\left(\sqrt{\mathbb{E}\left\{\|\textbf{w}_{0,k}\|^{2}\right\}},\ldots,\sqrt{\mathbb{E}\left\{\|\textbf{w}_{N_{\rm ue},k}\|^{2}\right\}}\right)$ and rewrite the per-AP power constraints in (51e) in SOC form in terms of $\bm{\rho}$ as $\left\|\textbf{F}_{k}\bm{\rho}\right\|\leq\sqrt{P_{\rm tx}},$ , for $k=1,\ldots,N_{\rm tx}$.

### v3

@@
-Joint Processing and Transmission Energy Optimization for ISAC in Cell-Free Massive MIMO with URLLC
+End-to-End Energy Saving in Cell-Free Massive MIMO ISAC for Ultra-Reliable Target-Aware Actuation
 Abstract
-In this paper, we explore the concept of integrated sensing and communication (ISAC) within a downlink cell-free massive MIMO (multiple-input multiple-output) system featuring multi-static sensing and users requiring ultra-reliable low-latency communications (URLLC).
-Our work focuses on devising two algorithms that concurrently optimize power and blocklength allocation.
-The objectives of these algorithms are to minimize the energy consumption solely for transmission and the overall end-to-end (E2E) energy consumption, which encompasses sensing/communication processing and transmission.
-This is achieved while meeting the necessary requirements for sensing and URLLC.
-To address the inherent non-convexity of these optimization problems, we apply feasible point pursuit - successive convex approximation (FPP-SCA), concave-convex programming (CCP), and fractional programming.
-Our numerical analysis compares the performance of these algorithms in ISAC scenarios and against a baseline URLLC-only scenario without sensing integration.
-Results show the E2E energy minimization algorithm’s effectiveness, particularly in scenarios without sensing.
-Additionally, our study underscores the increasing prominence of energy consumption associated with sensing processing tasks as the number of sensing receive access points rises, indicating potential for energy savings by selectively deactivating APs.
-Furthermore, the results emphasize that although increasing the number of transmit APs can improve URLLC satisfaction and sensing update rate by reducing blocklength, careful management is essential to avoid excessive energy consumption and reduced detection probability.
+[…]
 Index Terms:
@@
 I Introduction
-6G mobile networks are expected to offer various sensing-based applications such as autonomous vehicles, smart homes/cities/factories, remote healthcare, and location/environment-aware scenarios [1, 2].
-Many of these applications, such as autonomous vehicles and industrial Internet-of-things (IIoT), demand ultra-reliable delivering of the sensing information to the user equipments (UEs) with a minimum reliability of 99.999% and within a low end-to-end (E2E) delay of less than 10-150 ms [3, 4, 5].
+6G mobile networks are expected to offer various sensing-based applications such as autonomous vehicles, smart homes/cities/factories, remote healthcare, industrial Internet-of-things (IIoT), and robot control in target/environment-aware scenarios [1, 2].
+These applications introduce ultra-reliable target-aware actuation use cases, which involve scenarios where a communication and sensing system must detect the presence, position, and motion of a target (e.g., a human or object) with high accuracy and minimal delay, and then trigger an actuation response (e.g., controlling a robot, vehicle, or machine) with ultra-reliability and low latency.
+To support such use cases, the system must deliver sensing information to user equipment (UE) with a minimum reliability of 99.999% and an end-to-end (E2E) latency of less than 10–150 ms [3, 4, 5].
 This demand necessitates two key aspects defined in 6G networks: integrated sensing and communication (ISAC) together with ultra-reliable low-latency communication (URLLC).
-For most URLLC applications, short codewords are usually needed to satisfy latency constraints where codes with short blocklengths, e.g., 50-400 symbols are employed.
-Short blocklengths result in decoding errors; therefore, to ensure reliability, it is essential to consider the finite blocklength regime for modeling the decoding error probability (DEP) [6, 7, 8].
-In mission-critical applications such as traffic control and autonomous vehicles, real-time transmission of sensing information to URLLC UEs is essential, as the speed of information updates directly impacts the effectiveness of these applications.
-In this paper, we introduce the concept of the “sensing update rate (SUR),” which quantifies the number of sensing measurements and updates per time unit [9].
-The SUR is influenced by the blocklength, with longer blocklengths resulting in lower update rates due to the extended duration of each sensing operation.
-This dependency creates an opportunity to optimize blocklength to balance reliable communication with high SUR.
+In ultra-reliable target-aware actuation use cases such as traffic control and autonomous vehicles, real-time transmission of sensing information to URLLC UEs is essential, as the timeliness of updates directly affects the performance and safety.
+To capture the frequency at which sensing information must be updated, this paper introduces the concept of the “refreshing rate”, defined by 3GPP [6] as the number of sensing measurements and updates delivered per unit time [7].
+The refreshing rate is closely tied to the communication blocklength: longer blocklengths lead to lower refreshing rates due to the increased time required for each sensing operation.
+This trade-off highlights the potential for optimizing blocklength to achieve a balance between reliable communication and a high refreshing rate.
 Although URLLC and ISAC have typically been studied as separate domains, a joint design and optimization approach is required to meet the distinct requirements of both and maximize integration gains.
-Cell-free massive multiple-input multiple-output (CF-mMIMO) has been shown to be a good candidate to facilitate stringent requirements of URLLC applications [8, 10, 11] as it provides high reliability by handling large path-loss variations and improves cell-edge user performance.
-CF-mMIMO can also facilitate ISAC implementations leveraging distributed access points (APs) to deploy bi-static/ multi-static sensing eliminating the need for full-duplex capability at the APs [12, 13, 14, 15].
-The central processing of network functions within a centralized radio access network (C-RAN) architecture facilitates coordination among distributed APs, enabling joint transmission and processing in CF-mMIMO[16].
-C-RAN has evolved into virtualized C-RAN and, with the open interfaces, further into Open RAN (O-RAN), where virtualization enables the use of general-purpose servers for jointly processing of baseband and sensing signals.
-However, deploying these technologies in 6G is anticipated to result in substantial energy consumption.
-In CF-mMIMO networks, baseband processing combined with sensing tasks is expected to lead to higher energy usage in ISAC networks compared to conventional communication systems.
-The impact of sensing on energy consumption in such networks has not been well-studied.
-Existing studies on green CF-mMIMO primarily focus on optimizing transmit power consumption in radio access, with limited work addressing processing energy, aside from [17].
-Virtualized C-RAN architecture facilitates centralized processing and enables the sharing of processing resources between APs and tasks, paving the way for end-to-end (E2E) optimization of both processing and radio resources in ISAC networks.
-This motivates a comprehensive study of the E2E energy consumption in ISAC networks, considering not only energy consumption for transmission but also for processing tasks.
-The primary research question addressed in this paper is how to incorporate sensing in CF-mMIMO systems with URLLC while minimizing E2E energy consumption.
+Cell-free massive multiple-input multiple-output (CF-mMIMO) has emerged as a strong candidate for meeting the stringent requirements of URLLC applications [8, 9, 10].
+It offers high reliability by compensating for large path-loss variations and significantly improves the performance of cell-edge UEs.
+Moreover, CF-mMIMO is well-suited for implementing ISAC, as its distributed access points (APs) can support bi-static and multi-static sensing configurations [11, 12, 13, 14].
+These configurations eliminate the need for full-duplex capabilities at individual APs but require effective coordination among them.
+Centralized radio access network (C-RAN) architectures, which support centralized processing of network functions, enables joint transmission/reception and processing [15].
+This centralized approach is also essential for facilitating bi-static and multi-static sensing.
+While CF-mMIMO networks offer significant performance advantages, the integration of sensing functionalities is anticipated to substantially increase transmission energy consumption, resulting in up to a tenfold reduction in energy efficiency [16].
+This challenge is further compounded by the fact that the combined demands of baseband processing and sensing operations may result in higher processing energy consumption compared to conventional communication-only systems.
+Most existing research on green CF-mMIMO without sensing integration has primarily focused on optimizing transmit power, with relatively limited attention given to processing energy consumption, aside from a few notable exceptions [17, 18].
+Moreover, the impact of sensing functionalities on processing energy consumption in such networks remains insufficiently explored.
+It is important to mention that communication system efficiency is typically measured in terms of data rate while sensing tasks are evaluated by task-specific metrics such as target detection success rates.
+The execution of a sensing task involves not only signal transmission and reception of reflections from the target but also extensive processing to extract the desired information.
+This post-reception processing is a fundamental component of completing a sensing task, and thus must be explicitly accounted for in system design and analysis.
+This distinction highlights the need for a comprehensive E2E energy consumption analysis that accounts for both transmission and computation costs.
+The virtualization capabilities offered by C-RAN architectures facilitate centralized resource management and flexible allocation of processing resources across APs and tasks.
+This architectural flexibility presents a promising opportunity for E2E optimization of both processing and radio resources in ISAC networks.
+Motivated by these considerations, this paper addresses the following key research question: How can sensing be effectively integrated into CF-mMIMO systems with CRAN architecture and supporting URLLC, while minimizing E2E energy consumption?
 I-A Literature Review
@@
 Network availability is investigated in [22] and [23], and [24] proposes a power allocation algorithm to maximize the minimum signal-to-interference-plus-noise ratio (SINR).
-The authors in [24] propose two power optimization algorithms to provide URLLC for traditional ground UEs and unmanned aerial vehicles (UAVs).
-The objectives in [24] are maximizing either sum or minimum URLLC rate and they study neither energy minimization nor blocklength optimization.
-Max-min rate optimization is also studied in [11], where the authors adapt a special class of the conjugate beamforming for a CF-mMIMO with single-antenna APs.
+The authors in [25] propose two power optimization algorithms to provide URLLC for traditional ground UEs and unmanned aerial vehicles (UAVs) in a CF-mMIMO system.
+The objectives in [25] are maximizing either the sum or the minimum URLLC rate, and they study neither energy minimization nor blocklength optimization.
+Max-min rate optimization is also studied in [10], where the authors adapt a special class of conjugate beamforming for a CF-mMIMO with single-antenna APs.
 Moreover, the authors aim to address an energy efficiency optimization problem.
@@
 Cell-free networks as well as energy minimization are not studied in these works.
-Moreover, sensing requirements are not considered in [23, 21, 24, 11, 25, 26].
+Moreover, sensing requirements are not considered in [24, 22, 25, 10, 26, 27, 23].
 There are few works that jointly consider URLLC and ISAC.
 In [5], a joint precoding scheme is proposed to minimize transmit power, satisfying sensing and delay requirements.
-Moreover, joint ISAC beamforming and scheduling design is addressed in [27] for periodic and aperiodic traffic.
+Moreover, joint ISAC beamforming and scheduling design is addressed in [28] and [29] with a focus on the coexistence of periodic and aperiodic traffic to balance the tradeoff between their corresponding performance.
+The aperiodic traffic is triggered by sensing information about the stochastic environment.
 The consideration of E2E energy-awareness has been explored in various contexts, as reflected in prior works such as [30, 31, 17, 18, 32].
 In particular, [18] studied fully virtualized E2E power minimization problem for CF-mMIMO on O-RAN architecture by taking the radio, fronthaul, and processing resources into account.
-Joint user scheduling and power allocation scheme for cell-free ISAC systems is studied in [32] that aims to maximize the sum-rate of the communication UEs and the sensing targets.
+Joint UE scheduling and power allocation scheme for cell-free ISAC systems is studied in [33], which aims to maximize the sum rate of the communication UEs and the sensing targets.
 However, the authors do not consider either URLLC requirements or energy minimization.
-To the best of our knowledge, for the first time, ISAC in CF-mMIMO with URLLC UEs and with the focus on energy efficiency was studied in our previous work[33].
-In [33], we propose a successive convex approximation-based power allocation algorithm that maximizes energy efficiency while satisfying the sensing and URLLC requirements for a given blocklength.
-However, the impact of blocklength optimization is not considered in [33].
-In this paper, we extend [33] from an E2E energy minimization perspective and aim to address a joint blocklength and power optimization problem.
-Here, E2E refers to considering energy consumption at both radio site and the central cloud.
+TABLE I: NOVELTY COMPARISON OF OUR PAPER TO THE EXISTING WORKS IN LITERATURE
+[19, 20, 21] [22, 23, 24] [25] [10] [26, 27] [5] [28, 29] [30, 31, 17, 18, 32] [16] This paper
+ISAC
+URLLC
+MIMO
+Cell-free networks
+Multi-static sensing
+Target detection
+E2E energy consumption
+Blocklength optimization
+Table I presents a comparison highlighting the novelty of this work relative to existing studies.
+To the best of our knowledge, our previous work [16] was the first to investigate ISAC in CF-mMIMO systems with URLLC UEs, with a specific focus on energy efficiency.
+However, the impact of blocklength optimization was not addressed in that study.
+In this paper, we formulate and solve a joint blocklength and power optimization problem from an E2E energy minimization perspective.
+E2E energy consumption encompasses energy used both at the distributed radio APs and at the centralized cloud infrastructure.
+Figure 1: ISAC network architecture in CF-mMIMO with URLLC UEs and a joint resource allocation mechanism.
 I-B Contributions
+Integration of communication and sensing functionalities in a CF-mMIMO network is depicted in Fig.
+1.
+The objective is to jointly optimize blocklength and power allocation to minimize E2E energy consumption across both the radio access and edge cloud domains.
+URLLC requirements—such as reliability and latency—and sensing requirements—such as refreshing rate and accuracy—are jointly addressed through coordinated resource allocation.
+The edge cloud manages centralized processing for both sensing and communication, while distributed APs enable multi-static sensing alongside communication services.
+This architecture leverages shared radio and computation resources to facilitate joint performance optimization across sensing and communication domains.
 In this paper, we study E2E energy consumption in a CF-mMIMO system with URLLC UEs and multi-static sensing in a cluttered environment.
-We sense one candidate location of the target utilizing communication signals and an additional sensing signal.
-The sensing signals can cause interference for the UEs.
-Thus, we design the sensing precoding vectors to null the interference for the UEs.
-We consider a maximum DEP threshold, representing the reliability requirement, together with a maximum transmission delay threshold as the URLLC requirements and a minimum SINR as the sensing requirement.
-The latter is motivated by the fact that given a fixed false alarm probability, detection probability increases with a higher sensing SINR [34], [9, Chap.
-3 and 15].
-Different from [33], we conduct a comprehensive analysis of the processing requirements, specifically in terms of giga-operations per second (GOPS), and examine its impact on processing energy consumption.
+We sense a candidate location of the potential target utilizing communication signals and an additional sensing signal.
+Since the sensing signals can cause interference for the UEs, we employ null-space precoding vector for the sensing signal to null the interference to the UEs.
 The main contributions of this paper are outlined as follows:
-  • We propose two joint blocklength and power allocation algorithms aimed at minimizing transmission energy consumption and E2E energy consumption, respectively, while satisfying URLLC and sensing requirements.
+  • We conduct a comprehensive analysis of the processing requirements, specifically in terms of giga-operations per second (GOPS) for communication and sensing tasks at the cloud, and provide analytical expression of the cloud energy consumption.
+  • We present an E2E energy minimization algorithm by jointly optimizing the blocklength and power and compare it with two baseline algorithms.
+The first baseline algorithm aims to minimize the E2E energy consumption for only CF-mMIMO systems with URLLC without sensing capability.
+The second baseline considers sensing in the system while it aims to minimize only transmission energy consumption at the radio sites.
 To address the non-convex optimization challenges, we employ feasible point pursuit - successive convex approximation (FPP-SCA), concave-convex programming (CCP), and fractional programming.
-  • We analyze the impact of E2E energy minimization, covering both radio and processing energy consumption, within the proposed ISAC scenario and a comparative scenario that considers only URLLC requirements.
-  • We examine two distinct target detection schemes, comparing their performance in terms of E2E energy consumption, detection probability, and sensing update rate.
-  • We conduct sensitivity analyses to evaluate the effects of DEP threshold, sensing SINR threshold, and the number of ISAC transmit APs and sensing receive APs on E2E energy consumption, detection probability, and sensing update rate.
+  • We evaluate two distinct target detection schemes based on the maximum a posteriori ratio test (MAPRT), each with different processing requirements—namely, the clutter-unaware detector and the clutter-aware detector.
+Their performance is compared in terms of E2E energy consumption and detection probability.
+  • We conduct sensitivity analyses to evaluate the effects of communication and sensing requirements, number of antennas per AP, and sensing receive APs on E2E energy consumption and sensing performance.
 The rest of the paper is organized as follows: Section II introduces the system model.
-Section III provides URLLC analysis, considering DEP and delay in finite blocklength regime.
+Section III provides URLLC analysis, considering decoding error probability (DEP) and delay in the finite blocklength regime.
 Section IV covers the sensing analysis.
@@
 Optimization problems are presented in Section VI, followed by numerical results and conclusions in Sections VII and VIII, respectively.
+[…]
 II System Model
@@
 The system adopts a virtualized C-RAN architecture [34] for uplink channel estimation and downlink communication, as well as multi-static sensing as shown in Fig.
-1.
+2.
 All the APs are interconnected via fronthaul links to the edge cloud and operate in full synchronization.
+Notice that synchronization presents a major challenge in multi-static sensing, as asynchronous transmitters and receivers can lead to ambiguity and performance degradation.
+While this issue has been studied in the context of communication in cell-free networks [35], the synchronization mechanisms already developed for CF-mMIMO communication make it a promising infrastructure for implementing advanced sensing techniques.
+This motivates our focus on exploring other critical aspects of CF-mMIMO-based ISAC systems.
+Another important consideration is the significant fronthaul capacity requirement imposed by centralized signal processing for both communication and sensing functions [36].
+While this paper adopts a centralized processing framework to evaluate the full potential of CF-mMIMO ISAC systems, addressing fronthaul limitations remains an open research direction.
 We consider the original form of CF-mMIMO [8], wherein all the $N_{\rm tx}$ ISAC APs jointly serve the $N_{\rm ue}$ URLLC UEs by transmitting precoded signals containing both communication and sensing symbols.
-Figure 1: Illustration of an ISAC system in CF-mMIMO with URLLC UEs.
+Figure 2: ISAC system model in CF-mMIMO with URLLC UEs.
 Concurrently, the $N_{\rm rx}$ sensing receive APs engage in simultaneous sensing of the candidate location to detect the target.
@@
 Without loss of generality, we assume that we have a channel estimation phase in each transmission.
-II-A Downlink Communication
+II-A Downlink ISAC Transmission
 Let $s_{i}[m]$ and $s_{0}[m]$ represent the downlink communication symbol for UE $i$ and sensing symbol, respectively at time instance $m$.
@@
 Moreover, let $\rho_{i}\geq 0$ and $\rho_{0}\geq 0$ be, respectively, the power control coefficients for UE $i$ and the target, which are fixed throughout the transmission.
-Then, the transmitted signal from transmit AP $k\in\{1,\cdots,N_{\rm tx}\}$ at time instance $m$ is
-$\vspace{-1mm}\textbf{x}_{k}[m]=\sum_{i=0}^{N_{\rm ue}}\sqrt{\rho_{i}}\textbf{w}_{i,k}s_{i}[m]=\textbf{W}_{k}\textbf{D}_{\rm s}[m]\bm{\rho},\vspace{-1mm}$ (1)
+Then, the transmitted signal from transmit AP $k\in\{1,\ldots,N_{\rm tx}\}$ at time instance $m$ is
+$\vskip-2.84526pt\textbf{x}_{k}[m]=\sum_{i=0}^{N_{\rm ue}}\sqrt{\rho_{i}}\textbf{w}_{i,k}s_{i}[m]=\textbf{W}_{k}\textbf{D}_{\rm s}[m]\bm{\rho},\vskip-2.84526pt$ (1)
 where the vectors $\textbf{w}_{i,k}\in\mathbb{C}^{M}$ and $\textbf{w}_{0,k}\in\mathbb{C}^{M}$ are the transmit precoding vectors for transmit AP $k$ corresponding to UE $i$ and the sensing signal, respectively.
 In (1), $\textbf{W}_{k}=\begin{bmatrix}\textbf{w}_{0,k}&\textbf{w}_{1,k}&\cdots&\textbf{w}_{N_{\rm ue},k}\end{bmatrix}$, ${\textbf{D}}_{\rm s}[m]=\textrm{diag}\left(s_{0}[m],s_{1}[m],\ldots,s_{N_{\rm ue}}[m]\right)$ is the diagonal matrix containing the sensing and communication symbols, and $\bm{\rho}=[\sqrt{\rho_{0}}\ \ldots\ \sqrt{\rho_{N_{\rm ue}}}]^{T}$.
-The communication channels are modeled as spatially correlated Rician fading which are assumed to remain constant during each coherence block, and the channel realizations are independent of each other.
+The communication channels are modeled as spatially correlated Rician fading, which are assumed to remain constant during each coherence block, and the channel realizations are independent of each other.
 Let $\textbf{h}_{i,k}\in\mathbb{C}^{M}$ denote the channel between ISAC AP $k$ and UE $i$, modeled as
@@
 Both $\bar{\textbf{h}}_{i,k}$ and $\textbf{R}_{i,k}$ include the combined effect of geometric path loss and shadowing.
-We concatenate the channel vector $\textbf{h}_{i,k}$ in the collective channel vector
+We concatenate the channel vectors $\textbf{h}_{i,k}$ in the collective channel vector
 $\textbf{h}_{i}=\begin{bmatrix}\textbf{h}_{i,1}^{T}&\ldots&\textbf{h}_{i,N_{\rm tx}}^{T}\end{bmatrix}^{T}\in\mathbb{C}^{N_{\rm tx}M},$ (3)
@@
 $+\underbrace{\sqrt{\rho_{0}}\textbf{h}_{i}^{H}\textbf{w}_{0}s_{0}[m]}_{\textrm{Interference signal due to the sensing}}+\underbrace{n_{i}[m]}_{\textrm{Noise}},$ (4)
-where $n_{i}[m]\!\sim\!\mathcal{CN}(0,\sigma_{n}^{2})$ is the independent receiver noise at UE $i$ at time instance $m$ and the collective precoding vectors
-${\textbf{w}_{i}=\left[\textbf{w}_{i,1}^{T}\ \textbf{w}_{i,2}^{T}\ \ldots\ \textbf{w}_{i,N_{\rm tx}}^{T}\right]^{T}\in\mathbb{C}^{N_{\rm tx}M}},$ (5)
-for $i=1,\ldots,N_{\rm ue}$, and
-$\textbf{w}_{0}=\left[\textbf{w}_{0,1}^{T}\ \textbf{w}_{0,2}^{T}\ \ldots\ \textbf{w}_{0,N_{\rm tx}}^{T}\right]^{T}\in\mathbb{C}^{N_{\rm tx}M},$ (6)
-are the centralized precoding vectors.
-Using the independence of the data and sensing signals, the average transmit power for transmit AP $k\in\{1,\ldots,N_{\rm tx}\}$ is
-$P_{k}=\sum_{i=0}^{N_{\rm ue}}\rho_{i}\mathbb{E}\left\{\|\textbf{w}_{i,k}\|^{2}\right\},\quad k=1,\ldots,N_{\rm tx}$ (7)
-which should not exceed the maximum power limit $P_{\rm tx}$.
-In the subsequent section, we will present the specifics of the linear minimum mean-squared error (LMMSE)-based channel estimation.
-II-B LMMSE-Based Channel Estimation
-Within the URLLC transmission time comprising $L$ symbols, $L_{p}<L$ symbols are allocated for pilot transmission.
-A collection of $L_{p}$ mutually orthogonal pilot sequences is represented by $\bm{\phi}_{1},\ldots,\bm{\phi}_{L_{p}}\in\mathbb{C}^{L_{p}}$, where $\|\bm{\phi}_{t}\|^{2}=L_{p}$ for $t=1,\ldots,L_{p}$.
-These pilot sequences are assigned to the $N_{\rm ue}$ UEs.
-If the number of pilot sequences is less than the number of UEs (i.e., $L_{p}<N_{\rm ue}$), then each pilot sequence may be assigned to multiple UEs.
-Let $t_{i}\in\{1,\ldots,L_{p}\}$ denote the index of the pilot sequence that is assigned to UE $i$ and we define $\mathcal{P}_{i}=\{j\in\{1,\ldots,N_{\rm ue}\}|t_{j}=t_{i}\}$ as the set of UEs that share the same pilot with UE $i$.
-The received signal at the ISAC AP $k$ during the entire pilot transmission is denoted by $\textbf{Y}_{k}^{p}\in\mathbb{C}^{M\times L_{p}}$ where
-$\textbf{Y}_{k}^{p}=\sum_{i=1}^{N_{\rm ue}}\sqrt{\eta_{i}}\textbf{h}_{i,k}\bm{\phi}_{t_{i}}^{T}+\textbf{N}_{k},$ (8)
-where $\eta_{i}>0$ is the transmit power by UE $i$ and $\textbf{N}_{k}\in\mathbb{C}^{M\times L_{p}}$ is the received noise at AP $k$ with independent and identically distributed (i.i.d.) $\mathcal{CN}(0,\sigma_{n}^{2})$ entries.
-We assume that $\bar{\textbf{h}}_{i,k}$ and $\textbf{R}_{i,k}$, representing the long-term statistics for the channel in (2), are known while the random phase-shift $\varphi_{i,k}$ is unknown.
-With respect to estimating $\textbf{h}_{i,k}$, let us form $\textbf{y}_{t_{i},k}^{p}=\textbf{Y}_{k}^{p}\bm{\phi}_{t_{i}}^{*}\in\mathbb{C}^{M}$, where
-$\textbf{y}_{t_{i},k}^{p}=\underbrace{\sqrt{\eta_{i}}\,L_{p}\textbf{h}_{i,k}}_{\textrm{Desired part}}+\underbrace{\sum_{j\in\mathcal{P}_{i}\setminus\{i\}}\sqrt{\eta_{j}}\,L_{p}\textbf{h}_{j,k}}_{\textrm{Interference part}}+\underbrace{\textbf{n}_{t_{i},k}}_{\textrm{Noise}},$ (9)
-where $\textbf{n}_{t_{i},k}=\textbf{N}_{k}\bm{\phi}_{t_{i}}^{*}\sim\mathcal{CN}(\textbf{0},\sigma_{n}^{2}L_{p}\textbf{I}_{M})$.
-If the number of UEs is greater than the number of mutually orthogonal pilot sequences, then each pilot sequence may be assigned to multiple UEs using the pilot assignment algorithm in [35, Alg.
-4.1].
-The LMMSE estimate of $\textbf{h}_{i,k}$ is given by [37]
-$\hat{\textbf{h}}_{i,k}=\sqrt{\eta_{i}}\textbf{R}_{i,k}^{\prime}\left(\bm{\Psi}^{\prime}_{t_{i},k}\right)^{-1}\textbf{y}_{t_{i},k}^{p},$ (10)
-where $\textbf{R}_{i,k}^{\prime}$ and $\bm{\Psi}_{t_{i},k}^{\prime}$ are the spatial correlation matrix corresponding to channel $\textbf{h}_{i,k}$ and the correlation matrix of the received signal, respectively, given by
-$\textbf{R}_{i,k}^{\prime}\triangleq\textbf{R}_{i,k}+\bar{\textbf{h}}_{i,k}\bar{\textbf{h}}_{i,k}^{H}$ (11)
-$\bm{\Psi}^{\prime}_{t_{i},k}\triangleq\sum_{j\in\mathcal{P}_{i}}\eta_{j}L_{p}\textbf{R}_{j,k}^{\prime}+\sigma_{n}^{2}\textbf{I}_{M}.$ (12)
-It is worth noting that deriving MMSE channel estimation proves challenging due to the non-Gaussian distribution of channels resulting from random phase-shifts in the LOS paths.
-For the sake of tractability, we adhere to the use of LMMSE-based channel estimation in this paper.
-II-C Multi-Static Sensing
-We investigate multi-static sensing, involving multiple transmit and receive APs in the network.
-The network engages in target location sensing during the downlink phase.
-We assume the existence of a LOS connection between the target location and each transmit/receive AP.
+where $n_{i}[m]\!\sim\!\mathcal{CN}(0,\sigma_{n}^{2})$ is the independent receiver noise at UE $i$ at time instance $m$ and the collective precoding vectors $\textbf{w}_{i}=\left[\textbf{w}_{i,1}^{T}\ \textbf{w}_{i,2}^{T}\ \ldots\ \textbf{w}_{i,N_{\rm tx}}^{T}\right]^{T}\in\mathbb{C}^{N_{\rm tx}M},$ for $i=1,\ldots,N_{\rm ue}$, and $\textbf{w}_{0}=\left[\textbf{w}_{0,1}^{T}\ \textbf{w}_{0,2}^{T}\ \ldots\ \textbf{w}_{0,N_{\rm tx}}^{T}\right]^{T}\in\mathbb{C}^{N_{\rm tx}M}$ are the centralized precoding vectors.
+The communication and sensing transmit precoding vectors are obtained based on regularized zero forcing (RZF) and zero forcing (ZF) approaches, respectively.
+The unit-norm RZF precoding vector for UE $i$ is given as $\textbf{w}_{i}=\frac{\bar{\textbf{w}}_{i}}{\left\|\bar{\textbf{w}}_{i}\right\|}$, with
+$\bar{\textbf{w}}_{i}\!=\!\left(\sum\limits_{j=1}^{N_{\rm ue}}\hat{\textbf{h}}_{j}\hat{\textbf{h}}_{j}^{H}+\delta\textbf{I}_{N_{\rm tx}M}\right)^{\!\!-1}\!\!\hat{\textbf{h}}_{i},\quad i=1,\ldots,N_{\rm ue},$ (5)
+where $\delta$ is the regularization parameter, and $\hat{\textbf{h}}_{j}=\begin{bmatrix}\hat{\textbf{h}}_{j,1}^{T}&\ldots&\hat{\textbf{h}}_{j,N_{\rm tx}}^{T}\end{bmatrix}^{T}\in\mathbb{C}^{N_{\rm tx}M}$ is the linear minimum mean-squared error (LMMSE) channel estimate of the communication channel $\textbf{h}_{j}$, obtained as in [38]111 We omit the explanations from this paper due to the space limitation..
+If the number of UEs is larger than the number of mutually orthogonal pilot sequences, then each pilot sequence may be assigned to multiple UEs using the pilot assignment algorithm in [34, Algorithm 4.1].
+We aim to null the destructive interference from the sensing signal to the UEs by using the unit-norm ZF sensing precoding vector $\textbf{w}_{0}=\frac{\bar{\textbf{w}}_{0}}{\left\|\bar{\textbf{w}}_{0}\right\|}$, where
+$\bar{\textbf{w}}_{0}=\left(\textbf{I}_{N_{\rm tx}M}-\textbf{U}\textbf{U}^{H}\right)\textbf{h}_{0},$ (6)
+and U is the unitary matrix with the orthogonal columns that span the column space of the matrix $\begin{bmatrix}\hat{\textbf{h}}_{1}&\ldots&\hat{\textbf{h}}_{N_{\rm ue}}\end{bmatrix}$.
+$\textbf{h}_{0}=\begin{bmatrix}\sqrt{\beta_{1}}\textbf{a}^{T}(\varphi_{1},\vartheta_{1})&\ldots&\sqrt{\beta_{N_{\rm tx}}}\textbf{a}^{T}(\varphi_{N_{\rm tx}},\vartheta_{N_{\rm tx}})\end{bmatrix}^{T}\in\mathbb{C}^{N_{\rm tx}M}$ is the concatenated sensing channel between all the ISAC APs and the target, including the corresponding channel gains $\beta_{k}$ and the array response vectors $\textbf{a}(\varphi_{k},\vartheta_{k})$ for $k=1,\ldots,N_{\rm tx}$.
+II-B Multi-Static Sensing
+We employ multi-static sensing, utilizing multiple transmit and receive APs within the network.
+Target sensing is performed during the downlink phase.
+A LOS connection is assumed to exist between the target location and each participating transmit and receive AP.
 In the presence of the target, each receive AP captures both the reflected signals from the target and undesired signals, referred to as clutter.
@@
 Consequently, the interference signals correspond to the reflected paths through obstacles and are henceforth denoted as target-free channels.
-Let $\textbf{H}_{r,k}\in\mathbb{C}^{M\times M}$ denote the target-free channel matrix between transmit AP $k$ and receive AP $r$.
-We use the correlated Rayleigh fading model for the NLOS channels $\textbf{H}_{r,k}$, which is modeled using the Kronecker model [38].
-We define the random matrix $\textbf{W}_{{\rm ch},(r,k)}\in\mathbb{C}^{M\times M}$ with i.i.d.
-entries with $\mathcal{CN}(0,1)$ distribution.
+Let $\textbf{H}_{r,k}\in\mathbb{C}^{M\times M}$ denote the target-free channel matrix between transmit AP $k$ and receive AP $r$, which cause interference for sensing.
+We use the correlated Rayleigh fading model for the NLOS channels $\textbf{H}_{r,k}$, which is written as $\textbf{H}_{r,k}=\textbf{R}^{\frac{1}{2}}_{{\rm rx},(r,k)}\textbf{W}_{{\rm ch},{(r,k)}}\left(\textbf{R}^{\frac{1}{2}}_{{\rm tx},(r,k)}\right)^{T}$ using the Kronecker model [39].
+The matrix $\textbf{W}_{{\rm ch},(r,k)}\in\mathbb{C}^{M\times M}$ is a random matrix with independent and identically distributed (i.i.d.) entries with $\mathcal{CN}(0,1)$ distribution.
 The matrix $\textbf{R}_{{\rm rx},(r,k)}\in\mathbb{C}^{M\times M}$ represents the spatial correlation matrix corresponding to receive AP $r$ and with respect to the direction of transmit AP $k$.
 Similarly, $\textbf{R}_{{\rm tx},(r,k)}\in\mathbb{C}^{M\times M}$ is the spatial correlation matrix corresponding to transmit AP $k$ and with respect to the direction of receive AP $r$.
-The channel $\textbf{H}_{r,k}$ is written as
-$\textbf{H}_{r,k}=\textbf{R}^{\frac{1}{2}}_{{\rm rx},(r,k)}\textbf{W}_{{\rm ch},{(r,k)}}\left(\textbf{R}^{\frac{1}{2}}_{{\rm tx},(r,k)}\right)^{T},$ (13)
-where the channel gain is determined by the geometric path loss and shadowing, and is included in the spatial correlation matrices.
-The received signal at AP $r$ in the presence of the target and for $m\in\{1,\cdots,L_{d}\}$, can be formulated as
+The channel gain is determined by the geometric path loss and shadowing, and is included in the spatial correlation matrices.
+The received signal at AP $r$ in the presence of the target and for $m\in\{1,\ldots,L_{d}\}$, can be formulated as
 $\textbf{y}_{r}[m]=\sum_{k=1}^{N_{\rm tx}}\underbrace{\alpha_{r,k}\sqrt{\beta_{r,k}}\textbf{a}(\phi_{r},\theta_{r})\textbf{a}^{T}(\varphi_{k},\vartheta_{k})\textbf{x}_{k}[m]}_{\textrm{desired reflections from the target}}$
-$+\sum_{k=1}^{N_{\rm tx}}\underbrace{\textbf{H}_{r,k}\textbf{x}_{k}[m]}_{\textrm{undesired signals}}+\textbf{n}_{r}[m],$ (14)
+$+\sum_{k=1}^{N_{\rm tx}}\underbrace{\textbf{H}_{r,k}\textbf{x}_{k}[m]}_{\textrm{clutter}}+\textbf{n}_{r}[m],$ (7)
+where $\textbf{n}_{r}[m]\sim\mathcal{CN}(\textbf{0},\sigma_{n}^{2}\textbf{I}_{M})$ is the receiver noise at the $M$ antennas of receive AP $r$.
+The second term in (II-B) refers to clutter which acts as interference for the target detection.
+Here, $\beta_{r,k}$ is the channel gain, including the path loss from transmit AP $k$ to receive AP $r$ through the target and the variance of bi-static radar cross-section (RCS) of the target denoted by $\sigma_{\rm rcs}$.
+$\beta_{r,k}$ is computed according to the radar range equation for bi-static sensing in [7, Chap.
+2], given by
+$\beta_{r,k}=\frac{\lambda_{c}^{2}\,\sigma_{\rm rcs}}{(4\pi)^{3}d_{{\rm tx},k}^{2}d_{{\rm rx},r}^{2}}.$ (8)
 […]
-Following the same notation as in [34], in (II-C), the known part of each reflected path is denoted by ${\textbf{g}_{r,k}[m]\in\mathbb{C}^{M}}$, defined as
+Following the same notation as in [40], in (II-B), the known part of each reflected path is denoted by ${\textbf{g}_{r,k}[m]\in\mathbb{C}^{M}}$, defined as
 $\textbf{g}_{r,k}[m]\triangleq\sqrt{\beta_{r,k}}\textbf{a}(\phi_{r},\theta_{r})\textbf{a}^{T}(\varphi_{k},\vartheta_{k})\textbf{x}_{k}[m],$ (9)
@@
 Similarly, $\varphi_{k}$ and $\vartheta_{k}$ represent the azimuth and elevation angles from transmit AP $k$ to the target location.
-Each receive AP sends their respective signals $\textbf{y}_{r}[m]$, for $r=1,\ldots,N_{\rm rx}$, to the edge cloud, to form the collective received signal $\textbf{y}[m]=\begin{bmatrix}\textbf{y}_{1}^{T}[m]&\cdots&\textbf{y}_{N_{\rm rx}}^{T}[m]\end{bmatrix}^{T}$, which can be expressed as
+Each receive AP sends their respective signals $\textbf{y}_{r}[m]$, for $r=1,\ldots,N_{\rm rx}$, to the edge cloud to form the collective received signal $\textbf{y}[m]=\begin{bmatrix}\textbf{y}_{1}^{T}[m]&\ldots&\textbf{y}_{N_{\rm rx}}^{T}[m]\end{bmatrix}^{T}$, which can be expressed as
 $\textbf{y}[m]$ $=\underbrace{\mathrm{blkdiag}\left(\textbf{G}_{1}[m],\ldots,\textbf{G}_{N_{\rm rx}}[m]\right)}_{\triangleq\textbf{G}[m]}\bm{\alpha}$
-$+\underbrace{\left(\textbf{I}_{N_{\rm rx}}\otimes\left(\textbf{x}^{T}[m]\otimes\textbf{I}_{M}\right)\right)}_{\triangleq\textbf{X}[m]}\bm{\mathfrak{h}}+\textbf{n}[m],$ (16)
-where $\textbf{G}_{r}[m]=\begin{bmatrix}\textbf{g}_{r,1}[m]&\cdots&\textbf{g}_{r,N_{\rm tx}}[m]\end{bmatrix}\in\mathbb{C}^{M\times N_{\rm tx}}$, $\textbf{x}[m]=\begin{bmatrix}\textbf{x}^{T}_{1}[m]&\cdots&\textbf{x}^{T}_{N_{\rm tx}}[m]\end{bmatrix}^{T}\in\mathbb{C}^{N_{\rm tx}M}$, and $\mathfrak{h}\sim\mathcal{CN}\left(\textbf{0},\textbf{R}\right)$ is the vectorized target-free channels[34, Sec.
+$+\underbrace{\left(\textbf{I}_{N_{\rm rx}}\otimes\left(\textbf{x}^{T}[m]\otimes\textbf{I}_{M}\right)\right)}_{\triangleq\textbf{X}[m]}\bm{\mathfrak{h}}+\textbf{n}[m]$
+$=\textbf{G}[m]\bm{\alpha}+\textbf{X}[m]\bm{\mathfrak{h}}+\textbf{n}[m],$ (10)
+where $\textbf{G}_{r}[m]=\begin{bmatrix}\textbf{g}_{r,1}[m]&\ldots&\textbf{g}_{r,N_{\rm tx}}[m]\end{bmatrix}\in\mathbb{C}^{M\times N_{\rm tx}}$, $\textbf{x}[m]=\begin{bmatrix}\textbf{x}^{T}_{1}[m]&\ldots&\textbf{x}^{T}_{N_{\rm tx}}[m]\end{bmatrix}^{T}\in\mathbb{C}^{N_{\rm tx}M}$, and $\bm{\mathfrak{h}}\sim\mathcal{CN}\left(\textbf{0},\textbf{R}\right)$ is the vectorized target-free channel[40, Sec.
 V].
-II-D Target Detection
-We employ two maximum a posteriori ratio test (MAPRT) detectors with two levels of complexity: the clutter-unaware detector [13] and the clutter-aware detector [34, Lem.
-2].
+In addition, we have defined $\textbf{G}[m]=\mathrm{blkdiag}\left(\textbf{G}_{1}[m],\ldots,\textbf{G}_{N_{\rm rx}}[m]\right)$ and $\textbf{X}[m]=\left(\textbf{I}_{N_{\rm rx}}\otimes\left(\textbf{x}^{T}[m]\otimes\textbf{I}_{M}\right)\right)$.
+We employ two maximum a posteriori ratio test (MAPRT) detectors with two levels of complexity: the clutter-unaware detector and the clutter-aware detector, proposed in [12] and [40, Lem.
+2], respectively.
 The clutter-unaware detector ignores the presence of clutter for the sake of complexity reduction, whereas the clutter-aware detector accounts for unknown clutter.
@@
 $\textbf{E}=\sum_{m=1}^{L_{d}}\textbf{G}^{H}[m]\textbf{X}[m].$ (16)
-II-E Transmit ISAC Precoding Vectors
-The communication and sensing transmit precoding vectors are obtained based on regularized zero forcing (RZF) and zero forcing (ZF) approaches, respectively.
-The unit-norm RZF precoding vector for UE $i$ is given as $\textbf{w}_{i}=\frac{\bar{\textbf{w}}_{i}}{\left\|\bar{\textbf{w}}_{i}\right\|}$, with
-$\bar{\textbf{w}}_{i}\!=\!\left(\sum\limits_{j=1}^{N_{\rm ue}}\hat{\textbf{h}}_{j}\hat{\textbf{h}}_{j}^{H}+\delta\textbf{I}_{N_{\rm tx}M}\right)^{\!\!-1}\!\!\hat{\textbf{h}}_{i},\quad i=1,\ldots,N_{\rm ue},$ (23)
-where $\delta$ is the regularization parameter, and $\hat{\textbf{h}}_{j}=\begin{bmatrix}\hat{\textbf{h}}_{j,1}^{T}&\ldots&\hat{\textbf{h}}_{j,N_{\rm tx}}^{T}\end{bmatrix}^{T}\in\mathbb{C}^{N_{\rm tx}M}$ is the LMMSE channel estimate of the communication channel $\textbf{h}_{j}$.
-We aim to null the destructive interference from the sensing signal to the UEs by using the unit-norm ZF sensing precoding vector $\textbf{w}_{0}=\frac{\bar{\textbf{w}}_{0}}{\left\|\bar{\textbf{w}}_{0}\right\|}$, where
-$\bar{\textbf{w}}_{0}=\left(\textbf{I}_{N_{\rm tx}M}-\textbf{U}\textbf{U}^{H}\right)\textbf{h}_{0},$ (24)
-and U is the unitary matrix with the orthogonal columns that span the column space of the matrix $\begin{bmatrix}\hat{\textbf{h}}_{1}&\ldots&\hat{\textbf{h}}_{N_{\rm ue}}\end{bmatrix}$.
-$\textbf{h}_{0}=\begin{bmatrix}\sqrt{\beta_{1}}\textbf{a}^{T}(\varphi_{1},\vartheta_{1})&\ldots&\sqrt{\beta_{N_{\rm tx}}}\textbf{a}^{T}(\varphi_{N_{\rm tx}},\vartheta_{N_{\rm tx}})\end{bmatrix}^{T}\in\mathbb{C}^{N_{\rm tx}M}$ is the concatenated sensing channel between all the ISAC APs and the target, including corresponding path loss and the array response vectors $\textbf{a}(\varphi_{k},\vartheta_{k})$ for $k=1,\ldots,N_{\rm tx}$.
 III Reliability and Delay Analysis for URLLC
+For most URLLC applications, short codewords are usually needed to satisfy latency constraints where codes with short blocklengths, e.g., 50-400 symbols are employed.
+Short blocklengths result in decoding errors; therefore, to ensure reliability, it is essential to consider the finite blocklength regime for modeling the decoding error probability (DEP) [41, 42, 8].
+We consider a maximum DEP threshold, representing the reliability requirement, together with a maximum transmission delay threshold as the URLLC requirements.
 In this section, we derive an upper bound on the DEP and the transmission delay, both crucial aspects considered as URLLC requirements.
-In the finite blocklength regime, the communication data cannot be transmitted without error.
-From [10], ergodic data rate of UE $i$ can be approximated as
-$R_{i}\!\approx\!\mathbb{E}\Bigg{\{}\!\!\left(1\!-\!\beta\right)\!\log_{2}{\left(1+\mathsf{SINR}^{\rm(dl)}_{i}\right)}\!-\!\frac{Q^{-1}(\epsilon_{i})}{\ln(2)}\!\sqrt{\!\frac{(1\!-\!\beta)V_{i}}{L}}\Bigg{\}},\vspace{-0mm}$ (25)
-where $\beta=\frac{L_{p}}{L}$, $\epsilon_{i}$ denotes the DEP when transmitting $b_{i}$ bits to UE $i$, $\mathsf{SINR}^{\rm(dl)}_{i}$ is the instantaneous downlink communication signal-to-interference-plus-noise ratio (SINR) for UE $i$, $V_{i}=1-\left(1+\mathsf{SINR}^{\rm(dl)}_{i}\right)^{-2}$ is the channel dispersion, and $Q(\cdot)$ refers to the Gaussian Q-function.
+In the finite blocklength regime, the communication data cannot be transmitted without error. From [9], ergodic data rate of UE $i$ can be approximated as
+$R_{i}\!\approx\!\mathbb{E}\Bigg{\{}\!\!\left(1\!-\!\beta\right)\!\log_{2}{\left(1+\mathsf{SINR}^{\rm(dl)}_{i}\right)}\!-\!\frac{Q^{-1}(\epsilon_{i})}{\ln(2)}\!\sqrt{\!\frac{(1\!-\!\beta)V_{i}}{L}}\Bigg{\}},\vskip 0.0pt$ (17)
+where $\beta=\frac{L_{p}}{L}$, $\epsilon_{i}$ denotes the DEP when transmitting $b_{i}$ bits to UE $i$, $\mathsf{SINR}^{\rm(dl)}_{i}$ is the instantaneous downlink communication SINR for UE $i$, $V_{i}=1-\left(1+\mathsf{SINR}^{\rm(dl)}_{i}\right)^{-2}$ is the channel dispersion, and $Q(\cdot)$ refers to the Gaussian Q-function.
 Due to the fact that $V_{i}<1$, the ergodic data rate can be lower bounded by
 $R_{i}$ $\!\geq\!\left(1\!-\!\beta\right)\!\mathbb{E}\Bigg{\{}\!\!\log_{2}{\left(1+\mathsf{SINR}^{(\rm dl)}_{i}\right)}\!\!\Bigg{\}}\!-\!\frac{Q^{-1}(\epsilon_{i})}{\ln(2)}\sqrt{\frac{(1-\beta)}{L}}.$ (18)
-Moreover, given that only $\mathbb{E}\left\{\textbf{h}_{i}^{H}\textbf{w}_{i}\right\}$ is known at UE $i$, and according to [35, Thm. 6.1] and [34, Lem. 1],
+Moreover, given that only $\mathbb{E}\left\{\textbf{h}_{i}^{H}\textbf{w}_{i}\right\}$ is known at UE $i$, and according to [34, Thm. 6.1] and [40, Lem. 1],
 $\mathbb{E}\left\{\log_{2}{\left(1+\mathsf{SINR}^{\rm(dl)}_{i}\right)}\right\}\geq\log_{2}{\left(1+\overline{\mathsf{SINR}}^{(\rm dl)}_{i}\right)}$ (19)
@@
 This implies that the blocklength cannot exceed $D^{\rm(th)}_{i}B(1-\epsilon_{i}^{\rm(th)})$.
-Thus, we can define the maximum tolerable blocklength by $L_{\rm max}$, where
-$L_{\rm max}=\min\left\{D^{\rm(th)}_{i}B\left(1-\epsilon_{i}^{\rm(th)}\right)|i=1,\cdots,N_{\rm ue}\right\}.$ (34)
+Thus, we can define the maximum tolerable blocklength by $L_{\rm max,c}$, where
+$L_{\rm max,c}=\min\left\{D^{\rm(th)}_{i}B\left(1-\epsilon_{i}^{\rm(th)}\right)|i=1,\ldots,N_{\rm ue}\right\}.$ (26)
 IV Sensing Analysis
-We assess the sensing performance in terms of detection probability under a certain false alarm probability and sensing update rate in terms of the number of sensing measurement updates per time unit.
-It is numerically shown in [13] and [34], that maximizing the sensing SINR improves the target detection probability under a constant false alarm probability.
-When studying other sensing tasks, it is naturally desired to keep the sensing SINR above a required threshold denoted by $\gamma_{\rm s}$, i.e., $\mathsf{SINR}_{s}\geq\gamma_{\rm s}$.
-This fact motivates the optimization problems presented in Section VI.
-In the following subsections, we present sensing SINR and sensing update rate.
-IV-A Sensing SINR
-From [34] , the sensing SINR is given by
-$\mathsf{SINR_{\rm s}}=\frac{\bm{\rho}^{T}\textbf{A}\bm{\rho}}{L_{d}MN_{\rm rx}\sigma_{n}^{2}+\bm{\rho}^{T}\textbf{B}\bm{\rho}},$ (35)
-where $\bm{\rho}=[\sqrt{\rho_{0}}\ \ldots\ \sqrt{\rho_{N_{\rm ue}}}]^{T}$ and
-[…]
-×$\hskip 8.53581pt\times\textbf{W}_{k}\Bigg{)}\textbf{D}_{\rm s}[m],$ (36)
-B $=\sum_{m=1}^{L_{d}}\textbf{D}_{\rm s}^{H}[m]\left(\sum_{r=1}^{N_{\rm rx}}\sum_{k=1}^{N_{\rm tx}}\mathrm{tr}\left(\textbf{R}_{\mathrm{rx},(r,k)}\right)\textbf{W}_{k}^{H}\textbf{R}^{T}_{{\rm tx},(r,k)}\textbf{W}_{k}\right)$
-×$\hskip 8.53581pt\times\textbf{D}_{\rm s}[m].$ (37)
+We assess the sensing performance in terms of detection probability under a certain false alarm probability and refreshing rate in terms of the number of sensing measurement updates per time unit.
+Detection probability refers to the likelihood of correctly identifying the presence of a target, whereas false alarm probability denotes the likelihood of incorrectly detecting a target when none is present.
+Maximizing the sensing SINR enhances the probability of target detection under a fixed false alarm probability [7, Chapters 3 and 15].
+For other sensing tasks—such as estimating range, velocity, and similar parameters— it is also desired to maintain the sensing SINR above a required threshold denoted by $\gamma_{\rm s}$, i.e., $\mathsf{SINR}_{\rm s}\geq\gamma_{\rm s}$.This requirement serves as a key motivation for the optimization problem formulated in Section  VI.
+In this work, we consider a minimum sensing SINR and refreshing rate as the sensing requirements.
 The sensing SINR is a function of symbols, which vary randomly for different blocklength values.
@@
 Thus, the average sensing SINR would be
-$\overline{\mathsf{SINR}}_{s}$ $=\frac{M\bm{\rho}^{T}\textbf{A}_{D}\bm{\rho}}{MN_{\rm rx}\sigma_{n}^{2}+\bm{\rho}^{T}\textbf{B}_{D}\bm{\rho}},$ (38)
+$\overline{\mathsf{SINR}}_{\rm s}$ $=\frac{M\bm{\rho}^{T}\textbf{A}_{D}\bm{\rho}}{MN_{\rm rx}\sigma_{n}^{2}+\bm{\rho}^{T}\textbf{B}_{D}\bm{\rho}},$ (27)
 where $\textbf{A}_{D}$ and $\textbf{B}_{D}$ are diagonal matrices with
@@
 $\left[\textbf{B}_{D}\right]_{ii}$ $=\sum_{r=1}^{N_{\rm rx}}\sum_{k=1}^{N_{\rm tx}}\mathrm{tr}\left(\textbf{R}_{\mathrm{rx},(r,k)}\right)\left(\textbf{W}_{k}^{H}\textbf{R}^{T}_{{\rm tx},(r,k)}\textbf{W}_{k}\right)_{ii}.$ (29)
-IV-B Sensing Update Rate (SUR)
-In our system model, the sensing information is updated once throughout each transmission block. Thus, the sensing update rate, denoted as $R_{\rm s}$, is
+According to the 3GPP specification [6], refreshing rate is defined as the rate at which the sensing result is generated by the sensing system.
+It is the inverse of the time elapsed between two successive sensing results.
+In our system model, the sensing information is updated once throughout each transmission block.
+Thus, the sensing service latency is $\frac{L}{B}$ and the refreshing rate, denoted as $R_{\rm s}$, is
 $R_{\rm s}=\frac{B}{L}\quad(\rm updates/second).$ (30)
@@
 The total power consumption, taking into account both communication and sensing, can be expressed as
-$P_{\rm total}$ $=\underbrace{\sum_{k=1}^{N_{\rm tx}}\left(P_{\textrm{AP},0}^{\rm tx}+\Delta^{\rm tr}P_{k}\right)+\sum_{r=1}^{N_{\rm rx}}P_{\textrm{AP},0}^{\rm rx}}_{\textrm{radio site power consumption}}+P_{\rm cloud}$ (42)
+$P_{\rm total}$ $=\underbrace{\Delta^{\rm tr}P_{\rm tr}+\sum_{k=1}^{N_{\rm tx}}P_{\textrm{AP},0}^{\rm tx}+\sum_{r=1}^{N_{\rm rx}}P_{\textrm{AP},0}^{\rm rx}}_{\textrm{radio site power consumption}}+P_{\rm cloud}$ (31)
 where $\Delta^{\rm tr}$ is the slope of load-dependent transmission power consumption of each AP.
-The $P_{k}$ is the transmission power of the ISAC AP $k$ from (7).
+The average power consumption for transmission is given as $P_{\rm tr}=\sum_{j=0}^{N_{\rm ue}}\rho_{j}=\|\bm{\rho}\|^{2}$, where $\bm{\rho}=[\sqrt{\rho_{0}}\ \ldots\ \sqrt{\rho_{N_{\rm ue}}}]^{T}$ due to the unit-power centralized precoding vectors.
 The power consumption at the cloud is modeled as
-$P_{\rm cloud}$ $=P_{\rm fixed}+\frac{1}{\sigma_{\rm cool}}\Bigg{(}P_{\rm cloud,0}^{\rm proc}+\Delta_{\rm cloud}^{\rm proc}\frac{C_{\rm cloud}}{C_{\rm max}}\Bigg{)},$ (43)
-where $P_{\rm fixed}$ is the fixed power consumption at the cloud which is independent of the load.
-$P_{\rm cloud,0}^{\rm proc}$ is the processing power consumption in the idle mode.
+$P_{\rm cloud}$ $=P_{\rm fixed}+\frac{1}{\sigma_{\rm cool}}\Bigg{(}N_{\rm GPP}P_{\rm cloud,0}^{\rm proc}+\Delta_{\rm cloud}^{\rm proc}\frac{C_{\rm cloud}}{C_{\rm max}}\Bigg{)},$ (32)
+where $P_{\rm fixed}$ is the fixed power consumption at the cloud, which is independent of the load.
+$N_{\rm GPP}=\lceil\frac{C_{\rm cloud}}{C_{\rm max}}\rceil$ is the number of required general-purpose processors (GPPs), and $P_{\rm cloud,0}^{\rm proc}$ is the processing power consumption in the idle mode.
 $\sigma_{\rm cool}\in(0,1]$ and $\Delta_{\rm cloud}^{\rm proc}$ denote the cooling efficiency of the cloud and the slope of the load-dependent power consumption for processing at the digital unit (DU) in the cloud, respectively.
-Moreover, $C_{\rm max}$ and $C_{\rm cloud}\in[0,C_{\rm max}]$ are the maximum processing capacity of the processing resources and the total processing resource utilization in GOPS, respectively[30, 39].
+Moreover, $C_{\rm max}$ and $C_{\rm cloud}$ are the maximum processing capacity of the processing resources in each GPP and the total processing resource utilization in GOPS, respectively [18, 43].
 The processing resource utilization can be expressed as
@@
 We also consider memory overhead in arithmetic operation calculations by multiplying each operation by two as done in [44, 18].
-Hence, each complex multiplication is counted as $4\cdot 2=8$ operations in computing the total GOPS.
+Hence, each complex multiplication is counted as 42=84$\cdot$ 2=84 ⋅ 2 = 8 operations in computing the total GOPS.
 V-A GOPS Analysis of Digital Operations for Communication at the Cloud
 In this section, we analyze the GOPS for digital signal processing corresponding to the communication tasks including the uplink channel estimation and downlink transmission.
-To compute the number of real multiplications, we mainly follow the GOPS analysis in [36, App.
+To compute the number of real multiplications, we mainly follow the GOPS analysis in [37, App.
 B], [17].
@@
 To compute the channel estimates, we first obtain the vectors $\textbf{y}_{t_{i},k}^{p}$.
-From [36, App.
+From [37, App.
 B], the multiplication of one matrix of size $M\times L_{p}$ with a vector of size $L_{p}\times 1$ results in $ML_{p}$ complex multiplications.
@@
 However, this pre-computation can be neglected since the channel statistics are usually constant for a while and there is no need to compute them every coherence block.
-Finally, the multiplications in (10) costs $8M^{2}N_{\rm ue}N_{\rm tx}$ real operations.
+We use the LMMSE channel estimation approach in [38] which costs $8M^{2}N_{\rm ue}N_{\rm tx}$ real operations.
 Then, $C_{\textrm{ch-est}}$ is equal to
 $C_{\textrm{ch-est}}=\left\{\begin{matrix}(8ML_{p}+8M^{2})N_{\rm ue}N_{\rm tx},&L_{p}\geq N_{\rm ue}\\ 8ML_{p}^{2}N_{\rm tx}+8M^{2}N_{\rm ue}N_{\rm tx},&L_{p}<N_{\rm ue}.&\end{matrix}\right.$ (34)
-The number of real multiplications/divisions to compute centralized RZF precoding vector for all the UEs from [36, App. B] is
+The number of real multiplications/divisions to compute centralized RZF precoding vector for all the UEs from [37, App. B] is
 $C_{\textrm{prec-comp}}$ $=\underbrace{\frac{8\left((MN_{\rm tx})^{2}+MN_{\rm tx}\right)}{2}N_{\rm ue}}_{\textrm{inside the parentheses}}$
@@
 $\quad+N_{\rm ue}\cdot\left(\underbrace{8MN_{\rm tx}}_{\textrm{computing }||\textbf{w}_{i}||}+\underbrace{4MN_{\rm tx}}_{\textrm{normalization}}\right)$
-$=\left(12(MN_{\rm tx})^{2}+16MN_{\rm tx}\right)N_{\rm ue}$
-$\quad+\frac{8\left((MN_{\rm tx})^{3}-MN_{\rm tx}\right)}{3}.$ (46)
+$=\left(12(MN_{\rm tx})^{2}+16MN_{\rm tx}\right)N_{\rm ue}+\frac{8\left((MN_{\rm tx})^{3}-MN_{\rm tx}\right)}{3}.$ (35)
 Reciprocity calibration and multiplication of the symbols by the precoding vectors, each costs $8L_{d}MN_{\rm ue}N_{\rm tx}$ real operations [45, 18].
 Multiplying by the power coefficients also costs $4L_{d}MN_{\rm ue}N_{\rm tx}$.
-Finally, the GOPS corresponding to communication processing (i.e., channel estimation, precoding and reciprocity calibration) is computed as
-$C_{\textrm{proc}}^{\rm c}=\frac{B}{L10^{9}}\Bigg{(}C_{\textrm{ch-est}}+C_{\textrm{prec-comp}}+20L_{d}MN_{\rm ue}N_{\rm tx}\Bigg{)},$ (47)
-where we divided the total giga operations by the coherence time $L/B$.
+Finally, the GOPS corresponding to communication processing (i.e., channel estimation, precoding and reciprocity calibration) is computed as in (36) where we divided the total giga operations by the coherence time $L/B$.
+$C_{\textrm{proc}}^{\rm c}=\frac{B}{L10^{9}}\Bigg{(}C_{\textrm{ch-est}}+C_{\textrm{prec-comp}}+20L_{d}MN_{\rm ue}N_{\rm tx}\Bigg{)}.$ (36)
+TABLE II: COMPUTATIONAL COMPLEXITY FOR CLUTTER-AWARE DETECTOR
+Operation Computational complexity
+$\textbf{G}[m],\quad\forall m$ $20\,L_{d}\,MN_{\rm tx}N_{\rm rx}$
+a $8\,L_{d}\,M^{2}\,N_{\rm rx}\,N_{\rm tx}\,$
+b $8\,L_{d}\,M^{2}\,N_{\rm rx}\,N_{\rm tx}$
+C $4\,L_{d}MN_{\rm rx}\left(N_{\rm tx}^{2}+N_{\rm tx}\right)$
+D $4L_{d}\,\left((MN_{\rm tx})^{2}+MN_{\rm tx}\right)$
+E $8\,L_{d}\,M^{2}\,N_{\rm rx}\,N_{\rm tx}^{2}$
+$\begin{bmatrix}\textbf{C}&\textbf{E}\\ \textbf{E}^{H}&\textbf{D}\end{bmatrix}^{-1}$
+$\frac{8}{3}\left(\left((1+M^{2})N_{\rm tx}N_{\rm rx}\right)^{3}-(1+M^{2})N_{\rm tx}N_{\rm rx}\right)$
+$\textbf{D}^{-1}$ $\frac{8}{3}\left((M^{2}N_{\rm tx}N_{\rm rx})^{3}-M^{2}N_{\rm tx}N_{\rm rx}\right)$
+$\begin{bmatrix}\textbf{a}^{H}\!&\textbf{b}^{H}\end{bmatrix}\!\begin{bmatrix}(.)&(.)\\ (.)&(.)\end{bmatrix}\begin{bmatrix}\textbf{a}\\ \textbf{b}\end{bmatrix}$
+$8\left((1+M^{2})N_{\rm tx}N_{\rm rx}\right)^{2}+8(1+M^{2})N_{\rm tx}N_{\rm rx}$
 V-B GOPS Analysis of Digital Operations for Sensing at the Cloud
@@
 Multiplying the sensing symbols by the sensing precoding vector and the sensing power coefficient costs $12L_{d}MN_{\rm tx}$ real multiplications/divisions.
-The unitary matrix U in (24) (subspace spanned by the UE channel estimation vectors) is already obtained when computing the RZF precoding vectors by matrix inversion and the corresponding $\mathbf{L}\mathbf{D}\mathbf{L}^{H}$ decomposition [36, Lem.
+The unitary matrix U in (6) (subspace spanned by the UE channel estimation vectors) is already obtained when computing the RZF precoding vectors by matrix inversion and the corresponding $\mathbf{L}\mathbf{D}\mathbf{L}^{H}$ decomposition [37, Lem.
 B.2].
@@
 At the cloud, these signals are processed for a specific sensing application.
-In this work we consider target detection and assess the number of real multiplications/divisions required to compute the test statistics.
+In this work, we consider target detection and assess the number of real multiplications/divisions required to compute the test statistics.
 For target detection problems, we usually compute the test statistics and compare them with a threshold.
@@
 $C_{\rm proc}^{\rm s}$ $=\frac{B}{L10^{9}}\Bigg{(}\underbrace{12L_{d}MN_{\rm tx}}_{\textrm{precoding + power}}+C_{\textrm{prec-comp}}^{\rm s}+\underbrace{C_{\rm se-comp}}_{\textrm{target detection}}\Bigg{)},$ (38)
-where $C_{\rm se-comp}$ denotes the computational complexity for computing the detector test statistics. Given the clutter-unaware detector in (17), the computational complexity is computed as
+where $C_{\rm se-comp}$ denotes the computational complexity for computing the detector test statistics.
+Given the clutter-unaware detector in (11), the computational complexity is computed as
 $C_{\rm se-comp}^{\rm c-unaware}$ $=L_{d}\Bigg{(}\underbrace{20MN_{\rm tx}N_{\rm rx}}_{\rm{for}\,\textbf{ G}}+\underbrace{8\,M^{2}\,N_{\rm rx}\,N_{\rm tx}}_{\rm{for}\,\textbf{ a}}$
@@
 where we used the block diagonal structure of C in getting the inverse of it.
-The computational complexity for each step of computing the test statistics of clutter-aware detector in (18) are listed in Table I.
-TABLE I: Computational complexity with clutter-aware MAPRT detector
-Operation Computational complexity
-$\textbf{G}[m],\quad\forall m$ $\mathcal{O}\left(20\,L_{d}\,MN_{\rm tx}N_{\rm rx}\right)$
-a $\mathcal{O}\left(8\,L_{d}\,M^{2}\,N_{\rm rx}\,N_{\rm tx}\right)\,$
-b $\mathcal{O}\left(8\,L_{d}\,M^{2}\,N_{\rm rx}\,N_{\rm tx}\,\right)$
-C $\mathcal{O}\left(4\,L_{d}MN_{\rm rx}\left(N_{\rm tx}^{2}+N_{\rm tx}\right)\right)$
-D $\mathcal{O}\left(4L_{d}\,\left((MN_{\rm tx})^{2}+MN_{\rm tx}\right)\right)$
-E $\mathcal{O}\left(8\,L_{d}\,M^{2}\,N_{\rm rx}\,N_{\rm tx}^{2}\right)$
-$\begin{bmatrix}\textbf{C}&\textbf{E}\\ \textbf{E}^{H}&\textbf{D}\end{bmatrix}^{-1}$ $\mathcal{O}\left(\frac{8}{3}\left(\left((1+M^{2})N_{\rm tx}N_{\rm rx}\right)^{3}-(1+M^{2})N_{\rm tx}N_{\rm rx}\right)\right)$
-$\textbf{D}^{-1}$ $\mathcal{O}\left(\frac{8}{3}\left((M^{2}N_{\rm tx}N_{\rm rx})^{3}-M^{2}N_{\rm tx}N_{\rm rx}\right)\right)$
-$\begin{bmatrix}\textbf{a}^{H}&\textbf{b}^{H}\end{bmatrix}\begin{bmatrix}(.)&(.)\\ (.)&(.)\end{bmatrix}\begin{bmatrix}\textbf{a}\\ \textbf{b}\end{bmatrix}$ $\mathcal{O}\left(8\,\left(\left((1+M^{2})N_{\rm tx}N_{\rm rx}\right)^{2}+(1+M^{2})N_{\rm tx}N_{\rm rx}\right)\right)$
+The computational complexity for each step of computing the test statistics of clutter-aware detector in (12) are listed in Table II.
 VI Joint blocklength and Power Optimization
-In this paper, we aim to optimize the blocklength and the power control coefficients to minimize the energy consumption while URLLC and sensing requirements are satisfied.
-In the following subsections, we cast two non-convex optimization problems: i) minimizing only the energy consumption for transmission; ii) minimizing E2E energy consumption by taking into account the energy consumption at the radio site and the cloud.
+We aim to jointly optimize the blocklength and the power control coefficients to minimize the E2E energy consumption while URLLC and sensing requirements are satisfied.
 [^3] In addition to considering radio and cloud processing energy consumption, it is important to take fronthaul energy consumption into account when assessing E2E energy consumption.
 However, in our analysis, where the primary focus is on the interplay between processing and radio resources, we treat fronthaul energy consumption as a fixed component and do not include it in our considerations.
-The corresponding algorithms to these optimization problems are represented by SeURLLC+ and E2E SeURLLC+, respectively.
-VI-A Minimizing Energy Consumption for Transmission
-The optimization problem can be cast as
-[…]
-subject to $\epsilon_{i}\leq\epsilon^{\rm(ub)}_{i}\leq\epsilon^{\rm(th)}_{i},\quad\forall i>0$ (51b)
-$L\leq L_{\rm max},\quad$ (51c)
-$\overline{\mathsf{SINR}}_{s}\geq\gamma_{\rm s},\quad$ (51d)
-$P_{k}\leq P_{\rm tx},\quad\quad k=1,\cdots,N_{\rm tx}$ (51e)
-$P_{\rm tr}=\sum_{k=1}^{N_{\rm tx}}P_{k}$ is the average power consumption for transmission, given as $P_{\rm tr}=\sum_{j=0}^{N_{\rm ue}}\rho_{j}=\|\bm{\rho}\|^{2}$, where $\bm{\rho}=[\sqrt{\rho_{0}}\ \ldots\ \sqrt{\rho_{N_{\rm ue}}}]^{T}$, due to the unit-power centralized precoding vectors.
+The corresponding algorithm to this optimization problem is represented by E2E JPB-SeURLLC+.
+The optimization problem is formulated as follows:
+$\underset{\bm{\rho}\geq\textbf{0},L>L_{p}}{\textrm{minimize}}\quad$ $E_{\rm total}=\frac{L_{d}}{B}\Delta^{\rm tr}\|\bm{\rho}\|^{2}+\frac{L}{B}\left(P_{\rm FIXED}+\frac{1}{L}f_{1}+\frac{L_{d}}{L}f_{2}\right)\,$ (40a)
+subject to $\epsilon^{\rm(ub)}_{i}\leq\epsilon^{\rm(th)}_{i},\quad\forall i>0$ (40b)
+$L\leq L_{\rm max,c},\quad$ (40c)
+$R_{\rm s}\geq R_{\rm s}^{\rm th},\quad$ (40d)
+$\overline{\mathsf{SINR}}_{\rm s}\geq\gamma_{\rm s},\quad$ (40e)
+$P_{k}\leq P_{\rm tx},\quad\quad k=1,\ldots,N_{\rm tx}$ (40f)
+where
+$P_{\rm FIXED}=\sum_{k=1}^{N_{\rm tx}}P_{\textrm{AP},0}^{\rm tx}+\sum_{r=1}^{N_{\rm rx}}P_{\textrm{AP},0}^{\rm rx}+P_{\rm fixed}+N_{\rm GPP}\frac{P_{\rm cloud,0}^{\rm proc}}{\sigma_{\rm cool}}$ (41)
+$f_{1}=\frac{\Delta_{\rm cloud}^{\rm proc}}{\sigma_{\rm cool}C_{\rm max}}\frac{B}{10^{9}}\Bigg{(}C_{\textrm{ch-est}}+C_{\textrm{prec-comp}}+C_{\textrm{prec-comp}}^{\rm s}$
+$\quad+C_{\rm se-detector}\Bigg{)}$ (42)
+$f_{2}=\frac{\Delta_{\rm cloud}^{\rm proc}}{\sigma_{\rm cool}C_{\rm max}}\frac{B}{10^{9}}\left(20MN_{\rm ue}N_{\rm tx}+12MN_{\rm tx}+C_{{\rm se-prep}}\right),$ (43)
+where $C_{\rm se-detector}$ is the summation of the terms in $C_{\rm se-comp}$ that does not have the multiplying factor of $L_{d}$ as it is included in $f_{1}$.
+On the other hand, $C_{\rm se-prep}$ includes the terms with $L_{d}$ appearing in $f_{2}$.
 Constraints (40b) and (40c) correspond to the URLLC requirements.
-The $\gamma_{\rm s}$ is the required sensing SINR that is selected according to the target detection performance requirement and $P_{\rm tx}$ is the maximum transmit power per AP.
+Constraints (40d) and (40e) correspond to the sensing requirements, where $R_{s}^{\rm th}$ is the refreshing rate threshold and $\gamma_{\rm s}$ is the required sensing SINR that is selected according to the target detection performance requirement and $P_{\rm tx}$ is the maximum transmit power per AP.
+We define a new variable $\bar{L}$ where $L_{d}=L-L_{p}\leq\frac{1}{\bar{L}}$.
+Then, the objective function (40a) is equivalent to
+$\mathsf{F}\triangleq\frac{1}{B}\left(LP_{\rm FIXED}+\left(L-L_{p}\right)f_{2}+\Delta^{\rm tr}\frac{\|\bm{\rho}\|^{2}}{\bar{L}}\right).$ (44)
 The aforementioned problem is challenging to solve due to its non-convex nature and the high coupling of variables.
@@
 The optimal solution $\{\bm{\rho}^{\star},L^{\star}\}$ of the problem given below is also an optimal solution to (40):
-$\underset{\bm{\rho},\bm{\chi},\textbf{r}\geq\textbf{0},L>L_{p},\bar{L}>0}{\textrm{minimize}}\quad\frac{\|\bm{\rho}\|^{2}}{\overline{L}}\,$ (52a)
+$\underset{\bm{\rho},\bm{\chi},\textbf{r}\geq\textbf{0},L>L_{p},\bar{L}>0}{\textrm{minimize}}\quad\mathsf{F}$ (45a)
 […]
@@
 $\left\|\textbf{F}_{k}\bm{\rho}\right\|\leq\sqrt{P_{\rm tx}},\quad k=1,\ldots,N_{\rm tx},$ (45f)
-$L\leq L_{\rm max},$ (52g)
+$L\leq\min\left(L_{\rm max,c},\frac{B}{R_{\rm s}^{\rm th}}\right),$ (45g)
 where ${\bf F}_{k}=\mathrm{diag}\left(\sqrt{\mathbb{E}\left\{\|{\bf w}_{0,k}\|^{2}\right\}},\ldots,\sqrt{\mathbb{E}\left\{\|{\bf w}_{N_{\rm ue},k}\|^{2}\right\}}\right)$.
@@
 Finally, the convex problem that is solved at the $c^{\rm th}$ iteration becomes
-$\underset{\bm{\rho},\bm{\chi},\textbf{r}\geq\textbf{0},L,\bar{L}>L_{p},\chi_{0}\geq 0}{\textrm{minimize}}\quad\frac{\|\bm{\rho}\|^{2}}{\overline{L}}+\lambda\chi_{0}\,,$ (53a)
-subject $\textrm{subject to}\quad{L-L_{p}}\leq\frac{2}{\overline{L}^{(c-1)}}-\frac{\overline{L}}{\left(\overline{L}^{(c-1)}\right)^{2}},$ (53b)
+$\underset{\bm{\rho},\bm{\chi},\textbf{r}\geq\textbf{0},L>L_{p},\bar{L}>0,\chi_{0}\geq 0}{\textrm{minimize}}\quad\mathsf{F}+\lambda\chi_{0}$ (46a)
+$\textrm{subject to}\quad{L-L_{p}}\leq\frac{2}{\overline{L}^{(c-1)}}-\frac{\overline{L}}{\left(\overline{L}^{(c-1)}\right)^{2}},$ (46b)
 $\ln\left(1+\chi_{i}\right)-\chi_{i}+2\frac{\sqrt{\rho_{i}}^{(c-1)}\mathsf{b}_{i}^{2}\,\sqrt{\rho_{i}}}{r_{i}^{(c-1)}}-r_{i}\left(\frac{\sqrt{\rho_{i}}^{(c-1)}\mathsf{b}_{i}}{r_{i}^{(c-1)}}\right)^{2}$
@@
 […]
-The steps of FPP-SCA and CCP procedure are outlined in Algorithm 1.
+The steps of the FPP-SCA and CCP procedure, E2E JPB-SeURLLC+ algorithm, are outlined in Algorithm 1.
 We empirically observed that setting $\overline{L}^{(0)}=\frac{1}{L_{\rm max}-L_{p}}$, $\rho_{0}^{(0)}=0$ and $\sqrt{\rho_{i}}^{(0)}=10^{-3}\sqrt{P_{\rm tx}/N_{\rm ue}}$ for $i>0$ yields satisfactory results.
-Algorithm 1 FPP-SCA and CCP Procedure for Problem (52).
-1:Initialization: Set arbitrary $\bm{\rho}^{(0)},\textbf{r}^{(0)}\geq\textbf{0}$, the solution accuracy parameters $\epsilon,\epsilon_{\chi}>0$, $\overline{L}^{(0)}=\frac{1}{L_{\rm max}-L_{p}}$, and $\lambda>0$.
-Set the iteration counter to $c=0$, and a maximum iteration number $c_{\rm max}$.
-Moreover, let $\mathsf{F}^{(0)}=\infty$ be the initial value of objective function and $\Delta\mathsf{F}^{(c)}=\mathsf{F}^{(c)}-\mathsf{F}^{(c-1)}$ with $\mathsf{F}^{(-1)}=0$.
-2:$c\leftarrow c+1$.
-3:While $\Delta\mathsf{F}^{(c-1)}\geq\epsilon$ and $c\leq c_{\rm max}$, do
-4:Set $\bm{\rho}^{(c)}$, $\overline{L}^{(c)}$, and $\textbf{r}^{(c)}$ to the solution of the convex problem in (53), where the previous iterates $\bm{\rho}^{(c-1)}$, $\overline{L}^{(c-1)}$, and $\textbf{r}^{(c-1)}$ are taken as constant.
-5:If $\chi_{0}<\epsilon_{\chi}$, set $\chi_{0}=0$ for the next iteration.
-6:End while
-7:Output: The transmit power coefficients $\bm{\rho}^{(c)}$ and the blocklength $L^{(c)}$.
-VI-B Minimizing E2E Energy Consumption
-In this subsection, our objective is to minimize the E2E energy consumption. The optimization problem is formulated as follows
+Algorithm 1 FPP-SCA and CCP Procedure for Solving Problem (45), E2E JPB-SeURLLC+ algorithm
+1:Initialization: Initialize $\rho_{0}^{(0)}=0$, $\sqrt{\rho_{i}}^{(0)}=10^{-3}\sqrt{P_{\rm tx}/N_{\rm ue}}$ for $i>0$, $\textbf{r}^{(0)}\geq\textbf{0}$, and blocklength $\overline{L}^{(0)}=\frac{1}{L_{\rm max}-L_{p}}$.
+Set solution accuracy parameters $\epsilon,\epsilon_{\chi}>0$, and $\lambda>0$.
+Set iteration counter $c=0$, maximum number of iterations $c_{\rm max}$, the initial objective value $\mathsf{F}^{(0)}=\infty$ and define the improvement metric as $\Delta\mathsf{F}^{(c)}=\mathsf{F}^{(c-1)}-\mathsf{F}^{(c)}$, and $\Delta\mathsf{F}^{(0)}=\infty$.
 […]
+3:while $\Delta\mathsf{F}^{(c-1)}\geq\epsilon$ and $c\leq c_{\rm max}$ do
+4:  Solve the convexified problem in (46) using the previous iterates $\bm{\rho}^{(c-1)}$, $\overline{L}^{(c-1)}$, and $\textbf{r}^{(c-1)}$ as constants, and update $\bm{\rho}^{(c)}$, $\overline{L}^{(c)}$, and $\textbf{r}^{(c)}$ accordingly.
+5:  if $\chi_{0}<\epsilon_{\chi}$ then
+6:   Set $\chi_{0}=0$ for the next iteration.
+7:  end if
+8:  $c\leftarrow c+1$
+9:end while
+10:Output: Transmit power coefficients $\bm{\rho}^{(c)}$ and the optimized blocklength $L^{(c)}$.
+VII Numerical Results
 […]
-The total power consumption can be rewritten as
-$P_{\rm total}=P_{\rm FIXED}+\Delta^{\rm tr}\|\bm{\rho}\|^{2}+\frac{1}{L}f_{1}+\frac{L_{d}}{L}f_{2}$ (55)
-where
-$P_{\rm FIXED}=\sum_{k=1}^{N_{\rm tx}}P_{\textrm{AP},0}^{\rm tx}+\sum_{r=1}^{N_{\rm rx}}P_{\textrm{AP},0}^{\rm rx}+P_{\rm fixed}+\frac{P_{\rm cloud,0}^{\rm proc}}{\sigma_{\rm cool}}$ (56)
-$f_{1}=\frac{\Delta_{\rm cloud}^{\rm proc}}{\sigma_{\rm cool}C_{\rm max}}\frac{B}{10^{9}}\Bigg{(}C_{\textrm{ch-est}}+C_{\textrm{prec-comp}}+C_{\textrm{prec-comp}}^{\rm s}$
-$\quad+C_{\rm se-detector}\Bigg{)}$ (57)
-$f_{2}=\frac{\Delta_{\rm cloud}^{\rm proc}}{\sigma_{\rm cool}C_{\rm max}}\frac{B}{10^{9}}\left(20MN_{\rm ue}N_{\rm tx}+12MN_{\rm tx}+C_{{\rm se-prep},m}\right).$ (58)
-Similar to the approach of problem P1, we define a new variable $\bar{L}$ where $L-L_{p}\leq\frac{1}{\bar{L}}$. Then, the objective function (54a) is equivalent to
-$\mathsf{F_{2}}\triangleq\frac{1}{B}\left(L\left(P_{\rm FIXED}\right)+\left(L-L_{p}\right)f_{2}+\Delta^{\rm tr}\frac{\|\bm{\rho}\|^{2}}{\bar{L}}\right).$ (59)
-The non-convex optimization problem in (54) can be solved using the same approach in P1 as follows:
-$\underset{\bm{\rho},\bm{\chi},\textbf{r}\geq\textbf{0},L,\bar{L}>L_{p},\chi_{0}\geq 0}{\textrm{minimize}}\quad\mathsf{F_{2}}+\lambda\chi_{0},$ (60a)
+The large-scale fading coefficients, shadowing parameters, probability of LOS, and the Rician factors are simulated based on the 3GPP Urban Microcell model, defined in [47, Table B.1.2.1-1, Table B.1.2.1-2, Table B.1.2.2.1-4].
+The path losses for the Rayleigh fading target-free channels are also modeled by the 3GPP Urban Microcell model with the difference that the channel gains are multiplied by an additional scaling parameter equal to 0.30.30.3 to suppress the known parts of the target-free channels due to LOS and permanent obstacles [40].
+The sensing channel gains are computed by the bi-static radar range equation [7].
+The carrier frequency, the bandwidth, and the noise variance are set to 1.91.91.9 GHz, 200200200 KHz, and 114-114- 114 dBm, respectively.
+The number of pilot symbols is $L_{p}=10$.
+The regularization parameter $\delta$ in (5) is set to the noise variance.
+The spatial correlation matrices for the communication channels are generated by using the local scattering model in [34, Sec.
+2.5.3].
+The RCS of the target is modeled by the Swerling-I model with $\sigma_{\rm rcs}=0$ dBsm.
+For all the UEs, the packet size, maximum transmission delay, and DEP threshold are $b_{i}=256$ bits, $D_{i}^{\rm(th)}=1$ ms, and $\epsilon_{i}^{\rm(th)}=10^{-5}$, respectively.
+The sensing SINR threshold is $\gamma_{s}=0$ dB, the false alarm probability threshold is $P_{\rm fa}=0.03$, and the refreshing rate threshold is $R_{\rm s}^{\rm th}=10$ updates per second [6, Table 6.2-1], unless otherwise stated.
+The remaining parameters are detailed in Table III, where the values are consistent with those in [18].
+TABLE III: SIMULATION PARAMETERS
+$\Delta^{\rm tr},\sigma_{\rm cool}$ 4, 0.9 $P_{\rm fixed}$ 120 W
+$P_{\textrm{AP},0}^{\rm tx},P_{\textrm{AP},0}^{\rm rx}$ $6.8\cdot M$ W $P_{\rm cloud,0}^{\rm proc}$ 81 W
+$\Delta_{\rm cloud}^{\rm proc}$ 288 W $C_{\rm max}$ 700.94 GOPS
+We compare the performance of the proposed algorithm, E2E JPB-SeURLLC+, which aims to jointly optimize power and blocklength to minimize E2E energy consumption, against two benchmark schemes: (i) JPB-SeURLLC+, which minimizes only the transmission energy consumption, and (ii) E2E JPB-URLLC, which targets E2E energy minimization in URLLC systems without sensing integration.
+We also evaluate performance under both clutter-aware and clutter-unaware MAPRT detectors, as defined in (11) and (12), respectively.
+Our analysis focuses on the total energy consumed to complete one cycle of communication and sensing tasks.
+The communication task requires satisfying a maximum DEP threshold for each UE, while the sensing task involves detecting a target at a candidate location with a specified false alarm probability threshold.
+The total E2E energy includes ISAC transmission, sensing reception, and processing components for both tasks.
+Figure 3: Convergence condition for E2E JPB-SeURLLC+, JPB-SeURLLC+, and E2E JPB-URLLC algorithms with clutter-aware detector and $N_{\rm rx}=2$.
+Figure 4: (a) Detection probability, (b) transmission energy, and (c) E2E energy vs. sensing SINR threshold with $N_{\rm rx}=2$.
+Figure 5: Maximum blocklength threshold and network availability vs. refreshing rate threshold for $N_{\rm rx}=2$ and the E2E JPB-SeURLLC+ algorithm.
+Figure 6: (a) Detection probability, (b) transmission, and (c) total energy consumption vs. DEP threshold for $N_{\rm rx}=2$, $\gamma_{\rm s}=0$ dB, and $R_{\rm s}^{\rm th}=1000$ updates/s.
+Figure 7: Breakdown of energy consumption for completing one sensing and communication task with clutter-aware detector, $N_{\rm rx}=2$, and $\gamma_{\rm s}=0$ dB.
+Figure 8: (a) Computational complexity for one sensing task, (b) sensing GOPS, (c) number of required GPPs vs. blocklength.
+(a)(b)
+Figure 9: (a) Number of antenna elements per AP, and (b) total energy consumption for a minimum detection probability of 0.90.90.9 with E2E JPB-SeURLLC+ algorithm and $N_{\rm rx}=1,2$.
+Fig.
+3 illustrates the convergence behavior of the proposed and benchmark algorithms.
+The objective of the E2E algorithms is to minimize the total energy consumption, whereas the JPB-SeURLLC+ algorithm focuses solely on minimizing transmission energy consumption.
+In the figure, the values for JPB-SeURLLC+ are scaled by $10^{5}$ since they are significantly lower compared to the total energy values.
+In the algorithms, the solution accuracy parameters are set as $\epsilon=10^{-3}$, $\epsilon_{\chi}=10^{-6}$, $\lambda=10$, and the maximum iteration is set to $c_{\rm max}=30$.
+In the following, we first compare the performance of the proposed E2E JPB-SeURLLC+ algorithm with relevant benchmarks, followed by a comprehensive analysis of the proposed algorithm under various communication and sensing requirements.
+Then, we investigate the interplay between the number of receive APs, the number of antenna elements per AP, and the complexity level of the detector to balance the sensing performance and energy consumption in the system.
+Detection probability results are obtained empirically through Monte Carlo simulation.
+VII-A Impact of Sensing and Communication Requirements
+The effect of the sensing SINR threshold $\gamma_{\rm s}$ is evaluated in Fig.
+4 for $N_{\rm rx}=2$.
+As shown in Fig.
+4(a), detection probability generally increases with $\gamma_{s}$, but performance varies based on the detector and resource allocation strategy.
+Clutter-aware detectors consistently outperform clutter-unaware ones due to their advanced signal processing.
+As this detector estimates the target-free channels, its performance does not change significantly with increasing $\gamma_{\rm s}$ while the performance of clutter-unaware detector accordingly.
+The proposed E2E JPB-SeURLLC+ with a clutter-aware detector achieves the highest detection probability up to 0.980.980.98, while the same algorithm with a clutter-unaware detector reaches only detection probability of 0.630.630.63.
+The JPB-SeURLLC+ algorithm that prioritizes minimizing transmission energy consumption, as opposed to E2E energy minimization, tend to result in lower detection performance.
+This is because the E2E algorithms try to minimize the total energy consumption by reducing the blocklength.
+This approach allows higher transmission power levels–as shown in Fig.
+4(b)– to meet the requirements, which in turn enhance detection capability.
+In contrast, E2E JPB-URLLC fails to exceed 0.5 detection probability, even with a clutter-aware detector—highlighting the need to jointly optimize sensing and communication in URLLC systems.
 […]
-VII Numerical Results
-In this section, we present numerical results to evaluate the performance of the proposed joint blocklength and power allocation algorithms.
-We consider a total area of 500 $\rm m\times$ 500 $\rm m$, with the sensing location positioned at the center.
-The $N_{\rm tx}=16$ ISAC transmit APs are uniformly distributed in the area.
-We consider the number of sensing receive APs equal to $N_{\rm rx}=1$ or $N_{\rm rx}=2$ where the first AP is located at $(200,250)$ and the second one is located at $(300,250)$.
-The number of antenna elements per AP is set to $M=4$.
-The total number of URLLC UEs in the network is $N_{\rm ue}=8$.
-We set $P_{\rm tx}=100$ mW and the uplink pilot transmission power for each UE is $50$ mW.
-The large-scale fading coefficients, shadowing parameters, probability of LOS, and the Rician factors are simulated based on the 3GPP Urban Microcell model, defined in [43, Table B.1.2.1-1, Table B.1.2.1-2, Table B.1.2.2.1-4].
-The path losses for the Rayleigh fading target-free channels are also modeled by the 3GPP Urban Microcell model with the difference that the channel gains are multiplied by an additional scaling parameter equal to $0.3$ to suppress the known parts of the target-free channels due to LOS and permanent obstacles [34].
-The sensing channel gains are computed by the bi-static radar range equation [9].
-The carrier frequency, the bandwidth, and the noise variance are set to $1.9$ GHz, $200$ KHz, and $-114$ dBm, respectively.
-The number of pilot symbols is $L_{p}=10$.
-The spatial correlation matrices for the communication channels are generated by using the local scattering model in [35, Sec.
-2.5.3].
-The RCS of the target is modeled by the Swerling-I model with $\sigma_{\rm rcs}=0\,\rm dBsm$.
-For all the UEs, the packet size, maximum transmission delay, and DEP threshold are $b_{i}=256$ bits, $D_{i}^{\rm(th)}=1$ ms, and $\epsilon_{\rm th}=10^{-5}$, respectively.
-In the proposed algorithms, $\epsilon=10^{-3}$, $\epsilon_{\chi}=10^{-6}$, $\lambda=10$, and $c_{\rm max}=30$.
-The sensing SINR threshold is set to $\gamma_{s}=0$ dB, unless otherwise stated.
-The remaining parameters are detailed in Table II, where the values are consistent with those in [30], except for a scaling of the total processing capacity and the corresponding power of a DU by a factor of ten to accommodate all the processing in a single DU.
-TABLE II: Simulation Parameters
-$\Delta^{\rm tr},\sigma_{\rm cool}$ 4, 0.9 $P_{\rm fixed}$ 120 W
+Fig.
+5 illustrates the maximum blocklength threshold and network availability as a function of the refreshing rate threshold.
+The network availability represents the percentage of cases where the optimization problem is feasible, meaning that all the requirements can be satisfied [16].
+As the required refreshing rate exceeds 1000 updates/s, the system becomes increasingly constrained, resulting in a notable reduction in allowable blocklength.
+Specifically, the maximum blocklength is nearly halved when $R_{\rm s}^{\rm th}=2000$.
+However, such reduced blocklengths may not satisfy reliability requirements, leading to infeasibility.
+Accordingly, the network availability drops to 80% at 1800 updates/s and further declines to 60% at 2000 updates/s.
+Figs.
+6a–c illustrate the impact of the DEP threshold on sensing performance, transmission energy, and total energy consumption, respectively.
+As shown in Fig.
+6(a), stricter reliability requirements slightly enhance detection probability, since the communication task requires higher power and/or longer blocklengths to meet lower DEP thresholds, which in turn enhances the sensing performance.
+However, this improvement comes at the cost of increased energy consumption, as evident in Fig.
+6(b) and Fig.
+6(c).
+Notably, the E2E JPB-SeURLLC+ algorithm consistently outperforms the JPB-SeURLLC+ algorithm in terms of both sensing performance and energy consumption.
+It is worth mentioning that higher delay threshold allows the system to operate with higher blocklength.
+However, it does not effect the results since the algorithm still chooses a smaller blocklength to minimize the energy consumption.
+VII-B Impact of Detector and Number of APs
+Fig.
+7 presents a detailed breakdown of task-oriented energy consumption across system components and operations, including ISAC transmit APs, sensing receive APs, communication and sensing processing, and an “Others” category representing load-independent and idle-mode power consumption in the cloud.
+As shown, the majority of energy consumption is attributed to sensing processing tasks, ISAC transmit APs, and the cloud’s load-independent and idle-mode power usage.
+To further explore the system behavior, Figs.
+8a-c analyze the impact of blocklength, detector type, and the number of sensing receive APs on computational complexity, GOPS, and the number of required GPPs, respectively.
+This multi-dimensional analysis highlights how detector complexity and network configuration influence overall system efficiency.
+As shown in Fig.
+8(a) and Fig.
+8(b), clutter-aware detector with $N_{\rm rx}=2$ has the highest computational complexity and GOPS requirements among other configurations.
+Although longer blocklengths generally increase the number of operations for one task, they reduce the number of operations per unit time, thereby lowering GOPS requirements.
+This, in turn, reduces the number of active GPPs as shown in Fig.
+8(c) and consequently decreases energy consumption through the $N_{\rm GPP}P_{\rm cloud,0}^{\rm proc}$ term in equation (32).
+Clutter-unaware detector maintains much lower complexity compared to the clutter-aware detector.
+However, a comparison between the clutter-aware and clutter-unaware curves reveals distinct trends.
+For the clutter-aware detector, the computational complexity remains approximately constant as blocklength increases, whereas for the clutter-unaware detector, complexity grows significantly with larger blocklengths.
+Interestingly, the opposite trend is observed in the GOPS curves.
+This behavior stems from the interplay between blocklength-dependent and blocklength-independent terms in the computational complexity expressions—specifically, equation (V-B) for the clutter-unaware detector and Table II for the clutter-aware detector.
+Based on these observations, we conclude that for the clutter-aware detector, blocklength-independent terms dominate the computational complexity, whereas for the clutter-unaware detector, blocklength-dependent terms are the primary contributors.
 […]
-$\Delta_{\rm cloud}^{\rm proc}$ 740 W $C_{\rm max}$ 1800 GOPS
-Figure 2: (a) Transmission energy consumption, (b) E2E energy consumption, and (c) breakdown of energy consumption over one transmission for different optimization algorithms, for clutter-aware target detector with $N_{\rm rx}=2$ and $\gamma_{s}=0$ dB.
-We compare the performance of two proposed algorithms: i) SeURLLC+ aiming to minimize the transmission energy consumption, and ii) E2E SeURLLC+ aiming to minimize the E2E energy consumption.
-Additionally, we include two baseline algorithms for URLLC systems without sensing integration: (i) URLLC-only, which minimizes transmission energy consumption, and (ii) E2E URLLC-only, which minimizes E2E energy consumption.
-We further compare results using clutter-aware and clutter-unaware MAPRT detectors, as defined in (17) and (18), respectively.
-Fig.
-2(a) illustrates the transmission energy consumption for four algorithms with the clutter-aware detector.
-The results show that integrating sensing into the system necessitates higher transmission energy.
-Notably, the URLLC-only and SeURLLC+ achieve lower transmission energy compared to their E2E counterparts, as their primary objective is to minimize transmission energy consumption.
-The E2E energy consumption is depicted in Fig.
-2(b).
-We observe that E2E energy minimization results in a substantial reduction in total energy consumption by approximately $50\%$ in URLLC-only scenarios, while its impact is less pronounced in ISAC scenarios, resulting in only a marginal decrease.
-The E2E algorithms try to minimize the total energy consumption by reducing the blocklength.
-However, higher power coefficients are required to meet the reliability requirements for URLLC UEs.
-In SeURLLC+, higher power coefficients for sensing allow lower blocklength, narrowing the gap in optimal blocklengths between SeURLLC+ and E2E SeURLLC+.
-Fig.
-2c provides a detailed breakdown of energy consumption by system components and different operations, including ISAC transmit APs, sensing receive APs, communication and sensing processing, and “Others,” which represent load-independent and idle mode power consumption in the cloud.
-For both algorithms, most of the energy consumption is attributed to sensing processing tasks, ISAC APs, and load-independent and idle mode power consumption in the cloud.
-However, the E2E SeURLLC+ algorithm slightly reduces the corresponding energy consumption.
-Figure 3: (a) Blocklength, (b) E2E energy consumption, (c) detection probability, and (d) sensing update rate vs.
-decoding error probability threshold with clutter-aware target detector, $N_{\rm rx}=2$ and $\gamma_{s}=0$ dB.
-VII-A Impact of URLLC Requirements
-Fig.
-3 illustrates the blocklength, E2E energy consumption, detection probability and sensing update rate as functions of DEP threshold.
-Generally, the blocklength in Fig.
-3(a) decreases as the reliability threshold is relaxed in all algorithms except for URLLC-only.
-The URLLC-only algorithm consistently uses the maximum available blocklength, i.e., $L_{\rm max}$.
-By integrating sensing, the optimal blocklength is significantly reduced compared to URLLC-only, as meeting sensing constraint requires higher power consumption resulting in lower blocklength.
-Relaxing the reliability constraint would generally reduce the E2E energy consumption.
-However, ISAC scenarios demand higher energy consumption as it is expected as shown in Fig.
-3(b).
-Fig.
-3(c) and Fig.
-3(d) demonstrate the impact of the DEP threshold on sensing performance.
-Relaxing the DEP constraint slightly reduces detection probability while increasing the sensing update rate, due to smaller blocklengths at higher DEP thresholds.
-Notably, the E2E SeURLLC+ algorithm outperforms the SeURLLC+ algorithm in both sensing and communication performance.
-It is worth mentioning that higher delay threshold allows the system to operate with higher blocklength.
-However, it does not effect the results since the algorithm aims to minimize the energy consumption.
-Therefore, it still chooses a smaller blocklength.
-VII-B Impact of Sensing Requirement
-The impact of the sensing SINR threshold, $\gamma_{s}$, on the performance of the E2E SeURLLC+ algorithm is evaluated in Figs.
-4a-c in terms of E2E energy consumption, detection probability, and sensing update rate, respectively.
-Comparing Fig.
-4(a) with Fig.
-4(b), we observe that both detection probability and energy consumption generally increase with a higher SINR threshold.
-The clutter-aware detector with $N_{\rm rx}=2$ provides the highest detection probability, consistently above 0.9, whereas the clutter-unaware detector cannot achieve a detection probability over 0.6, despite offering the lowest energy consumption.
-Notably, using a clutter-aware detector while disabling one RX-AP maintains a detection probability above 0.8 with around 43% energy saving.
-This suggests a potential optimization strategy: turning off one sensing receive AP can halve the energy consumption, significantly enhancing overall energy efficiency.
-Higher sensing SINR threshold demands higher blocklength, resulting in lower sensing update rate as shown in Fig.
-4(c).
-Comparing Fig.
-4(b) and Fig.
-4(c), it is evident that increasing the sensing SINR threshold significantly decreases the sensing update rate without substantially improving detection probability.
-Figure 4: (a) E2E energy consumption, (b) detection probability, and (c) sensing update rate vs. sensing SINR threshold with E2E SeURLLC+ algorithm.
-VII-C Impact of Number of APs
-We explore the impact of the number of transmit and receive APs on the network performance in terms of E2E energy consumption, detection probability, and sensing update rate as shown in Figs.
-5a-c, respectively.
-The results in Fig.
-5(a) and Fig.
-5(b) indicate that increasing the number of receive APs generally raises both E2E energy consumption and detection probability.
-For $N_{\rm rx}=2$, the detection probability substantially improves with only a marginal raise in energy consumption.
-However, further increasing to $N_{\rm rx}=4$ offers little improvement in detection probability while significantly increasing energy consumption.
-Moreover, the results in Fig.
-5(c) show a consistent sensing update rate across all curves, especially for $N_{\rm tx}=8$ and $N_{\rm tx}=12$, indicating that $N_{\rm rx}$ has minimal impact on $R_{s}$.
-This suggests the optimization algorithm is more constrained by URLLC requirements than by sensing requirements.
-A comparison between $N_{\rm tx}=8$ and $N_{\rm tx}=12$ in Fig.
-5(c) reveals that a higher $N_{\rm tx}$ enhances URLLC satisfaction due to elevated SINR, resulting in a reduced blocklength and consequently a higher sensing update rate.
-It is observed that when $N_{\rm rx}=1$ or $N_{\rm rx}=2$, increasing the number of transmit APs does not significantly affect E2E energy consumption.
-This is because the rise in energy consumption from the additional APs is offset by the decrease in blocklength.
-Thus, the overall E2E energy consumption remains relatively unchanged.
-Conversely, when $N_{\rm rx}=4$, the influence of $N_{\rm tx}$ becomes more pronounced, as the energy consumption for sensing tasks increases non-linearly with the product of $N_{\rm tx}$ and $N_{\rm rx}$.
-When $N_{\rm tx}=16$, the system’s behavior concerning E2E energy consumption and detection probability follows a similar trend, showing an increase in energy consumption and a decrease in detection probability.
-However, $R_{s}$ declines compared to when $N_{\rm tx}=12$, indicating a higher blocklength for $N_{\rm tx}=16$.
-This occurs because common power coefficients are used to maintain a favorable beam pattern, and increasing the number of transmit APs may constrain power due to some APs being closer to the UEs.
-As a result, a higher blocklength is necessary to fulfill the reliability requirements.
-These findings underscore the importance of optimizing the number of transmit and receive APs.
-The results of detection probability is depicted in Fig.
-5(b).
-In general, a decrease in blocklength translates to a lower detection probability explaining the results for $N_{\rm tx}=12$.
-However, we observe that detection probability for $N_{\rm tx}=16$ still decreases although blocklength is higher compared to $N_{\rm tx}=12$.
-This is because the power coefficients required to satisfy the sensing SINR requirement might be lower with higher $N_{\rm tx}$, resulting in a more pronounced negative impact on detection probability.
-This negative impact outweighs the benefits of increased blocklength, leading to a lower detection probability.
-The findings emphasize the importance of optimizing the number of transmit and receive APs to balance energy consumption, detection probability and sensing update rate.
-Higher numbers of transmit APs can enhance URLLC satisfaction and reduce blocklength resulting in higher sensing update rate, but must be carefully managed to avoid an unnecessary increase in energy consumption and detection probability degradation.
-Figure 5: (a) E2E energy consumption, (b) detection probability, and (c) sensing update rate vs. the number of transmit ISAC APs for $N_{\rm rx}=1,2,4$ and E2E SeURLLC+ algorithm.
-VIII Conclusion
-In this paper, we proposed a joint blocklength and power control algorithm for a downlink CF-mMIMO system with multi-static sensing and URLLC UEs.
-We formulated two non-convex optimization problems focused on minimizing transmission and E2E energy consumption, and conducted GOPS analysis for both communication and sensing tasks.
-Numerical results demonstrated the effectiveness of the E2E energy minimization algorithm, especially in scenarios without sensing capabilities.
-Integrating sensing reduces the optimal blocklength compared to URLLC-only scenarios due to the higher power demands of sensing, which leads to a lower optimal blocklength.
-Moreover, the E2E algorithms minimize total energy consumption by further reducing the blocklength.
-Additionally, we observed that higher sensing SINR thresholds significantly lower the sensing update rate without substantially improving detection probability.
-Our findings also indicated that increasing the number of sensing receive APs raises sensing-related processing energy, suggesting that deactivating some APs can help reduce E2E energy consumption in CF-mMIMO ISAC systems.
-While increasing the number of transmit APs can enhance URLLC satisfaction and sensing update rate by reducing blocklength, careful management is necessary to prevent excessive energy consumption and degradation in detection probability.
+VIII Conclusion and Future Directions
+[…]
 Appendix A Proof of Theorem 1
 Let us define a new optimization variable, denoted by $\overline{L}$, where $L-L_{p}\leq 1/\overline{L}$.
-Then, minimizing the objective function is equivalent to minimizing the following convex function (quadratic-over-linear function) $\frac{\|\bm{\rho}\|^{2}}{\overline{L}}$ Minimizing this function, at the optimal solution, holds that $\overline{L}=1/L_{d}=1/(L-L_{p})$.
+Then, minimizing the objective function is equivalent to minimizing the convex function (quadratic-over-linear plus affine function) $\mathsf{F}$ in (44).
+Minimizing this function, at the optimal solution, leads to $\overline{L}=1/L_{d}=1/(L-L_{p})$.
 The reliability constraints in (40b) can be written in the form of
@@
 which will not destroy optimality since we want to minimize $r_{i}$ to increase the left-hand side of the SINR constraint.
-From (38), the sensing constraint in (51d) are expressed as
+From (27), the sensing constraint in (40e) are expressed as
 $\bm{\rho}^{T}\left(\gamma_{\rm s}\textbf{B}_{D}-M\,\textbf{A}_{D}\,\right)\bm{\rho}\!\leq\!-\gamma_{\rm s}MN_{\rm rx}\sigma_{n}^{2}.$ (51)
-Finally, we define $\textbf{F}_{k}=\textrm{diag}\left(\sqrt{\mathbb{E}\left\{\|\textbf{w}_{0,k}\|^{2}\right\}},\ldots,\sqrt{\mathbb{E}\left\{\|\textbf{w}_{N_{\rm ue},k}\|^{2}\right\}}\right)$ and rewrite the per-AP power constraints in (51e) in SOC form in terms of $\bm{\rho}$ as $\left\|\textbf{F}_{k}\bm{\rho}\right\|\leq\sqrt{P_{\rm tx}},$ , for $k=1,\ldots,N_{\rm tx}$.
+Finally, we define $\textbf{F}_{k}=\textrm{diag}\left(\sqrt{\mathbb{E}\left\{\|\textbf{w}_{0,k}\|^{2}\right\}},\ldots,\sqrt{\mathbb{E}\left\{\|\textbf{w}_{N_{\rm ue},k}\|^{2}\right\}}\right)$ and rewrite the per-AP power constraints in (40f) in SOC form in terms of $\bm{\rho}$ as $,\left\|\textbf{F}_{k}\bm{\rho}\right\|\leq\sqrt{P_{\rm tx}},$ for $k=1,\ldots,N_{\rm tx}$.